### v1

# Predict. Optimize. Revise. On Forecast and Policy Stability in Energy Management Systems

Abstract
This research addresses the challenge of integrating forecasting and optimization in energy management systems, focusing on the impacts of switching costs, forecast accuracy, and stability. It proposes a novel framework for analyzing online optimization problems with switching costs and enabled by deterministic and probabilistic forecasts. Through empirical evaluation and theoretical analysis, the research reveals the balance between forecast accuracy, stability, and switching costs in shaping policy performance. Conducted in the context of battery scheduling within energy management applications, it introduces a metric for evaluating probabilistic forecast stability and examines the effects of forecast accuracy and stability on optimization outcomes using the real-world case of the Citylearn 2022 competition. Findings indicate that switching costs significantly influence the trade-off between forecast accuracy and stability, highlighting the importance of integrated systems that enable collaboration between forecasting and operational units for improved decision-making. The study shows that committing to a policy for longer periods can be advantageous over frequent updates. Results also show a correlation between forecast stability and policy performance, suggesting that stable forecasts can mitigate switching costs. The proposed framework provides valuable insights for energy sector decision-makers and forecast practitioners when designing the operation of an energy management system.
Index Terms:
Energy Management Systems, Forecasting, Optimization, Switching Costs, Forecast Stability, Model Predictive Control
I Introduction
Managing energy assets in a grid system is a complex task that requires decision-making under uncertainty. The dynamics and stochasticity of the environment create a constantly evolving and inherently complex system. This reality necessitates making decisions with an incomplete knowledge of the future and for a limited time window, a scenario frequently approached through the lens of online optimization. In many applications, the future is reasonably predictable. Indeed, it is possible to make predictions about future cost function with a significant accuracy, particularly near future.
To effectively manage this, it is useful to combine the forecast of the problem environment with the optimization of the decision-making process. This combination of forecasting and decision-making is often referred to as the predict, then optimize problem, or an online optimization with predictions. The online optimization problems with predictions are commonly seen in energy applications such as electric vehicle charging [1] [2], battery scheduling [3][4] and energy and flexibility dispatch problems [5] [6]. In number of these applications, the decision-maker is faced with switching costs, which are the costs associated with updating the operational plan. For instance, transitioning from $x_{t-1}$ to $x_{t}$ typically incurs a switching cost or ramp cost, $d(x_{t}-x_{t-1})$. Consequently, the stage cost becomes time-coupled and is defined as: $C_{t}(x_{t-1},x_{t}):=f_{t}(x_{t})+d(x_{t}-x_{t-1})$.
In the energy sector, the switching costs can arise from various factors, such as the physical limitations of ramping up or down energy production [7], managing server load balancing [8] and payment of network fees associated with energy trading [9]. While there has been a significant amount of research on online optimization problems with predictions, there is a gap in the literature on the impact of switching costs on the interaction between forecast model and the optimization algorithm. We address this gap by studying two aspects of the combined system: time-coupling and forecast stability.
Switching costs are a direct consequence of time-coupling, when the action $x_{t}$ at time $t$ influences the future actions $x_{t+1}$ and system states $S_{t+1}$, as defined by Lin et al. [8]. Errors in forecasts act as ’switching incentive‘ compelling policy revisions using more accurate predictions aligned with up-to-date information and the actual state of the system. This perspective suggests a balancing act between forecast accuracy and switching costs. This paper explores how this trade-off influences the system’s design and operation, suggesting a nuanced view of forecasting’s role in optimization.
In principle, forecasts used in optimization can be updated every time new information becomes available, often span multiple steps ahead and are revised periodically. Transitioning from the concept of time-coupling, we speculate that switching costs not only underscore the importance of forecast accuracy but also elevate the value of forecast stability. In this context, a stable forecast with minimal variation despite periodic updates, minimizes the need for frequent policy changes, effectively reducing switching costs. Consequently, this research aims to answer the following questions:
  • Do switching costs influence the balance between forecast accuracy and stability?
  • What are the optimal strategies for forecast revision in the presence of such costs?
  • And how can integrated systems be designed to effectively balance these considerations for improved decision-making in energy management?
I-A Related work
The problem of decision making with no complete information is often called an online optimization problem. According to Powell [10], the problem involves two main challenges: modelling the sequence of decisions and designing a policy to maximize the objective. Our focus is on a type of optimization problem that depends on the current state, uses a direct lookahead approach, and includes switching costs. The exact details of this problem are given in the Problem Statement in the Section II-A.
Literature solutions for integrating forecasting and optimization are classified based on the linkage of these subproblems. There are three categories: Direct, Indirect, and Semi-direct methods.
Direct methods engage with the optimization problem during training of the forecast making it an integrated approach which is known as ‘predict and optimize’. Such approach is proposed in Elmachtoub and Grigas [11], with optimization cost integrated directly into the forecast loss function to align forecasts with end-use optimization. This optimizes the prediction model in the forecasting stage for better decision-making in the optimization stage. The computational burden is partially resolved by using a simplified loss function, however more complex problems can still pose significant computational challenges. There are also other works that design an alternative task-specific loss function, such as reported in Mandi et al. [12], and tested on a real-world energy management problem.
Indirect methods regard the two subproblems as separate, also referred in literature as ‘predict, then optimize’. This approach is regarded as standard and is widely applied in practice. In the study by Vanderschueren et al. [13], an empirical evaluation of direct and indirect approaches is conducted in the context of cost-sensitive classification. Although the study is not directly applicable to the scheduling problems, it is interesting to note that those authors find that the indirect approach where optimization is performed on the predictions of the classifier outperforms the integrated approach. Therefore, the effectiveness of cost-sensitive classification may not always align with the intuitively expected benefits of integrated training within the ’predict and optimize’.
The third category, Semi-direct methods, considers characteristics of the optimization problem but abstains from direct interaction during training of the forecast. These methods can take various forms, as potential improvements can occur at any stage outside of the training process. An example is found in the work by Kazmi and Paskevich [14], where those authors propose a Bayesian Optimization approach for integrating the downstream optimization problem into the hyperparameter tuning process of the forecast model.
There is a noticeable gap in the literature when it comes to studying operational aspects of forecasting model and optimization deployment. For forecasting, the frequency of the forecast revision can be interpreted as both the frequency of the forecast model deployment and the frequency of model training. The latter aspect is studied by Spiliotis and Petropoulos [15], where those authors investigate different scenarios of updating the model fit for univariate exponential smoothing (ES) and univariate gradient boosting models (LightGBM).
We are particularly interested in the applications of optimization with predictions in environments where switching costs are present. This topic has featured in research on online convex optimization (OCO) problems. In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm, also referred to as model predictive control (MPC). When rerun at every time step, this method commits to the first step of the future horizon while treating later decisions as advisory. The research conducted by Chen et al. [16] sheds light on the role of noisy prediction in the design of the OCO algorithm. Those authors propose a model that simulates realistic prediction errors. Using this model, they show analytically that an alternative Averaging Fixed Horizon Control (AFHC) algorithm can achieve sublinear regret and a constant competitive ratio, which indicate a desirable level of optimality in reference to the theoretical best policy. Figure 1 shows an overview of the RHC and AFHC algorithm designs, along with other algorithms discussed further. Those authors state that other algorithms can reach better performance if they can use the predictions more efficiently. A later work, Chen et al. [17], introduces a class of policies known as Committed Horizon Control (CHC), which extends the principles of Receding Horizon Control (RHC) and Adaptive Forward Horizon Control (AFHC) through the integration of a commitment concept. This concept involves specifying a predetermined number of steps that the algorithm executes. The analysis of the generalized problem reveals a trade-off set by the varying levels of switching costs and quality of the predictions. The general observation is that the level of commitment is problem-specific and depends on the nature of the switching costs and the quality of the predictions. Therefore, the results show that the commitment level should be tuned for maximizing the performance of the policy.
(a) RHC
(b) FHC
(c) AFHC
(d) CHC
Figure 1: Diagrams of the Online Optimization algorithms. RHC: Receding Horizon Control, FHC: Fixed Horizon Control, AFHC: Averaging Fixed Horizon Control, CHC: Committed Horizon Control.
Current work sets to explore the performance of commitment-based policies in the context of energy management systems. We are particularly interested in analyzing the forecast and planning stability, as a potential factor that influences the performance of the policy. A comprehensive study of forecast stability is made by Godahewa et al. [18]. Those authors categorize the stability into two types: vertical and horizontal. Vertical stability refers to the variation period-to-period variability between forecasts given for the same time period but issued from different origins. Horizontal stability describes the variation between predictions made at one instance within a prediction window. The lack of planning stability has long been recognized an important issue within research on supply chain management. The vertical stability in order forecasts is often referred to as the system nervousness. It is a well-known phenomenon that can lead to inefficient production and inventory management. The extra capacity implications are also referred to as the bullwhip effect [19]. The commonly used methods to obtain a stable plan include freezing intervals [20] [21], using smoothing techniques [22], and incorporating changing costs into the objective function [23] [24].
Freezing order plans is synonymous with the concept of commitment, discussed earlier in the context of OCO problems with switching costs. The freezing intervals are periods during which the plan is not revised, and the production is executed according to the plan. In supply chain management, the freezing is found to be effective in reducing the nervousness of the system [20]. However, it is also found that the freezing itself is a potential source of nervousness. It is shown by Tang and Grubbström [25] that forecast errors have an effect on the selection of optimal freezing intervals. Xie et al. [26] show that the interaction between forecast errors and the freezing intervals has a significant impact on the performance of the system.
Unlike supply chain management, energy management systems are more dynamic and require more frequent updates of forecasts and policies. Switching costs are also observed in energy, such as the ramping costs of generating units [7] [27] and toggling server load balancing [8] [28]. Furthermore, the literature is scarcer when considering the operational aspects of a combined forecasting and optimization problem. Guo et al. [29] review the historical evolution and state-of-the-art for the optimization methods for energy storage dispatch. They conclude that current research indicates that RHC performance relies strongly on the quality of predictions, and therefore, it may not be robust to uncertainty. Christianson et al. [30] analyze the problem of resource procurement and online economic dispatch for power system operators. They present a customized FHC algorithm — Feasible Fixed Horizon Control (FFHC) — which incorporates a regularization term to solve the multi-interval optimization problem while ensuring feasibility and robustness amidst the complexities introduced by ramping costs. A study by La Tona et al. [31] explores how the frequency of forecast revisions affects the Energy Management System’s (EMS) performance in online replanning tasks, focusing on battery control to align actual and declared consumption. Their findings suggest that more frequent forecasts enhance EMS performance, with diminishing returns at intervals of 60 minutes and shorter. The current work examines applications in energy storage optimization and the relationship between forecast and optimal commitment level.
I-B Contents and Contributions
Through a systematic review of the literature, it becomes evident that the effects of predictions on online optimization problems with switching costs are not well understood. This present research aims to address this gap by studying the impact of forecast stability and accuracy on the performance of the policy in the context of energy management systems. The principle contributions are as follows:
  • First, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs. We discuss the role of policy commitment and draw a parallel to stability. A trade off is demonstrated between forecast accuracy and switching costs, with the commitment level regulating the balance between the two.
  • We investigate the effect of forecast accuracy and stability on the downstream performance of the optimization policy. We extend the definition of forecast stability towards probabilistic forecasts and propose a metric for measuring horizontal and vertical stability of scenario sets.
  • An empirical evaluation of the stability of deterministic and probabilistic scenario sets is conducted. We use the case study of the integrated forecasting and optimization energy management problem from the Citylearn 2022 competition. Based on the case study, we analyze the relationship between forecast accuracy, stability and the performance of the policy. We show that running the policy for a longer period provides a simple mechanism to avoid switching costs while maintaining or improving the performance.
  • Lastly, the paper concludes with a discussion of the general implications of the results for the design of the energy management systems followed by suggested future research directions.
II Methodology
II-A Problem statement
This paper addresses the challenge of optimizing battery storage management within a multi-agent energy management system. The setting of the problem is first outlined in the CityLearn Challenge 2022 [32] and based on a real-world dataset from a grid-interactive community. More generally, we consider an online optimization with switching costs. The problem is solved with model predictive control (MPC) and full details of the objective score and relevant constraints are provided in Appendix A-A. The problem includes finding an optimal state-dependent policy $X_{t}^{MPC}(S_{t}|\theta^{LA})$. The decision-maker is faced with a sequence of decisions, where the decision at each step is based on the current state $S_{t}$ of the system and the direct lookahead approximation $\theta^{LA}$. Using the framework proposed by Powell [10], we use the following formulation:
$X_{t}^{MPC}(S_{t}|\theta^{LA})=\arg\min_{x_{t}}\Biggl{(}\sum_{i\in n}W_{i}\cdot\mathbb{C}(\textbf{x};\textbf{S})\Biggr{)}=$
$\arg\min_{x_{t+1},\ldots,x_{t+H}}\Biggl{(}\sum_{t^{\prime}=t+1}^{t+H}\sum_{i\in n}W_{i}\cdot\Bigl{[}\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ (1)
$+\beta\left\lVert\tilde{x}_{tt^{\prime}}-\tilde{x}_{t(t^{\prime}-1)}\right\rVert\Bigr{]}\Biggr{)}$
In this formulation:
  • $X_{t}^{MPC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharging $x_{t}^{neg}$ actions) made by the MPC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$. This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
  • $\mathbb{C}(\textbf{x};\textbf{S})$ is the cost function, which is a function of actions $\textbf{x}=x_{t},...x_{t+H}$ and states $\textbf{S}=S_{t},...S_{t+H}$.
  • $W_{s}$ Weight factor for scenario $s$, indicating the importance of different scenarios in the decision-making process. We assume that the weight factors are equal for all scenarios. In the deterministic case, the number of scenarios is equal to 1.
  • $\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ Represents the cost of the set of decisions $x_{t+1},...,x_{t+H}$ given the future forecast scenarios generated by the model $\theta^{LA}_{i,tt^{\prime}}$
[…]
Optimized battery schedules are evaluated using a set of KPIs. The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$. The optimization score is defined as:
Average $\text{Average Score}=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (2)
When grid-related Key Performance Indicators (KPIs) are included, the average includes the grid score $D$. The grid score is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$. The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting. Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, in this study we consider both optimizations with and without the grid score. The average score with the grid cost is defined as:
[…]
$\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$
All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario. The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in Appendix A-A.
II-B Optimization
The optimal policy is computed for the current state of the system and the forecast for the next $H$ time steps. An online optimization algorithm typically determines the optimal actions $x_{t+1},...,x_{t+H}$ given a prediction window of length $H$. Every $v$ time steps, the optimization is rerun to generate a new plan. This algorithm is referred to as a Fixed Horizon Control (FHC) algorithm.
In the context of the online optimization algorithms, the term “commitment” refers to the duration or number of time steps for which the optimized policy is committed to before the optimization process is rerun to generate a new plan. This concept plays a crucial role in determining how frequently the optimization is updated based on new information, accumulation of forecast errors and changes in the system’s state. An FHC algorithm with the shortest commitment is the Receding Horizon Control (RHC) algorithm. In RHC, the commitment is $v=1$, indicating that the optimization is rerun at every time step.
Similar to the optimization, the forecast is also run in a rolling origin fashion. The optimal policy is then applied to the system for the next $v_{F}$ time steps and the process is repeated every $v_{F}$ steps. In summary, both the forecast and the optimization have the same rolling origin setup and need to be updated at certain intervals while committing to $v$ steps. Ultimately, at each time step, the decision-maker is faced with a decision to either reuse the current plan or to revise it. Similarly, the forecast can be reused or revised. The resulting decision matrix is shown in Table I.
TABLE I: Decision Matrix illustrating the possible actions based on the decisions made in the forecasting and optimization stages.
Optimization Decision
Forecasting Decision Reuse Revise
Reuse Continue with current plan Evaluate
& adjust current plan
Revise Evaluate &
retain current plan
Re-evaluate both
forecasting & plan
II-C Theoretical Analysis of Performance Bounds
In this section, we analyze the upper bound on the performance of the FHC algorithm with a limited commitment $v$. We focus on the cost difference between the optimal policy and the policy produced by the FHC algorithm. Specifically, we examine the impact of the commitment $v$, which defines how often the policy and forecast are updated. Our goal is to show how the performance bound is influenced by a trade-off between forecast accuracy and switching costs.
To illustrate this, we consider two models of forecast error: independent and identically distributed (i.i.d.) noise, noise with exponentially decaying correlation and a probabilistic forecast. We start with the deterministic case (point forecast) and then extend to the probabilistic case (set of scenarios). Following the derivation in Chen et al. [17], we derive the average-case performance of the FHC algorithm by taking the expectation of the cost difference with respect to the prediction noise $(e(t))^{T}_{t=1}$.
Here, OPT represents the optimal policy that would achieve the lowest possible cost if there were perfect information and no prediction errors.
$\mathbb{E}\operatorname{cost}(FHC)\leq$ (4)
$\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2G\mathbb{E}\sum_{t=1}^{T}\left\|y_{t}-y_{t\mid t-\phi^{1}(t)}\right\|_{2}^{\alpha}$
$\mathbb{E}\operatorname{cost}(FHC)\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2GT\left\|f_{v}\right\|^{\alpha}.$ (5)
The term $\left\|f_{v}\right\|^{\alpha}$ is the $\alpha$-norm of the prediction error covariance, where $\alpha$ is the exponent of the Hölder condition. The two noise models (i.i.d. and exponentially decaying correlation) are explained below.
Independent and Identically Distributed Noise
The i.i.d. noise is a simple and understandable baseline for forecast error, without time dependency. For i.i.d. noise, we have:
$f(s)=\begin{cases}I,&s=0\\ 0,&\text{otherwise.}\end{cases}$
The statistical properties of prediction errors in this context are:
$\mathbb{E}\left[e(s)e(s)^{T}\right]=R_{e}$
$\operatorname{trace}(R_{e})=\sigma^{2}$
This leads to the upper bound on the FHC cost:
$\mathbb{E}\operatorname{cost}(FHC(v))\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2GT\sigma^{\alpha}$ (6)
The upper bound on the competitive difference depends on $v$ only in the first term (in the denominator). Therefore, the optimal cost is minimized when $v=H$, i.e., the commitment level reaches the maximum allowable duration and corresponds to the forecast horizon length. The first term represents switching costs, so a shorter commitment period leads to more frequent updates and higher switching costs. The second term, representing i.i.d. noise, is independent of $v$.
Exponentially Decaying Correlation
A more realistic model of prediction noise uses correlations that decay exponentially over time. For this, given $a<1$, the Frobenius norm of the correlation structure is:
$\|f(s)\|_{F}=\begin{cases}ca^{s},&s\geq 0\\ 0,&s<0\end{cases}$ (7)
We then express the prediction error covariance as:
$\left\|f_{v}\right\|^{2}$ $=\sum_{s=0}^{v}\operatorname{trace}(R_{e}f(s)^{T}f(s))=\sum_{s=0}^{v}\left\langle R_{e}^{1/2},f(s)\right\rangle^{2}$
$\leq\sum_{s=0}^{v}\left(\left\|R_{e}^{1/2}\right\|_{F}\|f(s)\|_{F}\right)^{2}$
$=\sum_{s=0}^{v}c^{2}\sigma^{2}a^{2s}=c^{2}\sigma^{2}\frac{1-a^{2v}}{1-a^{2}}$
Given that $h$ exhibits G-Lipschitz continuity ($\alpha=1$), the norm of the correlation function $\|f_{v}\|$ is bounded by:
$\left\|f_{v}\right\|\leq\frac{c\sigma(1-a^{2(v+1)}/2)}{1-a^{2}}$ (8)
Plugging this into the expression for the expected competitive difference of the FHC algorithm, we get:
$\mathbb{E}\operatorname{cost}(FHC(v))\leq$ (9)
$\mathbb{E}\operatorname{cost}(OPT)+2\left[\frac{T}{v}\right]\beta D+2GT\sigma^{\alpha}\frac{c(1-a^{2(v+1)})/2}{1-a^{2}}$
$\leq\mathbb{E}\operatorname{cost}(OPT)+2\left[\frac{T}{v}\right]\beta D+2GTc\sigma\left(\frac{a^{2v+2}-2}{2(a^{2}-1)}\right)$
To simplify, we group the parameters independent of $v$:
$A=2T\beta D\quad B=\frac{1}{a^{2}-1}\quad C=GTc\sigma a^{2}$
Thus, the expected competitive difference of the FHC algorithm is:
$\mathbb{E}\operatorname{cost}(FHC(v))\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{A}{v}+B(Ca^{2v}-1)$ (10)
The competitive difference behavior depends on two main components. The term $\frac{A}{v}$ decreases with $v$, while the term $B(Ca^{2v}-1)$ increases with $v$. Since $Ca^{2v}$ approaches 0 as $v$ increases, the second term is bound by $B$. Therefore, the minimal competitive difference is a trade-off between switching costs (first term) and forecast errors (second term).
Stochastic Optimization
For stochastic optimization with Fixed Horizon Control (FHC) applied across $n$ scenarios, we examine how the stochastic nature influences the competitive difference compared to an optimal solution. The expected cost for the stochastic algorithm across all scenarios is:
$\mathbb{E}\operatorname{cost}(FHC_{stochastic}(v))=\sum_{i=1}^{n}\mathbb{E}\operatorname{cost}(FHC(v,i))$ (11)
Taking the expectation of the competitive difference across all scenarios:
$\mathbb{E}\operatorname{cost}(FHC_{stochastic}(v))\leq$ (12)
$\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2G\mathbb{E}\left[\sum_{t=1}^{T}\left\|y_{t}-y_{t\mid t-\phi^{1}(t)}\right\|_{2}^{\alpha}\right]$
In the stochastic case, the forecast errors across the scenario ensemble would be lower in expectation than in any singular, deterministic scenario. This is because the stochastic approach averages out the errors across multiple scenarios, reducing the overall forecast error. Hence, the competitive difference of the FHC algorithm in the stochastic case will likely be lower, favoring a higher commitment level, as the first term (switching costs) decreases with $v$ at the same rate as in the deterministic case.
II-D Forecast
Following the solution implemented in the winning submission of the competition, the forecasts are generated with gradient boosting regression trees and the linear least squares regression model. The gradient boosting implementation is LightGBM [33]. Predictions are made for the next 24 hours with 24 independent models. The scenario sets are generated with added gaussian noise. The details of the forecast models and the generation of the scenario sets are outlined in the competition paper by Nweye et al. [32] and the code repository of the winner solution is available on GitLab [34]. The evaluation conducted in this study is available in the code repository [35].
While typical forecasting models use a building’s own historical data for training, in this case, we apply forecasts to out-of-sample buildings using a model trained on an isolated set of buildings. This setup follows the problem setting given in the Citylearn Challenge 2022 [32]. Although this train-test splitting is not typical in forecasting, the setup is still relevant to real-world scenarios, as the lack of historical data within the test subset is compensated by the availability of perfect weather forecast and similarity between the train and test sets.
II-D1 Rolling Origin Window
In practical applications, the forecast is issued periodically for a finite number of future time steps. This process is run iteratively in a rolling origin fashion, where the forecast is updated every $v_{f}$ time steps. A schematic of this process is shown in Figure 2. This approach involves continuously moving the time window forward by a certain step size after each forecasting iteration. For example, if we have a window size of 3 and a step size of 1, we would first generate the predictions and optimize for the steps 1-3, then at a second iteration, the forecasts are issued for the steps 2-4.
Figure 2: Vertical stability between rolling origin windows
in time-series forecast
The rolling horizon setup enables an update of the prediction using the most recent information. When the commitment period is shorter than the forecast horizon, the regular updates create overlaps between prediction windows, when the forecasts associated with a particular target time in the future are released in multiple revisions. The overlaps between windows with a rolling origin are shown in the Figure 2. The overlapping interval is highlighted in light grey.
II-D2 Stability
As a potential factor influencing the performance of the policy, we are interested in the stability of the forecast. The stability of the forecast is defined as the consistency of the predictions over time. Following the categorization, first proposed by Godahewa et al. [18], we distinguish vertical and horizontal stability. The difference between the two concepts is illustrated in Figure 3. We define the vertical stability as the difference between the predictions produced at different origins for the same time period, marked with the overlap area in Figure 2. If the forecast revisions are vertically stable, the corresponding adjustment costs in policy are set to be minimal. Horizontal stability characterizes the difference between predictions generated at the same origin. The horizontally stable forecasts are meant to reduce fluctuations between time steps within a prediction window.
(a) Vertical Stability
(b) Horizontal Stability
Figure 3: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions
Stability is evaluated for deterministic forecasts using the stability mean absolute change (MAC). The formulations of the stability metrics vary for vertical and horizontal types. Following the general framework of Van Belle et al. [36], we evaluate stability for deterministic forecasts. In line with the definition of sMAPC in [36], we define the absolute scale stability metric, Mean Absolute Change (MAC). The metric measures the change between the predictions between overlapping rolling windows — ${MAC}_{V}$, or the change of predictions generated at the same origin — ${MAC}_{H}$
$\operatorname{MAC}_{V}=\frac{\sum_{i=1}^{H-1}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i\mid t-1}\right|}{H-1}$ (13)
$\operatorname{MAC}_{H}=\frac{\sum_{i=2}^{H}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i-1\mid t}\right|}{H-1}$ (14)
where $\hat{y}_{t+i\mid t}$ is the forecast for time $t+i$ generated at time $t$, $H$ is the forecast horizon, and $t$ is the current time step.
II-D3 Stability of Probabilistic Forecasts
The Earth Mover’s Distance (EMD), also known as the Wasserstein distance, quantifies the minimum amount of “work” required to transform one distribution into another, making it particularly suitable for assessing the stability of scenario sets. In a stochastic optimization problem, the choice of EMD for measuring the stability of probabilistic forecasts is based on its advantageous properties for comparing empirical distributions. Firstly, EMD is a symmetric measure and holds a triangle inequality, which makes it a valid metric. Secondly, it does not require both distributions to be on the same probability space, which makes it effective even when the two sets of forecasts do not overlap—for instance, when an initial forecast predicts a sunny day and high photovoltaic generation, and an updated forecast adjusts the prediction to a cloudy day and low generation. The EMD provides insights into the extent of the adjustments between the old predictions and the new ones. Lastly, EMD takes into account the dimension of the metric space, which makes it equivalent to the mean absolute change (MAC) used for point forecasts.
Several other metrics, such as Kullback-Leibler (KL) divergence and Jensen-Shannon distance, can measure statistical distance between distributions. However, EMD is particularly suitable for comparing scenario sample sets due to the reasons outlined above. Formally, in a 1-dimensional case, where $P$ and $Q$ are empirical sets of samples $P_{1},...,P_{n}$ and $Q_{1},...,Q_{m}$, the EMD is defined as:
$\operatorname{EMD}(P,Q)=\frac{1}{n}\sum_{i=1}^{n}\left\|X_{(i)}-Y_{(i)}\right\|$ (15)
where $X_{(i)}$ and $Y_{(i)}$ are the sorted samples of $P$ and $Q$ respectively.
This metric measures the dissimilarity between two sets of scenarios. To assess the stability of forecasts, the EMD can be computed for both horizontal and vertical stability. Horizontal stability (${EMD}_{H}$) measures the similarity between predictions within the same horizon, while vertical stability (${EMD}_{V}$) evaluates the distance between predictions generated for the same period at different origins:
$EMD_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (16)
$EMD_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (17)
where $\hat{y}_{t+i\mid t,j}$ is the $j$-th scenario of the forecast for time step $t+i$ generated at time $t$, and $N$ is the number of scenarios.
II-D4 Accuracy of Determinstic Forecasts
For deterministic forecasts, the MAE is utilized to quantify the forecast accuracy. It calculates the average absolute difference between each forecasted value and the corresponding actual value. Mathematically, it is expressed as:
$\operatorname{MAE}=\Biggl{(}\frac{1}{H}\sum_{i=1}^{H}\left|y_{i}-\hat{y}_{i}\right|\Biggr{)}$ (18)
where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the forecasted value at time $i$.
II-D5 Accuracy of Probabilistic Forecasts
For probabilistic forecasts, we employ the Energy Score (ES). The ES is a multivariate generalization of the continuous ranked probability score (CRPS), a widely used metric for evaluating probabilistic forecasts. The ES measures the distance between the forecasted distribution and the actual value. The lower the ES, the better the forecast. The ES is defined as:
$ES=\left(\frac{1}{N}\sum_{j=1}^{N}\left\|y_{i}-\hat{y}_{ij}\right\|^{p}-\frac{1}{2N^{2}}\sum_{j=1}^{N}\sum_{k=1}^{N}\left\|\hat{y}_{ij}-\hat{y}_{ik}\right\|^{p}\right)^{\frac{1}{p}}$ (19)
where $y_{i}$ is the actual value, $\hat{y}_{ij}$ is the $j$-th scenario of the forecasted value at time $i$, $N$ is the number of scenarios, and $p$ is the order of the norm. The second term in the equation is the mean pairwise distance between the scenarios.
III Experimentation and Results
Compared to theoretical analysis, practical setups are more complex. They often deal with non-trivial correlation structures in forecast errors and the potential for non-convex behaviors in optimization. This complexity is due to the necessity for cooperative behavior across multiple buildings. Additionally, various temporal dependencies emerge from switching costs and system feedback. To confirm the initial theoretical findings, we proceed with evaluating the performance of the FHC algorithm with different combinations of forecast and optimization commitment in a real-world scenario.
According to the decision matrix in Figure I, at every time step the forecast and optimization plans can be updated or retained. Therefore, the forecast and optimization run with commitment periods $v_{F}$ and $v_{O}$. respectively. We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment. The result is shown in Tables II(a) and II(b). The scores in the tables indicate that updating the optimization plan more frequently than the forecast offers no added benefit. The optimal performance is achieved when the forecast and optimization are updated at the same frequency. Therefore, in the following analysis, we set $v=v_{F}=v_{O}$.
(a) FHC algorithm with deterministic forecast
$v_{O}$
$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
1 0.899
2 0.910 0.904
3 0.911 0.907 0.904
4 0.912 0.908 0.907 0.905
5 0.913 0.910 0.908 0.908 0.903
6 0.912 0.910 0.907 0.908 0.908 0.906
7 0.912 0.910 0.907 0.910 0.908 0.909 0.902
8 0.910 0.907 0.908 0.905 0.906 0.908 0.902 0.903
9 0.911 0.909 0.906 0.908 0.908 0.908 0.904 0.908 0.903
10 0.913 0.909 0.909 0.907 0.906 0.908 0.903 0.906 0.905 0.905
11 0.912 0.910 0.909 0.910 0.908 0.909 0.905 0.906 0.906 0.907 0.903
12 0.917 0.911 0.911 0.909 0.908 0.910 0.904 0.907 0.907 0.907 0.904 0.908
(b) FHC algorithm with stochastic forecast
$v_{O}$
$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
1 0.875
2 0.888 0.875
3 0.894 0.886 0.873
4 0.895 0.885 0.882 0.874
5 0.899 0.892 0.885 0.886 0.875
6 0.901 0.891 0.881 0.883 0.884 0.876
7 0.901 0.896 0.887 0.888 0.886 0.883 0.875
8 0.900 0.894 0.887 0.884 0.884 0.885 0.882 0.874
9 0.902 0.894 0.883 0.887 0.884 0.882 0.883 0.880 0.876
10 0.903 0.894 0.888 0.887 0.881 0.884 0.883 0.880 0.883 0.876
11 0.902 0.895 0.889 0.887 0.885 0.885 0.884 0.880 0.884 0.882 0.876
12 0.905 0.895 0.887 0.884 0.886 0.880 0.884 0.880 0.883 0.882 0.882 0.879
TABLE II: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12
Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability. In Figure 4, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods. The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings. Similar evaluation is done for the in-sample buildings and shown in Appendix B. It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods. The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours. The hypothesis is that this is due to the nature of the implemented method of noise simulation. The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
The plots of the optimization scores show different behavior depending on the presence of the switching costs. The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship. The optimal commitment period remains at 1 for the deterministic forecast, while the stochastic forecast enables an optimal performance at a longer commitment period.
These observations are supported by the correlation coefficients in the tables III(a) and III(b) with the correlation coefficients between the forecast properties and the optimization scores. The accuracy metrics are in positive correlation with the optimization score. When switching costs are factored in, the inverse correlation between the vertical stability and the cost becomes less pronounced. Yet, the correlation between the horizontal stability and the cost changes from negative to positive correlation when switching costs are included.
Accuracy
Optimization Score: without switching costs
Stability: vertical (top) and horizontal (bottom)
Optimization Score: with switching costs
Point Forecast + Deterministic MPC
Probabilistic forecast + Stochastic MPC
Figure 4: Accuracy (Top), Horizontal (Middle) and Vertical Stability (Bottom row) of stochastic forecast over different revision periods with data from out-of-sample buildings. The optimization score is shown with and without switching costs. The forecast error is measured with MAE and Energy Score, while the stability is measured with MAC and EMD. The optimization score is measured with the cost function of the FHC algorithm.
(a) FHC algorithm with deterministic forecast
Correlation Coef. $MAE$ $MAC_{V}$ $MAC_{H}$
$cost(FHC(v))$ 0.95 -0.94 -0.27
$cost(FHC(v))$ w/ switching costs 0.70 -0.62 0.17
(b) FHC algorithm with stochastic forecast
Correlation Coef. $ES$ $EMD_{V}$ $EMD_{H}$
$cost(FHC_{stochastic}(v))$ 0.63 -0.86 -0.29
$cost(FHC_{stochastic}(v))$ w/ switching costs 0.71 -0.06 0.47
TABLE III: Correlation Coefficients for Forecast Metrics and Optimization Score
IV Discussion
The observations from the results are consistent with the theoretical analysis of the upper bound on the competitive difference of the FHC algorithm. We observe that the optimal commitment period is dependent on the forecast properties. The inclusion of switching costs in the optimization problem changes the relationship between desired forecast properties and the optimization score. The forecast error is broadly lower with shorter commitment periods, however, the horizontal and vertical stability become also relevant. In the theoretical analysis, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs. Through testing with real data, we observe that switching costs can be partially mitigated by committing to more stable forecasts and optimization plans.
The addition of switching costs makes the algorithmic problem harder as it forces current actions to depend on beliefs about future cost functions. In Figure 5, we can see that the shape of the final load curve is significantly different when switching costs are included into the optimization. Without switching costs, the battery is charged and discharged while following the price and emissions cost signals. The resulting net load curve (in green) is compared to the perfect scenario (in yellow), where the battery is charged and discharged following the perfect forecast. In the case with switching costs, the optimal policy is regulating the net load curve to a smoother shape with less drops and peaks. The charging and discharging actions are the means of such regulation. However, due to the error in the forecast, the actual net load curve (in green) is not as smooth and flat as the optimal policy (in yellow). It generally follows the optimal policy, but with oscillation around it.
(a) Without switching costs
(b) With switching costs
Figure 5: The difference in net load between MPC with and without switching costs. Actual net load is shown in yellow. The base load is shown in green. The net load following the perfect forecast is shown in red. The electricity price is shown in blue.
Minimal commitment periods $v$ of the control algorithm, combined with imperfect predictions, can lead to a scenario where the control strategy is constantly “chasing” the optimal state based on imperfect prediction. This can aggravate the impact of switching costs, as the control actions may change frequently in response to new information, even when such changes are not cost-effective in the long run. While traditional forward-looking RHC formulations can include terms in the objective function to penalize switching costs, the longer commitment period provides a native mechanism that respects past actions of the FHC algorithm. This mechanism can reduce the need for frequent adjustments, thus avoiding some of the switching costs.
We identify two reasons for underperformance of the algorithm with a minimal commitment period (RHC, equivalent to FHC with $v_{O}=1$). Firstly, when the policy is computed for a period in a problem with switching cost, the past decisions are calculated as optimal with respect to future decisions given the forecast, however, after the policy is reviewed at the next iteration with an updated forecast, the optimality of the executed decision at the recent time step does not hold, since the future policies within the same horizon are reviewed. A simple illustrative example would be the battery is discharged first with an outlook for charging in the future when the price is more favorable. Once the forecast and policy updates, the updated policy is discharging again, however the battery has been discharged earlier. The second reason is that while the objective cost function does include the term for the switching costs $\beta||{x_{t}}-x_{t-1}||$, the switching costs are computed pairwise between loads at close time instances. However, since the total charging capacity of batteries is finite and limited, it is advantageous to regulate switching costs within longer action windows.
When prediction is inaccurate in the look-ahead window, the RHC is more effective than FHC with a commitment period. If we make predictions accurate for a small window $\gamma$ and then inaccurate for the remaining $(\omega-\gamma)$ steps of the lookahead window, the fixed horizon policy is affected by the inaccurate predictions whereas RHC only acts on the correct ones. This is typically the case for a multi-step horizon forecast when the nearest steps are predicted at a higher accuracy. The situation changes for a stochastic MPC, generally hedged better against prediction errors. We can see from the results that FHC with longer commitment period becomes viable when the optimization is computed for a set of scenarios, as opposed to a deterministic point forecast.
V Conclusion
In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems. In summary, our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and severity of policy revisions, thereby mitigating the financial and operational impacts of switching costs. Our findings have significant implications for the design of energy management systems. They suggest that enhancing the stability of forecasts can improve system performance by reducing the need for frequent adjustments and managing the trade-offs between immediate and future resource allocation more effectively. Furthermore, we propose a novel approach to evaluating the stability of probabilistic forecasts using the Earth Mover’s Distance, which provides a robust measure of the dissimilarity between scenario sets.
Looking ahead, there are questions remaining to be addressed. For future research, we recommend further exploration into methods that enhance forecast stability without compromising accuracy. It would be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading. Furthermore, from literature and current findings, there are indications that methods like AFHC and stochastic FHC, show more robustness in hedging against the forecast errors by aggregating policies with different underlying scenarios. However, they require more computational resources and are more complex to implement. An interesting avenue for future research would be defining the criteria for problems where these methods are most effective. These efforts would contribute to a deeper understanding of forecast-dependent optimization and improve the operational strategies of energy management systems facing an evolving energy landscape.
Appendix A Cost Function and Optimization
A-A Optimization score
Optimized battery schedules are evaluated using a set of KPIs, each of which is targeted for minimization. The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$. The optimization score is defined as:
Average $\text{Average Score}=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (20)
When grid-related KPIs are included, the average includes the grid score $D$. The grid score is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$. The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting. Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, in this study we consider both optimizations with and without the grid score. The average score with the grid cost is defined as:
[…]
$\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$
All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario.
The normalized electricity cost, $C$, as delineated in Eq. 22, is the ratio of electricity spending in a given policy, $c_{policy}$, to the spending in a reference scenario without battery intervention, $c_{{nobattery}}$ The cost metric $c$ is further defined in Eq. 23 as the aggregate of the non-negative product of district-level net electricity price, $E_{h}\times T_{h}(\$)$, where $E_{h}$ denotes the electricity consumption of the district at the hour $h$ and $T_{h}$ specifies the electricity rate corresponding to that hour.
$C=\frac{c_{{submission}}}{c_{{nobattery}}}$ (22)
$c=\sum_{h=0}^{n-1}\max\left(0,E_{h}\times T_{h}\right)$ (23)
Similarly, the normalized carbon emissions, $G$, is defined in Eq. 24 as the ratio of district carbon emissions for a given policy, $g_{{policy}}$, relative to the emissions in the aforementioned baseline scenario, $g_{{nobattery}}$. The emission metric $g$ is elaborated in Eq. 25 as the sum of carbon emissions, measured in $\left({kg}_{\mathrm{CO}_{2}\mathrm{e}}/\mathrm{kWh}\right)$, given by $E_{h}\times O_{h}$. Here, $O_{h}$ represents the carbon intensity for the hour $h$.
$G=\frac{g_{{submission}}}{g_{{nobattery}}}$ (24)
$g=\sum_{h=0}^{n-1}\max\left(0,E_{h}\times O_{h}\right)$ (25)
Lastly, the evaluation metric includes a grid-related KPI, $D$. The metric follows grid-level objectives, such as the minimization of ramping and load factor. It is defined as the mean of the normalized ramping KPI, $R$, and the Load Factor KPI, $(1-L)$. The formulation, scaled by the grid cost in the baseline no-battery scenario, is given in Eq. 26.
$D=\operatorname{avg}\left(\frac{R_{\text{entry}}}{R_{\text{no\ battery}}},\frac{1-L_{\text{entry}}}{1-L_{\text{no\ battery}}}\right)$ (26)
The ramping KPI, $R$, reflects the smoothness of the district’s load profile. A low $R$ indicates a gradual increase in grid electricity demand even after self-generation becomes unavailable in the evening and early morning, while a high $R$ indicates abrupt changes in load on the grid, which may lead to unscheduled strain on grid infrastructure and potential blackouts due to supply deficits. It is calculated as the sum of the absolute difference of net electricity consumption between consecutive time steps:
$R=\sum_{t=1}^{8760}|E_{t}-E_{t-1}|$ (27)
The Load Factor, $L$, indicates the efficiency of electricity consumption and is bounded between 0 (very inefficient) and 1 (highly efficient). Thus, the goal is to minimize $(1-L)$. $L$ is calculated as the average ratio of monthly average to maximum net electricity consumption:
$L=\left(\frac{1}{12}\sum_{m=0}^{11}\left(\frac{\sum_{h=0}^{729}E_{730m+h}}{730\max(E_{730m},\ldots,E_{730m+729})}\right)\right)$ (28)
A-B MPC Formulation
$\mathbf{C}(S_{t},x_{t},\theta^{LA}_{t})=\sum_{i\in{price,carbon,grid}}\mathbf{C}_{i}(\tilde{S}{tt^{\prime}},\tilde{x}_{tt^{\prime}})$ (29)
$=$ (30)
subject to
$SOC_{min}\leq SOC_{t}\leq SOC_{max},\quad\forall t,$ (31)
$x_{t}=x_{t}^{pos}+x_{t}^{neg},\quad\forall t,$ (32)
$-P_{max}\leq x_{t}^{neg}\leq 0,\quad\forall t,$ (33)
$SOC_{t}=SOC_{t-1}+$ (34)
$\eta_{charging}\cdot x_{t}^{pos}-\frac{1}{\eta_{discharging}}\cdot x_{t}^{neg},\quad\forall t.$ (35)
In this formulation:
  • $X_{t}^{MPC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the MPC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
  • $W_{s}$ Weight factor for scenario $s$, indicating the importance of different scenarios in the decision-making process. We assume that the weight factors are equal for all scenarios. In the deterministic case, the number of scenarios is equal to 1.
  • $C_{price}$, $C_{carbon}$, and $C_{grid}$ are cost functions representing the cost of electricity, the cost associated with carbon emissions, and the cost related to grid reliance, respectively. Each of these costs depends on the action taken ($x_{t}^{pos}$ and $x_{t}^{neg}$) and the current system state $S_{t}$.
  • $SOC_{t}$ denotes the state of charge of the battery at time $t$, with $SOC_{min}$ and $SOC_{max}$ being the minimum and maximum allowable states of charge, respectively.
  • $P_{max}$ is the maximum power with which the battery can be charged or discharged.
  • $\eta_{charging}$ and $\eta_{discharging}$ are the charging and discharging efficiencies of the battery, respectively.
Appendix B In-sample Performance
(a) Performance of the FHC algorithm with deterministic forecast. Study over combinations of forecast and optimization commitment between 1 and 12
$v_{O}$
$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
1 0.486
2 0.489 0.489
3 0.491 0.492 0.491
4 0.492 0.492 0.494 0.492
5 0.493 0.494 0.494 0.494 0.493
6 0.493 0.493 0.493 0.495 0.495 0.493
7 0.493 0.494 0.495 0.495 0.495 0.495 0.493
8 0.493 0.493 0.495 0.493 0.495 0.496 0.495 0.493
9 0.494 0.495 0.494 0.495 0.495 0.495 0.495 0.496 0.494
10 0.494 0.494 0.495 0.495 0.494 0.495 0.496 0.495 0.496 0.494
11 0.494 0.495 0.495 0.495 0.495 0.495 0.495 0.495 0.496 0.496 0.494
12 0.495 0.495 0.495 0.495 0.495 0.495 0.496 0.496 0.495 0.496 0.496 0.495
(b) Performance of the FHC algorithm with stochastic forecast. Study over combinations of forecast and optimization commitment between 1 and 12
$v_{O}$
$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
1 0.798
2 0.805 0.800
3 0.809 0.806 0.803
4 0.809 0.805 0.807 0.802
5 0.811 0.807 0.806 0.805 0.805
6 0.812 0.807 0.805 0.805 0.806 0.805
7 0.810 0.808 0.807 0.805 0.806 0.808 0.805
8 0.808 0.806 0.808 0.802 0.807 0.807 0.807 0.804
9 0.811 0.809 0.806 0.805 0.806 0.807 0.808 0.806 0.805
10 0.811 0.807 0.807 0.805 0.806 0.807 0.808 0.806 0.808 0.806
11 0.810 0.808 0.808 0.805 0.807 0.808 0.807 0.807 0.808 0.808 0.806
12 0.813 0.809 0.807 0.805 0.806 0.808 0.807 0.806 0.807 0.808 0.808 0.807
TABLE IV: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12 (with data from in-sample buildings)
Accuracy
Optimization Score: without switching costs
Stability: vertical (top) and horizontal (bottom)
Optimization Score: with switching costs
Point Forecast + Deterministic MPC
Probabilistic forecast + Stochastic MPC
Figure 6: Accuracy (Top), Horizontal (Middle) and Vertical Stability (Bottom row) of stochastic forecast over different revision periods with data from in-sample buildings

### v2

@@
 This reality necessitates making decisions with an incomplete knowledge of the future and for a limited time window, a scenario frequently approached through the lens of online optimization.
-In many applications, the future is reasonably predictable.
-Indeed, it is possible to make predictions about future cost function with a significant accuracy, particularly near future.
 To effectively manage this, it is useful to combine the forecast of the problem environment with the optimization of the decision-making process.
@@
 I-A Related work
-The problem of decision making with no complete information is often called an online optimization problem.
-According to Powell [10], the problem involves two main challenges: modelling the sequence of decisions and designing a policy to maximize the objective.
-Our focus is on a type of optimization problem that depends on the current state, uses a direct lookahead approach, and includes switching costs.
-The exact details of this problem are given in the Problem Statement in the Section II-A.
 Literature solutions for integrating forecasting and optimization are classified based on the linkage of these subproblems. There are three categories: Direct, Indirect, and Semi-direct methods.
@@
 The extra capacity implications are also referred to as the bullwhip effect [18].
-The commonly used methods to obtain a stable plan include freezing intervals [20] [21], using smoothing techniques [22], and incorporating changing costs into the objective function [23] [24].
-Freezing order plans is synonymous with the concept of commitment, discussed earlier in the context of OCO problems with switching costs.
-The freezing intervals are periods during which the plan is not revised, and the production is executed according to the plan.
-In supply chain management, the freezing is found to be effective in reducing the nervousness of the system [20].
-However, it is also found that the freezing itself is a potential source of nervousness.
-It is shown by Tang and Grubbström [25] that forecast errors have an effect on the selection of optimal freezing intervals.
-Xie et al.
-[26] show that the interaction between forecast errors and the freezing intervals has a significant impact on the performance of the system.
+The commonly used methods to obtain a stable plan include increased commitment, referred as freezing intervals [19] [20], using smoothing techniques [21], and incorporating changing costs into the objective function [22] [23].
 Unlike supply chain management, energy management systems are more dynamic and require more frequent updates of forecasts and policies.
@@
 The principle contributions are as follows:
-  • First, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs.
-We discuss the role of policy commitment and draw a parallel to stability.
-A trade off is demonstrated between forecast accuracy and switching costs, with the commitment level regulating the balance between the two.
+  • First, we formally define the problem and the metric used for evaluating the downstream performance, both with and without switching costs.
+We also conduct a theoretical analysis of the performance bounds of the Fixed Horizon Control (FHC) algorithm, highlighting how the commitment level influences the trade-off between forecast accuracy and switching costs.
   • We investigate the effect of forecast accuracy and stability on the downstream performance of the optimization policy.
@@
 More generally, we consider an online optimization with switching costs.
-The problem is solved with model predictive control (MPC) and full details of the objective score and relevant constraints are provided in Appendix A-A.
+The problem is solved with model predictive control (MPC) and full details of the objective score and relevant constraints are provided in Online Appendix A [^1] Appendix A.
+[Online] Available: https://arxiv.org/abs/2407.03368.
 The problem includes finding an optimal state-dependent policy $X_{t}^{MPC}(S_{t}|\theta^{LA})$.
@@
 $\arg\min_{x_{t+1},\ldots,x_{t+H}}\Biggl{(}\sum_{t^{\prime}=t+1}^{t+H}\sum_{i\in n}W_{i}\cdot\Bigl{[}\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ (1)
-$+\beta\left\lVert\tilde{x}_{tt^{\prime}}-\tilde{x}_{t(t^{\prime}-1)}\right\rVert\Bigr{]}\Biggr{)}$
+$+\beta\left\lVert\tilde{y}_{tt^{\prime}}-\tilde{y}_{t(t^{\prime}-1)}\right\rVert\Bigr{]}\Biggr{)}$
 In this formulation:
@@
 All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario.
-The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in Appendix A-A.
+The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in Online Appendix A  [^2] Appendix A.
+[Online] Available: https://arxiv.org/abs/2407.03368.
+Furthermore, In the Online Appendix C  [^3] Appendix C.
+[Online] Available: https://arxiv.org/abs/2407.03368, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs.
+We discuss the role of policy commitment and draw a parallel to stability.
+A trade off is demonstrated between forecast accuracy and switching costs, with the commitment level regulating the balance between the two.
 II-B Optimization
@@
 forecasting & plan
-II-C Theoretical Analysis of Performance Bounds
+II-C Forecast
+Following the solution implemented in the winning submission of the competition, the forecasts are generated with gradient boosting regression trees and the linear least squares regression model.
+The gradient boosting implementation is LightGBM [31].
+Predictions are made for the next 24 hours with 24 independent models.
+The scenario sets are generated with added gaussian noise.
+The details of the forecast models and the generation of the scenario sets are outlined in the competition paper by Nweye et al.
+[29] and the code repository of the winner solution is available on GitLab [32].
+The evaluation conducted in this study is available in the code repository [33].
+While typical forecasting models use a building’s own historical data for training, in this case, we apply forecasts to out-of-sample buildings using a model trained on an isolated set of buildings.
+This setup follows the problem setting given in the Citylearn Challenge 2022 [29].
+Although this train-test splitting is not typical in applied tasks, the setup is still relevant to real-world scenarios, as the lack of historical data within the test subset is compensated by the availability of perfect weather forecast and similarity between the train and test sets.
+II-C1 Rolling Origin Window
+In practical applications, the forecast is issued periodically for a finite number of future time steps.
+This process is run iteratively in a rolling origin fashion, where the forecast is updated every $v_{f}$ time steps.
+A schematic of this process is shown in Figure 2.
+This approach involves continuously moving the time window forward by a certain step size after each forecasting iteration.
+For example, if we have a window size of 3 and a step size of 1, we would first generate the predictions and optimize for the steps 1-3, then at a second iteration, the forecasts are issued for the steps 2-4.
+Figure 2: Vertical stability between rolling origin windows
+in time-series forecast
+The rolling horizon setup enables an update of the prediction using the most recent information.
+When the commitment period is shorter than the forecast horizon, the regular updates create overlaps between prediction windows, when the forecasts associated with a particular target time in the future are released in multiple revisions.
+The overlaps between windows with a rolling origin are shown in the Figure 2.
+The overlapping interval is highlighted in light grey.
+II-C2 Stability
+As a potential factor influencing the performance of the policy, we are interested in the stability of the forecast.
+The stability of the forecast is defined as the consistency of the predictions over time.
+Following the categorization, first proposed by Godahewa et al.
+[17], we distinguish vertical and horizontal stability.
+The difference between the two concepts is illustrated in Figure 3.
+We define the vertical stability as the difference between the predictions produced at different origins for the same time period, marked with the overlap area in Figure 2.
+If the forecast revisions are vertically stable, the corresponding adjustment costs in policy are set to be minimal.
+Horizontal stability characterizes the difference between predictions generated at the same origin.
+The horizontally stable forecasts are meant to reduce fluctuations between time steps within a prediction window.
+(a) Vertical Stability
+(b) Horizontal Stability
+Figure 3: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions
+Stability is evaluated for deterministic forecasts using the stability mean absolute change (MAC).
+The formulations of the stability metrics vary for vertical and horizontal types.
+In line with the definition of sMAPC in [34], we define the absolute scale stability metric, Mean Absolute Change (MAC).
+The metric measures the change between the predictions between overlapping rolling windows — ${MAC}_{V}$, or the change of predictions generated at the same origin — ${MAC}_{H}$
+$\operatorname{MAC}_{V}=\frac{\sum_{i=1}^{H-1}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i\mid t-1}\right|}{H-1}$ (4)
+$\operatorname{MAC}_{H}=\frac{\sum_{i=2}^{H}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i-1\mid t}\right|}{H-1}$ (5)
+where $\hat{y}_{t+i\mid t}$ is the forecast for time $t+i$ generated at time $t$, $H$ is the forecast horizon, and $t$ is the current time step.
+II-C3 Stability of Probabilistic Forecasts
+The Earth Mover’s Distance (EMD), or Wasserstein distance, quantifies the minimum ”work” required to transform one distribution into another, making it ideal for assessing the stability of scenario sets in stochastic optimization.
+EMD is chosen for its symmetric nature, adherence to the triangle inequality, and ability to compare distributions on different probability spaces, making it effective even when forecasts differ significantly.
+It also accounts for the metric space dimension, akin to the mean absolute change (MAC) used for point forecasts.
+Other metrics like Kullback-Leibler (KL) divergence and Jensen-Shannon distance measure statistical distance, but EMD is preferred for scenario sample sets.
+In a 1-dimensional case with empirical sets $P$ and $Q$, EMD is defined as:
+$\operatorname{EMD}(P,Q)=\frac{1}{n}\sum_{i=1}^{n}\left\|P_{(i)}-Q_{(i)}\right\|$ (6)
+where $P_{(i)}$ and $Q_{(i)}$ are the sorted samples of $P$ and $Q$ respectively.
+This metric measures the dissimilarity between two sets of scenarios.
+To assess the stability of forecasts, the EMD can be computed for both horizontal and vertical stability:
+$EMD_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (7)
+$EMD_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (8)
+where $\hat{y}_{t+i\mid t,j}$ is the $j$-th scenario of the forecast for time step $t+i$ generated at time $t$, and $N$ is the number of scenarios.
+II-C4 Accuracy of Determinstic Forecasts
+For deterministic forecasts, the MAE is utilized to quantify the forecast accuracy.
+It calculates the average absolute difference between each forecasted value and the corresponding actual value.
+Mathematically, it is expressed as:
+$\operatorname{MAE}=\Biggl{(}\frac{1}{H}\sum_{i=1}^{H}\left|y_{i}-\hat{y}_{i}\right|\Biggr{)}$ (9)
+where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the forecasted value at time $i$.
+II-C5 Accuracy of Probabilistic Forecasts
+For probabilistic forecasts, we employ the Energy Score (ES).
+The ES is a multivariate generalization of the continuous ranked probability score (CRPS), a widely used metric for evaluating probabilistic forecasts.
+The ES measures the distance between the forecasted distribution and the actual value.
+The lower the ES, the better the forecast.
+The ES is defined as:
+$ES=\left(\frac{1}{N}\sum_{j=1}^{N}\left\|y_{i}-\hat{y}_{ij}\right\|^{p}-\frac{1}{2N^{2}}\sum_{j=1}^{N}\sum_{k=1}^{N}\left\|\hat{y}_{ij}-\hat{y}_{ik}\right\|^{p}\right)^{\frac{1}{p}}$ (10)
+where $y_{i}$ is the actual value, $\hat{y}_{ij}$ is the $j$-th scenario of the forecasted value at time $i$, $N$ is the number of scenarios, and $p$ is the order of the norm.
+The second term in the equation is the mean pairwise distance between the scenarios.
+III Experimentation and Results
+Compared to theoretical analysis, practical setups are more complex.
+They often deal with non-trivial correlation structures in forecast errors and the potential for non-convex behaviors in optimization.
+This complexity is due to the necessity for cooperative behavior across multiple buildings.
+Additionally, various temporal dependencies emerge from switching costs and system feedback.
+To confirm the initial theoretical findings, we proceed with evaluating the performance of the FHC algorithm with different combinations of forecast and optimization commitment in a real-world scenario.
+According to the decision matrix in Figure I, at every time step the forecast and optimization plans can be updated or retained.
+Therefore, the forecast and optimization run with commitment periods $v_{F}$ and $v_{O}$.
+respectively.
+We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment.
+The result is shown in Tables II(a) and II(b).
+The scores in the tables indicate that updating the optimization plan more frequently than the forecast offers no added benefit.
+The optimal performance is achieved when the forecast and optimization are updated at the same frequency.
+Therefore, in the following analysis, we set $v=v_{F}=v_{O}$.
+(a) FHC algorithm with deterministic forecast
+$v_{O}$
+$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
+1 0.899
+2 0.910 0.904
+3 0.911 0.907 0.904
+4 0.912 0.908 0.907 0.905
+5 0.913 0.910 0.908 0.908 0.903
+6 0.912 0.910 0.907 0.908 0.908 0.906
+7 0.912 0.910 0.907 0.910 0.908 0.909 0.902
+8 0.910 0.907 0.908 0.905 0.906 0.908 0.902 0.903
+9 0.911 0.909 0.906 0.908 0.908 0.908 0.904 0.908 0.903
+10 0.913 0.909 0.909 0.907 0.906 0.908 0.903 0.906 0.905 0.905
+11 0.912 0.910 0.909 0.910 0.908 0.909 0.905 0.906 0.906 0.907 0.903
+12 0.917 0.911 0.911 0.909 0.908 0.910 0.904 0.907 0.907 0.907 0.904 0.908
+(b) FHC algorithm with stochastic forecast
+$v_{O}$
+$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
+1 0.875
+2 0.888 0.875
+3 0.894 0.886 0.873
+4 0.895 0.885 0.882 0.874
+5 0.899 0.892 0.885 0.886 0.875
+6 0.901 0.891 0.881 0.883 0.884 0.876
+7 0.901 0.896 0.887 0.888 0.886 0.883 0.875
+8 0.900 0.894 0.887 0.884 0.884 0.885 0.882 0.874
+9 0.902 0.894 0.883 0.887 0.884 0.882 0.883 0.880 0.876
+10 0.903 0.894 0.888 0.887 0.881 0.884 0.883 0.880 0.883 0.876
+11 0.902 0.895 0.889 0.887 0.885 0.885 0.884 0.880 0.884 0.882 0.876
+12 0.905 0.895 0.887 0.884 0.886 0.880 0.884 0.880 0.883 0.882 0.882 0.879
+TABLE II: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12
+Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability.
+In Figure 4, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods.
+The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings.
+Similar evaluation is done for the in-sample buildings and shown in Online Appendix D  [^4] Appendix D.
+[Online] Available: https://arxiv.org/abs/2407.03368.
+It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods.
+The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
+For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours.
+The hypothesis is that this is due to the nature of the implemented method of noise simulation.
+The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
+The plots of the optimization scores show different behavior depending on the presence of the switching costs.
+The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship.
+The optimal commitment period remains at 1 for the deterministic forecast, while the stochastic forecast enables an optimal performance at a longer commitment period.
+These observations are supported by the correlation coefficients in the tables III(a) and III(b) with the correlation coefficients between the forecast properties and the optimization scores.
+The accuracy metrics are in positive correlation with the optimization score.
+When switching costs are factored in, the inverse correlation between the vertical stability and the cost becomes less pronounced.
+Yet, the correlation between the horizontal stability and the cost changes from negative to positive correlation when switching costs are included.
+Accuracy
+Stability: vertical (top) and horizontal (bottom)
+Optimization Score: without switching costs
+Optimization Score: with switching costs
+Point Forecast + Deterministic MPC
+Probabilistic forecast + Stochastic MPC
+Figure 4: Accuracy (Top), Horizontal (Second) and Vertical Stability (Third row) of stochastic forecast over different revision periods with data from out-of-sample buildings.
+The optimization score is shown with and without switching costs.
+The forecast error is measured with MAE and Energy Score, while the stability is measured with MAC and EMD.
+The optimization score is measured with the cost function of the FHC algorithm.
+(a) FHC algorithm with deterministic forecast
+Correlation Coef. $MAE$ $MAC_{V}$ $MAC_{H}$
+$cost(FHC(v))$ 0.95 -0.94 -0.27
+$cost(FHC(v))$ w/ switching costs 0.70 -0.62 0.17
+(b) FHC algorithm with stochastic forecast
+Correlation Coef. $ES$ $EMD_{V}$ $EMD_{H}$
+$cost(FHC_{stochastic}(v))$ 0.63 -0.86 -0.29
+$cost(FHC_{stochastic}(v))$ w/ switching costs 0.71 -0.06 0.47
+TABLE III: Correlation Coefficients for Forecast Metrics and Optimization Score
+IV Discussion
+The observations from the results are consistent with the theoretical analysis of the upper bound on the competitive difference of the FHC algorithm.
+We observe that the optimal commitment period is dependent on the forecast properties.
+The inclusion of switching costs in the optimization problem changes the relationship between desired forecast properties and the optimization score.
+The forecast error is broadly lower with shorter commitment periods, however, the horizontal and vertical stability become also relevant.
+In the theoretical analysis in Online Appendix C, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs.
+Through testing with real data, we observe that switching costs can be partially mitigated by committing to more stable forecasts and optimization plans.
+The addition of switching costs makes the algorithmic problem harder as it forces current actions to depend on beliefs about future cost functions.
+In Figure 5, we can see that the shape of the final load curve is significantly different when switching costs are included into the optimization.
+Without switching costs, the battery is charged and discharged while following the price and emissions cost signals.
+The resulting net load curve (in green) is compared to the perfect scenario (in yellow), where the battery is charged and discharged following the perfect forecast.
+In the case with switching costs, the optimal policy is regulating the net load curve to a smoother shape with less drops and peaks.
+The charging and discharging actions are the means of such regulation.
+However, due to the error in the forecast, the actual net load curve (in green) is not as smooth and flat as the optimal policy (in yellow).
+It generally follows the optimal policy, but with oscillation around it.
+(a) Without switching costs
+(b) With switching costs
+Figure 5: The difference in net load between MPC with and without switching costs.
+Actual net load is shown in yellow.
+The base load is shown in green.
+The net load following the perfect forecast is shown in red.
+The electricity price is shown in blue.
+Minimal commitment periods $v$ of the control algorithm, combined with imperfect predictions, can lead to a scenario where the control strategy is constantly “chasing” the optimal state based on imperfect prediction.
+This can aggravate the impact of switching costs, as the control actions may change frequently in response to new information, even when such changes are not cost-effective in the long run.
+While traditional forward-looking RHC formulations can include terms in the objective function to penalize switching costs, the longer commitment period provides a native mechanism that respects past actions of the FHC algorithm.
+This mechanism can reduce the need for frequent adjustments, thus avoiding some of the switching costs.
+We identify two reasons for underperformance of the algorithm with a minimal commitment period (RHC, equivalent to FHC with $v_{O}=1$).
+Firstly, when the policy is computed for a period in a problem with switching cost, the past decisions are calculated as optimal with respect to future decisions given the forecast, however, after the policy is reviewed at the next iteration with an updated forecast, the optimality of the executed decision at the recent time step does not hold, since the future policies within the same horizon are reviewed.
+A simple illustrative example would be the battery is discharged first with an outlook for charging in the future when the price is more favorable.
+Once the forecast and policy updates, the updated policy is discharging again, however the battery has been discharged earlier.
+The second reason is that while the objective cost function does include the term for the switching costs $\beta||{x_{t}}-x_{t-1}||$, the switching costs are computed pairwise between loads at close time instances.
+However, since the total charging capacity of batteries is finite and limited, it is advantageous to regulate switching costs within longer action windows.
+When prediction is inaccurate in the look-ahead window, the RHC is more effective than FHC with a commitment period.
+If we make predictions accurate for a small window $\gamma$ and then inaccurate for the remaining $(\omega-\gamma)$ steps of the lookahead window, the fixed horizon policy is affected by the inaccurate predictions whereas RHC only acts on the correct ones.
+This is typically the case for a multi-step horizon forecast when the nearest steps are predicted at a higher accuracy.
+The situation changes for a stochastic MPC, generally hedged better against prediction errors.
+We can see from the results that FHC with longer commitment period becomes viable when the optimization is computed for a set of scenarios, as opposed to a deterministic point forecast.
+V Conclusion
+In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems.
+In summary, our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and severity of policy revisions, thereby mitigating the financial and operational impacts of switching costs.
+Our findings have significant implications for the design of energy management systems.
+They suggest that enhancing the stability of forecasts can improve system performance by reducing the need for frequent adjustments and managing the trade-offs between immediate and future resource allocation more effectively.
+Furthermore, we propose a novel approach to evaluating the stability of probabilistic forecasts using the Earth Mover’s Distance, which provides a robust measure of the dissimilarity between scenario sets.
+Looking ahead, there are questions remaining to be addressed.
+For future research, we recommend further exploration into methods that enhance forecast stability without compromising accuracy.
+It would be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading.
+Furthermore, from literature and current findings, there are indications that methods like AFHC and stochastic FHC, show more robustness in hedging against the forecast errors by aggregating policies with different underlying scenarios.
+However, they require more computational resources and are more complex to implement.
+An interesting avenue for future research would be defining the criteria for problems where these methods are most effective.
+These efforts would contribute to a deeper understanding of forecast-dependent optimization and improve the operational strategies of energy management systems facing an evolving energy landscape.
+Appendix A Cost Function and Optimization
+A-A Optimization score
+Optimized battery schedules are evaluated using a set of KPIs, each of which is targeted for minimization.
+The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$.
+The optimization score is defined as:
+Average $\text{Average Score}=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (11)
+When grid-related KPIs are included, the average includes the grid score $D$.
+The grid score is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$.
+The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting.
+Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, in this study we consider both optimizations with and without the grid score.
+The average score with the grid cost is defined as:
+[…]
+$\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$
+All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario.
+The normalized electricity cost, $C$, as delineated in Eq.
+13, is the ratio of electricity spending in a given policy, $c_{policy}$, to the spending in a reference scenario without battery intervention, $c_{{nobattery}}$ The cost metric $c$ is further defined in Eq.
+14 as the aggregate of the non-negative product of district-level net electricity price, $E_{h}\times T_{h}(\$)$, where $E_{h}$ denotes the electricity consumption of the district at the hour $h$ and $T_{h}$ specifies the electricity rate corresponding to that hour.
+$C=\frac{c_{{submission}}}{c_{{nobattery}}}$ (13)
+$c=\sum_{h=0}^{n-1}\max\left(0,E_{h}\times T_{h}\right)$ (14)
+Similarly, the normalized carbon emissions, $G$, is defined in Eq.
+15 as the ratio of district carbon emissions for a given policy, $g_{{policy}}$, relative to the emissions in the aforementioned baseline scenario, $g_{{nobattery}}$.
+The emission metric $g$ is elaborated in Eq.
+16 as the sum of carbon emissions, measured in $\left({kg}_{\mathrm{CO}_{2}\mathrm{e}}/\mathrm{kWh}\right)$, given by $E_{h}\times O_{h}$.
+Here, $O_{h}$ represents the carbon intensity for the hour $h$.
+$G=\frac{g_{{submission}}}{g_{{nobattery}}}$ (15)
+$g=\sum_{h=0}^{n-1}\max\left(0,E_{h}\times O_{h}\right)$ (16)
+Lastly, the evaluation metric includes a grid-related KPI, $D$.
+The metric follows grid-level objectives, such as the minimization of ramping and load factor.
+It is defined as the mean of the normalized ramping KPI, $R$, and the Load Factor KPI, $(1-L)$.
+The formulation, scaled by the grid cost in the baseline no-battery scenario, is given in Eq.
+17.
+$D=\operatorname{avg}\left(\frac{R_{\text{entry}}}{R_{\text{no\ battery}}},\frac{1-L_{\text{entry}}}{1-L_{\text{no\ battery}}}\right)$ (17)
+The ramping KPI, $R$, reflects the smoothness of the district’s load profile.
+A low $R$ indicates a gradual increase in grid electricity demand even after self-generation becomes unavailable in the evening and early morning, while a high $R$ indicates abrupt changes in load on the grid, which may lead to unscheduled strain on grid infrastructure and potential blackouts due to supply deficits.
+It is calculated as the sum of the absolute difference of net electricity consumption between consecutive time steps:
+$R=\sum_{t=1}^{8760}|E_{t}-E_{t-1}|$ (18)
+The Load Factor, $L$, indicates the efficiency of electricity consumption and is bounded between 0 (very inefficient) and 1 (highly efficient).
+Thus, the goal is to minimize $(1-L)$.
+$L$ is calculated as the average ratio of monthly average to maximum net electricity consumption:
+$L=\left(\frac{1}{12}\sum_{m=0}^{11}\left(\frac{\sum_{h=0}^{729}E_{730m+h}}{730\max(E_{730m},\ldots,E_{730m+729})}\right)\right)$ (19)
+A-B MPC Formulation
+$\mathbf{C}(S_{t},x_{t},\theta^{LA}_{t})=\sum_{i\in{price,carbon,grid}}\mathbf{C}_{i}(\tilde{S}{tt^{\prime}},\tilde{x}_{tt^{\prime}})$ (20)
+$=$ (21)
+subject to
+$SOC_{min}\leq SOC_{t}\leq SOC_{max},\quad\forall t,$ (22)
+$x_{t}=x_{t}^{pos}+x_{t}^{neg},\quad\forall t,$ (23)
+$-P_{max}\leq x_{t}^{neg}\leq 0,\quad\forall t,$ (24)
+$SOC_{t}=SOC_{t-1}+$ (25)
+$\eta_{charging}\cdot x_{t}^{pos}-\frac{1}{\eta_{discharging}}\cdot x_{t}^{neg},\quad\forall t.$ (26)
+In this formulation:
+  • $X_{t}^{MPC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the MPC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
+  • $W_{s}$ Weight factor for scenario $s$, indicating the importance of different scenarios in the decision-making process.
+We assume that the weight factors are equal for all scenarios.
+In the deterministic case, the number of scenarios is equal to 1.
+  • $C_{price}$, $C_{carbon}$, and $C_{grid}$ are cost functions representing the cost of electricity, the cost associated with carbon emissions, and the cost related to grid reliance, respectively.
+Each of these costs depends on the action taken ($x_{t}^{pos}$ and $x_{t}^{neg}$) and the current system state $S_{t}$.
+  • $SOC_{t}$ denotes the state of charge of the battery at time $t$, with $SOC_{min}$ and $SOC_{max}$ being the minimum and maximum allowable states of charge, respectively.
+  • $P_{max}$ is the maximum power with which the battery can be charged or discharged.
+  • $\eta_{charging}$ and $\eta_{discharging}$ are the charging and discharging efficiencies of the battery, respectively.
+A-C Theoretical Analysis of Performance Bounds
 In this section, we analyze the upper bound on the performance of the FHC algorithm with a limited commitment $v$.
@@
 Hence, the competitive difference of the FHC algorithm in the stochastic case will likely be lower, favoring a higher commitment level, as the first term (switching costs) decreases with $v$ at the same rate as in the deterministic case.
-II-D Forecast
-Following the solution implemented in the winning submission of the competition, the forecasts are generated with gradient boosting regression trees and the linear least squares regression model.
-The gradient boosting implementation is LightGBM [33].
-Predictions are made for the next 24 hours with 24 independent models.
-The scenario sets are generated with added gaussian noise.
-The details of the forecast models and the generation of the scenario sets are outlined in the competition paper by Nweye et al.
-[32] and the code repository of the winner solution is available on GitLab [34].
-The evaluation conducted in this study is available in the code repository [35].
-While typical forecasting models use a building’s own historical data for training, in this case, we apply forecasts to out-of-sample buildings using a model trained on an isolated set of buildings.
-This setup follows the problem setting given in the Citylearn Challenge 2022 [32].
-Although this train-test splitting is not typical in forecasting, the setup is still relevant to real-world scenarios, as the lack of historical data within the test subset is compensated by the availability of perfect weather forecast and similarity between the train and test sets.
-II-D1 Rolling Origin Window
-In practical applications, the forecast is issued periodically for a finite number of future time steps.
-This process is run iteratively in a rolling origin fashion, where the forecast is updated every $v_{f}$ time steps.
-A schematic of this process is shown in Figure 2.
-This approach involves continuously moving the time window forward by a certain step size after each forecasting iteration.
-For example, if we have a window size of 3 and a step size of 1, we would first generate the predictions and optimize for the steps 1-3, then at a second iteration, the forecasts are issued for the steps 2-4.
-Figure 2: Vertical stability between rolling origin windows
-in time-series forecast
-The rolling horizon setup enables an update of the prediction using the most recent information.
-When the commitment period is shorter than the forecast horizon, the regular updates create overlaps between prediction windows, when the forecasts associated with a particular target time in the future are released in multiple revisions.
-The overlaps between windows with a rolling origin are shown in the Figure 2.
-The overlapping interval is highlighted in light grey.
-II-D2 Stability
-As a potential factor influencing the performance of the policy, we are interested in the stability of the forecast.
-The stability of the forecast is defined as the consistency of the predictions over time.
-Following the categorization, first proposed by Godahewa et al.
-[18], we distinguish vertical and horizontal stability.
-The difference between the two concepts is illustrated in Figure 3.
-We define the vertical stability as the difference between the predictions produced at different origins for the same time period, marked with the overlap area in Figure 2.
-If the forecast revisions are vertically stable, the corresponding adjustment costs in policy are set to be minimal.
-Horizontal stability characterizes the difference between predictions generated at the same origin.
-The horizontally stable forecasts are meant to reduce fluctuations between time steps within a prediction window.
-(a) Vertical Stability
-(b) Horizontal Stability
-Figure 3: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions
-Stability is evaluated for deterministic forecasts using the stability mean absolute change (MAC).
-The formulations of the stability metrics vary for vertical and horizontal types.
-Following the general framework of Van Belle et al.
-[36], we evaluate stability for deterministic forecasts.
-In line with the definition of sMAPC in [36], we define the absolute scale stability metric, Mean Absolute Change (MAC).
-The metric measures the change between the predictions between overlapping rolling windows — ${MAC}_{V}$, or the change of predictions generated at the same origin — ${MAC}_{H}$
-$\operatorname{MAC}_{V}=\frac{\sum_{i=1}^{H-1}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i\mid t-1}\right|}{H-1}$ (13)
-$\operatorname{MAC}_{H}=\frac{\sum_{i=2}^{H}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i-1\mid t}\right|}{H-1}$ (14)
-where $\hat{y}_{t+i\mid t}$ is the forecast for time $t+i$ generated at time $t$, $H$ is the forecast horizon, and $t$ is the current time step.
-II-D3 Stability of Probabilistic Forecasts
-The Earth Mover’s Distance (EMD), also known as the Wasserstein distance, quantifies the minimum amount of “work” required to transform one distribution into another, making it particularly suitable for assessing the stability of scenario sets.
-In a stochastic optimization problem, the choice of EMD for measuring the stability of probabilistic forecasts is based on its advantageous properties for comparing empirical distributions.
-Firstly, EMD is a symmetric measure and holds a triangle inequality, which makes it a valid metric.
-Secondly, it does not require both distributions to be on the same probability space, which makes it effective even when the two sets of forecasts do not overlap—for instance, when an initial forecast predicts a sunny day and high photovoltaic generation, and an updated forecast adjusts the prediction to a cloudy day and low generation.
-The EMD provides insights into the extent of the adjustments between the old predictions and the new ones.
-Lastly, EMD takes into account the dimension of the metric space, which makes it equivalent to the mean absolute change (MAC) used for point forecasts.
-Several other metrics, such as Kullback-Leibler (KL) divergence and Jensen-Shannon distance, can measure statistical distance between distributions.
-However, EMD is particularly suitable for comparing scenario sample sets due to the reasons outlined above.
-Formally, in a 1-dimensional case, where $P$ and $Q$ are empirical sets of samples $P_{1},...,P_{n}$ and $Q_{1},...,Q_{m}$, the EMD is defined as:
-$\operatorname{EMD}(P,Q)=\frac{1}{n}\sum_{i=1}^{n}\left\|X_{(i)}-Y_{(i)}\right\|$ (15)
-where $X_{(i)}$ and $Y_{(i)}$ are the sorted samples of $P$ and $Q$ respectively.
-This metric measures the dissimilarity between two sets of scenarios.
-To assess the stability of forecasts, the EMD can be computed for both horizontal and vertical stability.
-Horizontal stability (${EMD}_{H}$) measures the similarity between predictions within the same horizon, while vertical stability (${EMD}_{V}$) evaluates the distance between predictions generated for the same period at different origins:
-$EMD_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (16)
-$EMD_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (17)
-where $\hat{y}_{t+i\mid t,j}$ is the $j$-th scenario of the forecast for time step $t+i$ generated at time $t$, and $N$ is the number of scenarios.
-II-D4 Accuracy of Determinstic Forecasts
-For deterministic forecasts, the MAE is utilized to quantify the forecast accuracy.
-It calculates the average absolute difference between each forecasted value and the corresponding actual value.
-Mathematically, it is expressed as:
-$\operatorname{MAE}=\Biggl{(}\frac{1}{H}\sum_{i=1}^{H}\left|y_{i}-\hat{y}_{i}\right|\Biggr{)}$ (18)
-where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the forecasted value at time $i$.
-II-D5 Accuracy of Probabilistic Forecasts
-For probabilistic forecasts, we employ the Energy Score (ES).
-The ES is a multivariate generalization of the continuous ranked probability score (CRPS), a widely used metric for evaluating probabilistic forecasts.
-The ES measures the distance between the forecasted distribution and the actual value.
-The lower the ES, the better the forecast.
-The ES is defined as:
-$ES=\left(\frac{1}{N}\sum_{j=1}^{N}\left\|y_{i}-\hat{y}_{ij}\right\|^{p}-\frac{1}{2N^{2}}\sum_{j=1}^{N}\sum_{k=1}^{N}\left\|\hat{y}_{ij}-\hat{y}_{ik}\right\|^{p}\right)^{\frac{1}{p}}$ (19)
-where $y_{i}$ is the actual value, $\hat{y}_{ij}$ is the $j$-th scenario of the forecasted value at time $i$, $N$ is the number of scenarios, and $p$ is the order of the norm.
-The second term in the equation is the mean pairwise distance between the scenarios.
-III Experimentation and Results
-Compared to theoretical analysis, practical setups are more complex.
-They often deal with non-trivial correlation structures in forecast errors and the potential for non-convex behaviors in optimization.
-This complexity is due to the necessity for cooperative behavior across multiple buildings.
-Additionally, various temporal dependencies emerge from switching costs and system feedback.
-To confirm the initial theoretical findings, we proceed with evaluating the performance of the FHC algorithm with different combinations of forecast and optimization commitment in a real-world scenario.
-According to the decision matrix in Figure I, at every time step the forecast and optimization plans can be updated or retained.
-Therefore, the forecast and optimization run with commitment periods $v_{F}$ and $v_{O}$.
-respectively.
-We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment.
-The result is shown in Tables II(a) and II(b).
-The scores in the tables indicate that updating the optimization plan more frequently than the forecast offers no added benefit.
-The optimal performance is achieved when the forecast and optimization are updated at the same frequency.
-Therefore, in the following analysis, we set $v=v_{F}=v_{O}$.
-(a) FHC algorithm with deterministic forecast
-$v_{O}$
-$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
-1 0.899
-2 0.910 0.904
-3 0.911 0.907 0.904
-4 0.912 0.908 0.907 0.905
-5 0.913 0.910 0.908 0.908 0.903
-6 0.912 0.910 0.907 0.908 0.908 0.906
-7 0.912 0.910 0.907 0.910 0.908 0.909 0.902
-8 0.910 0.907 0.908 0.905 0.906 0.908 0.902 0.903
-9 0.911 0.909 0.906 0.908 0.908 0.908 0.904 0.908 0.903
-10 0.913 0.909 0.909 0.907 0.906 0.908 0.903 0.906 0.905 0.905
-11 0.912 0.910 0.909 0.910 0.908 0.909 0.905 0.906 0.906 0.907 0.903
-12 0.917 0.911 0.911 0.909 0.908 0.910 0.904 0.907 0.907 0.907 0.904 0.908
-(b) FHC algorithm with stochastic forecast
-$v_{O}$
-$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
-1 0.875
-2 0.888 0.875
-3 0.894 0.886 0.873
-4 0.895 0.885 0.882 0.874
-5 0.899 0.892 0.885 0.886 0.875
-6 0.901 0.891 0.881 0.883 0.884 0.876
-7 0.901 0.896 0.887 0.888 0.886 0.883 0.875
-8 0.900 0.894 0.887 0.884 0.884 0.885 0.882 0.874
-9 0.902 0.894 0.883 0.887 0.884 0.882 0.883 0.880 0.876
-10 0.903 0.894 0.888 0.887 0.881 0.884 0.883 0.880 0.883 0.876
-11 0.902 0.895 0.889 0.887 0.885 0.885 0.884 0.880 0.884 0.882 0.876
-12 0.905 0.895 0.887 0.884 0.886 0.880 0.884 0.880 0.883 0.882 0.882 0.879
-TABLE II: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12
-Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability.
-In Figure 4, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods.
-The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings.
-Similar evaluation is done for the in-sample buildings and shown in Appendix B.
-It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods.
-The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
-For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours.
-The hypothesis is that this is due to the nature of the implemented method of noise simulation.
-The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
-The plots of the optimization scores show different behavior depending on the presence of the switching costs.
-The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship.
-The optimal commitment period remains at 1 for the deterministic forecast, while the stochastic forecast enables an optimal performance at a longer commitment period.
-These observations are supported by the correlation coefficients in the tables III(a) and III(b) with the correlation coefficients between the forecast properties and the optimization scores.
-The accuracy metrics are in positive correlation with the optimization score.
-When switching costs are factored in, the inverse correlation between the vertical stability and the cost becomes less pronounced.
-Yet, the correlation between the horizontal stability and the cost changes from negative to positive correlation when switching costs are included.
-Accuracy
-Optimization Score: without switching costs
-Stability: vertical (top) and horizontal (bottom)
-Optimization Score: with switching costs
-Point Forecast + Deterministic MPC
-Probabilistic forecast + Stochastic MPC
-Figure 4: Accuracy (Top), Horizontal (Middle) and Vertical Stability (Bottom row) of stochastic forecast over different revision periods with data from out-of-sample buildings.
-The optimization score is shown with and without switching costs.
-The forecast error is measured with MAE and Energy Score, while the stability is measured with MAC and EMD.
-The optimization score is measured with the cost function of the FHC algorithm.
-(a) FHC algorithm with deterministic forecast
-Correlation Coef. $MAE$ $MAC_{V}$ $MAC_{H}$
-$cost(FHC(v))$ 0.95 -0.94 -0.27
-$cost(FHC(v))$ w/ switching costs 0.70 -0.62 0.17
-(b) FHC algorithm with stochastic forecast
-Correlation Coef. $ES$ $EMD_{V}$ $EMD_{H}$
-$cost(FHC_{stochastic}(v))$ 0.63 -0.86 -0.29
-$cost(FHC_{stochastic}(v))$ w/ switching costs 0.71 -0.06 0.47
-TABLE III: Correlation Coefficients for Forecast Metrics and Optimization Score
-IV Discussion
-The observations from the results are consistent with the theoretical analysis of the upper bound on the competitive difference of the FHC algorithm.
-We observe that the optimal commitment period is dependent on the forecast properties.
-The inclusion of switching costs in the optimization problem changes the relationship between desired forecast properties and the optimization score.
-The forecast error is broadly lower with shorter commitment periods, however, the horizontal and vertical stability become also relevant.
-In the theoretical analysis, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs.
-Through testing with real data, we observe that switching costs can be partially mitigated by committing to more stable forecasts and optimization plans.
-The addition of switching costs makes the algorithmic problem harder as it forces current actions to depend on beliefs about future cost functions.
-In Figure 5, we can see that the shape of the final load curve is significantly different when switching costs are included into the optimization.
-Without switching costs, the battery is charged and discharged while following the price and emissions cost signals.
-The resulting net load curve (in green) is compared to the perfect scenario (in yellow), where the battery is charged and discharged following the perfect forecast.
-In the case with switching costs, the optimal policy is regulating the net load curve to a smoother shape with less drops and peaks.
-The charging and discharging actions are the means of such regulation.
-However, due to the error in the forecast, the actual net load curve (in green) is not as smooth and flat as the optimal policy (in yellow).
-It generally follows the optimal policy, but with oscillation around it.
-(a) Without switching costs
-(b) With switching costs
-Figure 5: The difference in net load between MPC with and without switching costs.
-Actual net load is shown in yellow.
-The base load is shown in green.
-The net load following the perfect forecast is shown in red.
-The electricity price is shown in blue.
-Minimal commitment periods $v$ of the control algorithm, combined with imperfect predictions, can lead to a scenario where the control strategy is constantly “chasing” the optimal state based on imperfect prediction.
-This can aggravate the impact of switching costs, as the control actions may change frequently in response to new information, even when such changes are not cost-effective in the long run.
-While traditional forward-looking RHC formulations can include terms in the objective function to penalize switching costs, the longer commitment period provides a native mechanism that respects past actions of the FHC algorithm.
-This mechanism can reduce the need for frequent adjustments, thus avoiding some of the switching costs.
-We identify two reasons for underperformance of the algorithm with a minimal commitment period (RHC, equivalent to FHC with $v_{O}=1$).
-Firstly, when the policy is computed for a period in a problem with switching cost, the past decisions are calculated as optimal with respect to future decisions given the forecast, however, after the policy is reviewed at the next iteration with an updated forecast, the optimality of the executed decision at the recent time step does not hold, since the future policies within the same horizon are reviewed.
-A simple illustrative example would be the battery is discharged first with an outlook for charging in the future when the price is more favorable.
-Once the forecast and policy updates, the updated policy is discharging again, however the battery has been discharged earlier.
-The second reason is that while the objective cost function does include the term for the switching costs $\beta||{x_{t}}-x_{t-1}||$, the switching costs are computed pairwise between loads at close time instances.
-However, since the total charging capacity of batteries is finite and limited, it is advantageous to regulate switching costs within longer action windows.
-When prediction is inaccurate in the look-ahead window, the RHC is more effective than FHC with a commitment period.
-If we make predictions accurate for a small window $\gamma$ and then inaccurate for the remaining $(\omega-\gamma)$ steps of the lookahead window, the fixed horizon policy is affected by the inaccurate predictions whereas RHC only acts on the correct ones.
-This is typically the case for a multi-step horizon forecast when the nearest steps are predicted at a higher accuracy.
-The situation changes for a stochastic MPC, generally hedged better against prediction errors.
-We can see from the results that FHC with longer commitment period becomes viable when the optimization is computed for a set of scenarios, as opposed to a deterministic point forecast.
-V Conclusion
-In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems.
-In summary, our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and severity of policy revisions, thereby mitigating the financial and operational impacts of switching costs.
-Our findings have significant implications for the design of energy management systems.
-They suggest that enhancing the stability of forecasts can improve system performance by reducing the need for frequent adjustments and managing the trade-offs between immediate and future resource allocation more effectively.
-Furthermore, we propose a novel approach to evaluating the stability of probabilistic forecasts using the Earth Mover’s Distance, which provides a robust measure of the dissimilarity between scenario sets.
-Looking ahead, there are questions remaining to be addressed.
-For future research, we recommend further exploration into methods that enhance forecast stability without compromising accuracy.
-It would be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading.
-Furthermore, from literature and current findings, there are indications that methods like AFHC and stochastic FHC, show more robustness in hedging against the forecast errors by aggregating policies with different underlying scenarios.
-However, they require more computational resources and are more complex to implement.
-An interesting avenue for future research would be defining the criteria for problems where these methods are most effective.
-These efforts would contribute to a deeper understanding of forecast-dependent optimization and improve the operational strategies of energy management systems facing an evolving energy landscape.
-Appendix A Cost Function and Optimization
-A-A Optimization score
-Optimized battery schedules are evaluated using a set of KPIs, each of which is targeted for minimization.
-The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$.
-The optimization score is defined as:
-Average $\text{Average Score}=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (20)
-When grid-related KPIs are included, the average includes the grid score $D$.
-The grid score is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$.
-The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting.
-Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, in this study we consider both optimizations with and without the grid score.
-The average score with the grid cost is defined as:
-[…]
-$\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$
-All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario.
-The normalized electricity cost, $C$, as delineated in Eq.
-22, is the ratio of electricity spending in a given policy, $c_{policy}$, to the spending in a reference scenario without battery intervention, $c_{{nobattery}}$ The cost metric $c$ is further defined in Eq.
-23 as the aggregate of the non-negative product of district-level net electricity price, $E_{h}\times T_{h}(\$)$, where $E_{h}$ denotes the electricity consumption of the district at the hour $h$ and $T_{h}$ specifies the electricity rate corresponding to that hour.
-$C=\frac{c_{{submission}}}{c_{{nobattery}}}$ (22)
-$c=\sum_{h=0}^{n-1}\max\left(0,E_{h}\times T_{h}\right)$ (23)
-Similarly, the normalized carbon emissions, $G$, is defined in Eq.
-24 as the ratio of district carbon emissions for a given policy, $g_{{policy}}$, relative to the emissions in the aforementioned baseline scenario, $g_{{nobattery}}$.
-The emission metric $g$ is elaborated in Eq.
-25 as the sum of carbon emissions, measured in $\left({kg}_{\mathrm{CO}_{2}\mathrm{e}}/\mathrm{kWh}\right)$, given by $E_{h}\times O_{h}$.
-Here, $O_{h}$ represents the carbon intensity for the hour $h$.
-$G=\frac{g_{{submission}}}{g_{{nobattery}}}$ (24)
-$g=\sum_{h=0}^{n-1}\max\left(0,E_{h}\times O_{h}\right)$ (25)
-Lastly, the evaluation metric includes a grid-related KPI, $D$.
-The metric follows grid-level objectives, such as the minimization of ramping and load factor.
-It is defined as the mean of the normalized ramping KPI, $R$, and the Load Factor KPI, $(1-L)$.
-The formulation, scaled by the grid cost in the baseline no-battery scenario, is given in Eq.
-26.
-$D=\operatorname{avg}\left(\frac{R_{\text{entry}}}{R_{\text{no\ battery}}},\frac{1-L_{\text{entry}}}{1-L_{\text{no\ battery}}}\right)$ (26)
-The ramping KPI, $R$, reflects the smoothness of the district’s load profile.
-A low $R$ indicates a gradual increase in grid electricity demand even after self-generation becomes unavailable in the evening and early morning, while a high $R$ indicates abrupt changes in load on the grid, which may lead to unscheduled strain on grid infrastructure and potential blackouts due to supply deficits.
-It is calculated as the sum of the absolute difference of net electricity consumption between consecutive time steps:
-$R=\sum_{t=1}^{8760}|E_{t}-E_{t-1}|$ (27)
-The Load Factor, $L$, indicates the efficiency of electricity consumption and is bounded between 0 (very inefficient) and 1 (highly efficient).
-Thus, the goal is to minimize $(1-L)$.
-$L$ is calculated as the average ratio of monthly average to maximum net electricity consumption:
-$L=\left(\frac{1}{12}\sum_{m=0}^{11}\left(\frac{\sum_{h=0}^{729}E_{730m+h}}{730\max(E_{730m},\ldots,E_{730m+729})}\right)\right)$ (28)
-A-B MPC Formulation
-$\mathbf{C}(S_{t},x_{t},\theta^{LA}_{t})=\sum_{i\in{price,carbon,grid}}\mathbf{C}_{i}(\tilde{S}{tt^{\prime}},\tilde{x}_{tt^{\prime}})$ (29)
-$=$ (30)
-subject to
-$SOC_{min}\leq SOC_{t}\leq SOC_{max},\quad\forall t,$ (31)
-$x_{t}=x_{t}^{pos}+x_{t}^{neg},\quad\forall t,$ (32)
-$-P_{max}\leq x_{t}^{neg}\leq 0,\quad\forall t,$ (33)
-$SOC_{t}=SOC_{t-1}+$ (34)
-$\eta_{charging}\cdot x_{t}^{pos}-\frac{1}{\eta_{discharging}}\cdot x_{t}^{neg},\quad\forall t.$ (35)
-In this formulation:
-  • $X_{t}^{MPC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the MPC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
-  • $W_{s}$ Weight factor for scenario $s$, indicating the importance of different scenarios in the decision-making process.
-We assume that the weight factors are equal for all scenarios.
-In the deterministic case, the number of scenarios is equal to 1.
-  • $C_{price}$, $C_{carbon}$, and $C_{grid}$ are cost functions representing the cost of electricity, the cost associated with carbon emissions, and the cost related to grid reliance, respectively.
-Each of these costs depends on the action taken ($x_{t}^{pos}$ and $x_{t}^{neg}$) and the current system state $S_{t}$.
-  • $SOC_{t}$ denotes the state of charge of the battery at time $t$, with $SOC_{min}$ and $SOC_{max}$ being the minimum and maximum allowable states of charge, respectively.
-  • $P_{max}$ is the maximum power with which the battery can be charged or discharged.
-  • $\eta_{charging}$ and $\eta_{discharging}$ are the charging and discharging efficiencies of the battery, respectively.
-Appendix B In-sample Performance
+A-D In-sample Performance
 (a) Performance of the FHC algorithm with deterministic forecast. Study over combinations of forecast and optimization commitment between 1 and 12

### v3

@@
 The proposed framework provides valuable insights for energy sector decision-makers and forecast practitioners when designing the operation of an energy management system.
-Index Terms:
+keywords:
 Energy Management Systems, Forecasting, Optimization, Switching Costs, Forecast Stability, Model Predictive Control
-I Introduction
-Managing energy assets in a grid system is a complex task that requires decision-making under uncertainty.
-The dynamics and stochasticity of the environment create a constantly evolving and inherently complex system.
-This reality necessitates making decisions with an incomplete knowledge of the future and for a limited time window, a scenario frequently approached through the lens of online optimization.
-To effectively manage this, it is useful to combine the forecast of the problem environment with the optimization of the decision-making process.
-This combination of forecasting and decision-making is often referred to as the predict, then optimize problem, or an online optimization with predictions.
-The online optimization problems with predictions are commonly seen in energy applications such as electric vehicle charging [1] [2], battery scheduling [3][4] and energy and flexibility dispatch problems [5] [6].
-In number of these applications, the decision-maker is faced with switching costs, which are the costs associated with updating the operational plan.
-For instance, transitioning from $x_{t-1}$ to $x_{t}$ typically incurs a switching cost or ramp cost, $d(x_{t}-x_{t-1})$.
-Consequently, the stage cost becomes time-coupled and is defined as: $C_{t}(x_{t-1},x_{t}):=f_{t}(x_{t})+d(x_{t}-x_{t-1})$.
-In the energy sector, the switching costs can arise from various factors, such as the physical limitations of ramping up or down energy production [7], managing server load balancing [8] and payment of network fees associated with energy trading [9].
-While there has been a significant amount of research on online optimization problems with predictions, there is a gap in the literature on the impact of switching costs on the interaction between forecast model and the optimization algorithm.
-We address this gap by studying two aspects of the combined system: time-coupling and forecast stability.
-Switching costs are a direct consequence of time-coupling, when the action $x_{t}$ at time $t$ influences the future actions $x_{t+1}$ and system states $S_{t+1}$, as defined by Lin et al.
-[8].
-Errors in forecasts act as ’switching incentive‘ compelling policy revisions using more accurate predictions aligned with up-to-date information and the actual state of the system.
-This perspective suggests a balancing act between forecast accuracy and switching costs.
-This paper explores how this trade-off influences the system’s design and operation, suggesting a nuanced view of forecasting’s role in optimization.
-In principle, forecasts used in optimization can be updated every time new information becomes available, often span multiple steps ahead and are revised periodically.
-Transitioning from the concept of time-coupling, we speculate that switching costs not only underscore the importance of forecast accuracy but also elevate the value of forecast stability.
-In this context, a stable forecast with minimal variation despite periodic updates, minimizes the need for frequent policy changes, effectively reducing switching costs.
+{highlights}
+The paper evaluates the impact of forecast stability and accuracy on the downstream performance of energy management systems under switching costs.
+A novel metric, Scenario Distribution Change (SDC), is proposed to assess the stability of probabilistic scenario sets, extending the Wasserstein distance.
+Forecast accuracy and stability are shown to influence the optimal policy commitment level. There is a theoretical trade-off between switching costs and forecast errors.
+Empirical analysis is conducted using the CityLearn 2022 competition case study, highlighting the trade-offs between forecast updates, stability, and operational outcomes.
+The longer than minimal commitment periods are found to improve system performance in case of stochastic optimization with probabilistic scenario sets.
+1 Introduction
+Managing energy assets within a grid system presents a challenging task characterized by decision-making under uncertainty.
+The inherent dynamics and stochasticity of the environment make this a complex system, constantly evolving and demanding decisions to be made with incomplete knowledge of the future over limited time windows, a scenario typically addressed through online optimization.
+To effectively navigate these challenges, it is beneficial to integrate forecasting of the problem environment with the optimization of decision-making processes.
+This integration, often referred to as the predict, then optimize approach, is prevalent in energy applications like electric vehicle charging [1, 2], battery scheduling [3, 4], and energy and flexibility dispatch problems [5, 6].
+In many of these applications, decision-makers face switching costs — expenses incurred from updating operational plans.
+For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down [7], managing server load balancing [8], or covering network fees associated with energy trading [9].
+Despite extensive research on online optimization with predictions, there remains a notable gap concerning how switching costs affect the synergy between forecasting models and optimization algorithms.
+This study addresses this gap by exploring the concepts of time-coupling and forecast stability within such systems.
+Time-coupling, where one decision directly affects subsequent decisions and system states, gives rise to switching costs.
+Forecast inaccuracies often lead to policy revisions, requiring adjustments to better align with the current system state and updated information.
+Striking a balance between forecast accuracy and the impact of switching costs is therefore crucial.
+This paper explores how this balance shapes system design and informs operational strategies.
+Building on this interplay between forecast accuracy and switching costs, we propose that the stability of forecasts plays an equally critical role.
+Since forecasts in optimization are typically updated with each new data acquisition—covering multiple future steps and undergoing regular revisions—stability becomes essential.
+A forecast that remains consistent across updates reduces the need for frequent strategic adjustments, thereby mitigating the impact of switching costs.
 Consequently, this research aims to answer the following questions:
+  • 1.
     Do switching costs influence the balance between forecast accuracy and stability?
+  • 2.
     What are the optimal strategies for forecast revision in the presence of such costs?
-  • And how can integrated systems be designed to effectively balance these considerations for improved decision-making in energy management?
-I-A Related work
+  • 3.
+    How can integrated systems be designed to effectively balance these considerations for improved decision-making in energy management?
+1.1 Related work
 Literature solutions for integrating forecasting and optimization are classified based on the linkage of these subproblems. There are three categories: Direct, Indirect, and Semi-direct methods.
@@
 An example is found in the work by Kazmi and Paskevich [13], where those authors propose a Bayesian Optimization approach for integrating the downstream optimization problem into the hyperparameter tuning process of the forecast model.
+Other example is [3], where the authors propose iterating attention over training data statically and dynamically to improve the forecast model performance in the downstream task.
 There is a noticeable gap in the literature when it comes to studying operational aspects of forecasting model and optimization deployment.
@@
 This topic has featured in research on online convex optimization (OCO) problems.
-In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm, also referred to as model predictive control (MPC).
+In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm.
+Model Predictive Control (MPC) is the most common implementation of RHC that employs a dynamic model of the process and is widely used in online optimization.
 When rerun at every time step, this method commits to the first step of the future horizon while treating later decisions as advisory.
+The Fixed Horizon Control (FHC) algorithm is a generalization of the algorithm that commits to a fixed number of steps $v$ in the future before re-optimizing the policy.
+The difference between the two algorithms is illustrated in Figure 1.
+Throughout this paper, we will use the term FHC to align with the focus on commitment-specific optimization.
+(a) FHC
+(b) RHC
+Figure 1: Diagrams of the Online Optimization algorithms. RHC: Receding Horizon Control, FHC: Fixed Horizon Control
 The research conducted by Chen et al.
-[15] sheds light on the role of noisy prediction in the design of the OCO algorithm.
-Those authors propose a model that simulates realistic prediction errors.
-Using this model, they show analytically that an alternative Averaging Fixed Horizon Control (AFHC) algorithm can achieve sublinear regret and a constant competitive ratio, which indicate a desirable level of optimality in reference to the theoretical best policy.
-Figure 1 shows an overview of the RHC and AFHC algorithm designs, along with other algorithms discussed further.
-Those authors state that other algorithms can reach better performance if they can use the predictions more efficiently.
-A later work, Chen et al.
-[16], introduces a class of policies known as Committed Horizon Control (CHC), which extends the principles of Receding Horizon Control (RHC) and Adaptive Forward Horizon Control (AFHC) through the integration of a commitment concept.
-This concept involves specifying a predetermined number of steps that the algorithm executes.
-The analysis of the generalized problem reveals a trade-off set by the varying levels of switching costs and quality of the predictions.
-The general observation is that the level of commitment is problem-specific and depends on the nature of the switching costs and the quality of the predictions.
-Therefore, the results show that the commitment level should be tuned for maximizing the performance of the policy.
-(a) RHC
-(b) FHC
-(c) AFHC
-(d) CHC
-Figure 1: Diagrams of the Online Optimization algorithms. RHC: Receding Horizon Control, FHC: Fixed Horizon Control, AFHC: Averaging Fixed Horizon Control, CHC: Committed Horizon Control.
-Current work sets to explore the performance of commitment-based policies in the context of energy management systems.
-We are particularly interested in analyzing the forecast and planning stability, as a potential factor that influences the performance of the policy.
-A comprehensive study of forecast stability is made by Godahewa et al.
-[17].
-Those authors categorize the stability into two types: vertical and horizontal.
-Vertical stability refers to the variation period-to-period variability between forecasts given for the same time period but issued from different origins.
-Horizontal stability describes the variation between predictions made at one instance within a prediction window.
-The lack of planning stability has long been recognized an important issue within research on supply chain management.
-The vertical stability in order forecasts is often referred to as the system nervousness.
-It is a well-known phenomenon that can lead to inefficient production and inventory management.
-The extra capacity implications are also referred to as the bullwhip effect [18].
-The commonly used methods to obtain a stable plan include increased commitment, referred as freezing intervals [19] [20], using smoothing techniques [21], and incorporating changing costs into the objective function [22] [23].
-Unlike supply chain management, energy management systems are more dynamic and require more frequent updates of forecasts and policies.
-Switching costs are also observed in energy, such as the ramping costs of generating units [7] [24] and toggling server load balancing [8] [25].
-Furthermore, the literature is scarcer when considering the operational aspects of a combined forecasting and optimization problem.
-Guo et al.
-[26] review the historical evolution and state-of-the-art for the optimization methods for energy storage dispatch.
-They conclude that current research indicates that RHC performance relies strongly on the quality of predictions, and therefore, it may not be robust to uncertainty.
-Christianson et al.
-[27] analyze the problem of resource procurement and online economic dispatch for power system operators.
-They present a customized FHC algorithm — Feasible Fixed Horizon Control (FFHC) — which incorporates a regularization term to solve the multi-interval optimization problem while ensuring feasibility and robustness amidst the complexities introduced by ramping costs.
-A study by La Tona et al.
-[28] explores how the frequency of forecast revisions affects the Energy Management System’s (EMS) performance in online replanning tasks, focusing on battery control to align actual and declared consumption.
-Their findings suggest that more frequent forecasts enhance EMS performance, with diminishing returns at intervals of 60 minutes and shorter.
-The current work examines applications in energy storage optimization and the relationship between forecast and optimal commitment level.
-I-B Contents and Contributions
-Through a systematic review of the literature, it becomes evident that the effects of predictions on online optimization problems with switching costs are not well understood.
-This present research aims to address this gap by studying the impact of forecast stability and accuracy on the performance of the policy in the context of energy management systems.
+[15] highlights the challenges posed by noisy predictions in designing Online Convex Optimization (OCO) algorithms.
+Several solutions, such as Averaging Fixed Horizon Control (AFHC) and its generalized counterpart, Committed Horizon Control (CHC) [16], propose aggregation strategies over multiple revisions of plans to achieve sublinear regret and improve robustness to prediction errors.
+Similarly, the Feasible Fixed Horizon Control (FFHC) algorithm [17] incorporates a regularization term to handle multi-interval optimization problems while ensuring feasibility and robustness against ramping costs.
+While these approaches demonstrate the benefits of aggregating policies over prediction updates, the current work shifts focus to understanding the effects of commitment on forecast stability and downstream optimization in the presence of switching costs.
+Unlike these aggregation-based methods, this study seeks to explore how the commitment level impacts system performance, with a particular emphasis on the relationship between forecast stability and switching costs.
+Forecast stability has been recognized as a critical factor in planning.
+Building on the comprehensive framework by Godahewa et al.
+[18], which categorizes stability into vertical (variations across forecast updates for the same period) and horizontal (variations within a prediction window), this work investigates these dimensions in the context of energy management systems.
+Unlike supply chain management, where freezing intervals [19, 20] and smoothing techniques [21] have been widely used to address system nervousness and the bullwhip effect [22], energy systems require more dynamic and frequent forecast updates.
+Nevertheless, switching costs are observed in energy management systems, such as ramping costs for generating units [7, 23] or toggling server load balancing [8, 24].
+However, the literature lacks a systematic exploration of how commitment levels influence the interplay between forecast stability and switching costs.
+By focusing on these aspects, the current research sets the stage for advancing the design and operation of energy management systems, particularly in applications involving energy storage optimization and real-time decision-making.
+1.2 Contents and Contributions
 The principle contributions are as follows:
+  • 1.
     First, we formally define the problem and the metric used for evaluating the downstream performance, both with and without switching costs.
 We also conduct a theoretical analysis of the performance bounds of the Fixed Horizon Control (FHC) algorithm, highlighting how the commitment level influences the trade-off between forecast accuracy and switching costs.
-  • We investigate the effect of forecast accuracy and stability on the downstream performance of the optimization policy.
+  • 2.
+    The effect of forecast accuracy and stability on the downstream performance of the optimization policy is analyzed.
 We extend the definition of forecast stability towards probabilistic forecasts and propose a metric for measuring horizontal and vertical stability of scenario sets.
-  • An empirical evaluation of the stability of deterministic and probabilistic scenario sets is conducted.
+  • 3.
+    We conduct an empirical evaluation of the stability of point forecasts and probabilistic scenario sets, using deterministic and stochastic optimization approaches, respectively.
 We use the case study of the integrated forecasting and optimization energy management problem from the Citylearn 2022 competition.
@@
 We show that running the policy for a longer period provides a simple mechanism to avoid switching costs while maintaining or improving the performance.
+  • 4.
     Lastly, the paper concludes with a discussion of the general implications of the results for the design of the energy management systems followed by suggested future research directions.
-II Methodology
-II-A Problem statement
+2 Methodology
+2.1 Problem statement
+Figure 2: Grid diagram illustrating the coordinated battery scheduling problem.
 This paper addresses the challenge of optimizing battery storage management within a multi-agent energy management system.
 The setting of the problem is first outlined in the CityLearn Challenge 2022 [25] and based on a real-world dataset from a grid-interactive community.
+The setup of the problem is shown in the Figure 2.
 More generally, we consider an online optimization with switching costs.
-The problem is solved with model predictive control (MPC) and full details of the objective score and relevant constraints are provided in Online Appendix A [^1] Appendix A.
-[Online] Available: https://arxiv.org/abs/2407.03368.
-The problem includes finding an optimal state-dependent policy $X_{t}^{MPC}(S_{t}|\theta^{LA})$.
+The problem is solved with the fixed horizon control (FHC), the variation of model predictive control (MPC) with a fixed commitment horizon.
+The full details of the objective score and relevant constraints are provided in A.1.
+The problem includes finding an optimal state-dependent policy $X_{t}^{FHC}(S_{t}|\theta^{LA})$.
 The decision-maker is faced with a sequence of decisions, where the decision at each step is based on the current state $S_{t}$ of the system and the direct lookahead approximation $\theta^{LA}$.
 Using the framework proposed by Powell [26], we use the following formulation:
-$X_{t}^{MPC}(S_{t}|\theta^{LA})=\arg\min_{x_{t}}\Biggl{(}\sum_{i\in n}W_{i}\cdot\mathbb{C}(\textbf{x};\textbf{S})\Biggr{)}=$
+$X_{t}^{FHC}(S_{t}|\theta^{LA})=\arg\min_{x_{t}}\Biggl{(}\sum_{i\in n}W_{i}\cdot\mathbb{C}(\textbf{x};\textbf{S})\Biggr{)}=$
 $\arg\min_{x_{t+1},\ldots,x_{t+H}}\Biggl{(}\sum_{t^{\prime}=t+1}^{t+H}\sum_{i\in n}W_{i}\cdot\Bigl{[}\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ (1)
@@
 In this formulation:
-  • $X_{t}^{MPC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharging $x_{t}^{neg}$ actions) made by the MPC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$.
+  • 1.
+    $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharging $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$.
 This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
+  • 2.
     $\mathbb{C}(\textbf{x};\textbf{S})$ is the cost function, which is a function of actions $\textbf{x}=x_{t},...x_{t+H}$ and states $\textbf{S}=S_{t},...S_{t+H}$.
+  • 3.
     $W_{s}$ Weight factor for scenario $s$, indicating the importance of different scenarios in the decision-making process.
@@
 In the deterministic case, the number of scenarios is equal to 1.
+  • 4.
     $\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ Represents the cost of the set of decisions $x_{t+1},...,x_{t+H}$ given the future forecast scenarios generated by the model $\theta^{LA}_{i,tt^{\prime}}$
+  • 5.
 […]
@@
 All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario.
-The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in Online Appendix A  [^2] Appendix A.
-[Online] Available: https://arxiv.org/abs/2407.03368.
-Furthermore, In the Online Appendix C  [^3] Appendix C.
-[Online] Available: https://arxiv.org/abs/2407.03368, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs.
+The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in A.2.
+Furthermore, In A.3, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs.
+Following the line of proof in Chen et al.
+[16], we derive the upper bound on the competitive difference between the FHC algorithm and the optimal policy.
+The analysis is conducted for the deterministic case, where the forecast is a point forecast, then extended to the probabilistic case, where the forecast is represented as a set of scenarios.
 We discuss the role of policy commitment and draw a parallel to stability.
 A trade off is demonstrated between forecast accuracy and switching costs, with the commitment level regulating the balance between the two.
-II-B Optimization
+2.2 Optimization
 The optimal policy is computed for the current state of the system and the forecast for the next $H$ time steps.
@@
 Similarly, the forecast can be reused or revised.
-The resulting decision matrix is shown in Table I.
-TABLE I: Decision Matrix illustrating the possible actions based on the decisions made in the forecasting and optimization stages.
+The resulting decision matrix is shown in Table 1.
+Table 1: Decision Matrix illustrating the possible actions based on the decisions made in the forecasting and optimization stages.
 Optimization Decision
@@
 forecasting & plan
-II-C Forecast
+2.3 Forecast
 Following the solution implemented in the winning submission of the competition, the forecasts are generated with gradient boosting regression trees and the linear least squares regression model.
@@
 Although this train-test splitting is not typical in applied tasks, the setup is still relevant to real-world scenarios, as the lack of historical data within the test subset is compensated by the availability of perfect weather forecast and similarity between the train and test sets.
-II-C1 Rolling Origin Window
+2.3.1 Rolling Origin Window
 In practical applications, the forecast is issued periodically for a finite number of future time steps.
@@
 The overlapping interval is highlighted in light grey.
-II-C2 Stability
+2.3.2 Stability
 As a potential factor influencing the performance of the policy, we are interested in the stability of the forecast.
 The stability of the forecast is defined as the consistency of the predictions over time.
-Following the categorization, first proposed by Godahewa et al.
-[17], we distinguish vertical and horizontal stability.
+Following the categorization, first proposed in [18], we distinguish vertical and horizontal stability.
 The difference between the two concepts is illustrated in Figure 4.
-We define the vertical stability as the difference between the predictions produced at different origins for the same time period, marked with the overlap area in Figure 2.
+(a) Vertical Stability
+(b) Horizontal Stability
+Figure 4: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions.
+The vertical stability is the difference between the predictions produced at different origins for the same time period.
 If the forecast revisions are vertically stable, the corresponding adjustment costs in policy are set to be minimal.
@@
 The horizontally stable forecasts are meant to reduce fluctuations between time steps within a prediction window.
-(a) Vertical Stability
-(b) Horizontal Stability
-Figure 3: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions
-Stability is evaluated for deterministic forecasts using the stability mean absolute change (MAC).
-The formulations of the stability metrics vary for vertical and horizontal types.
-In line with the definition of sMAPC in [34], we define the absolute scale stability metric, Mean Absolute Change (MAC).
-The metric measures the change between the predictions between overlapping rolling windows — ${MAC}_{V}$, or the change of predictions generated at the same origin — ${MAC}_{H}$
+Similarly to the evaluation in [30], a stability metric is measured for point forecasts using the mean absolute change (MAC).
+The metric measures the change between the predictions issued at different origins — ${MAC}_{V}$, or the change of predictions within the horizon window generated at the same origin — ${MAC}_{H}$
 $\operatorname{MAC}_{V}=\frac{\sum_{i=1}^{H-1}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i\mid t-1}\right|}{H-1}$ (4)
@@
 where $\hat{y}_{t+i\mid t}$ is the forecast for time $t+i$ generated at time $t$, $H$ is the forecast horizon, and $t$ is the current time step.
-II-C3 Stability of Probabilistic Forecasts
-The Earth Mover’s Distance (EMD), or Wasserstein distance, quantifies the minimum ”work” required to transform one distribution into another, making it ideal for assessing the stability of scenario sets in stochastic optimization.
-EMD is chosen for its symmetric nature, adherence to the triangle inequality, and ability to compare distributions on different probability spaces, making it effective even when forecasts differ significantly.
-It also accounts for the metric space dimension, akin to the mean absolute change (MAC) used for point forecasts.
-Other metrics like Kullback-Leibler (KL) divergence and Jensen-Shannon distance measure statistical distance, but EMD is preferred for scenario sample sets.
-In a 1-dimensional case with empirical sets $P$ and $Q$, EMD is defined as:
-$\operatorname{EMD}(P,Q)=\frac{1}{n}\sum_{i=1}^{n}\left\|P_{(i)}-Q_{(i)}\right\|$ (6)
-where $P_{(i)}$ and $Q_{(i)}$ are the sorted samples of $P$ and $Q$ respectively.
-This metric measures the dissimilarity between two sets of scenarios.
-To assess the stability of forecasts, the EMD can be computed for both horizontal and vertical stability:
-$EMD_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (7)
-$EMD_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (8)
-where $\hat{y}_{t+i\mid t,j}$ is the $j$-th scenario of the forecast for time step $t+i$ generated at time $t$, and $N$ is the number of scenarios.
-II-C4 Accuracy of Determinstic Forecasts
-For deterministic forecasts, the MAE is utilized to quantify the forecast accuracy.
+2.3.3 Stability of Probabilistic Forecasts
+In stochastic optimization, scenario stability measures the consistency of probabilistic forecast updates over time.
+We propose the Scenario Distribution Change (SDC) metric, adapting the Wasserstein Distance, also referred to as Earth Mover Distance, to assess the stability of scenario sets.
+SDC quantifies the average change between successive scenario updates, providing a clear measure of stability for probabilistic forecasts in dynamic systems.
+The Wasserstein distance formulation is symmetric, adheres to the triangle inequality, and effectively compares probabilistic scenario sets.
+Unlike metrics such as Kullback-Leibler (KL) divergence or Jensen-Shannon distance, SDC holds for non-overlapping distributions, making it particularly relevant for applications requiring forecast stability.
+By incorporating the dimension of the metric space, SDC is conceptually analogous to the Mean Absolute Change (MAC) used for point forecasts.
+To assess stability across time, SDC evaluates both horizontal stability (within a forecast horizon) and vertical stability (across forecast updates).
+For $N$ number of scenarios, these are defined as:
+Horizontal Stability (SDCH):
+$\operatorname{SDC}_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (6)
+Vertical Stability (SDCV):
+$\operatorname{SDC}_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (7)
+2.3.4 Accuracy of Determinstic Forecasts
+For point forecasts, the MAE is utilized to quantify the forecast accuracy.
 It calculates the average absolute difference between each forecasted value and the corresponding actual value.
@@
 where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the forecasted value at time $i$.
-II-C5 Accuracy of Probabilistic Forecasts
+2.3.5 Accuracy of Probabilistic Forecasts
 For probabilistic forecasts, we employ the Energy Score (ES).
@@
 The second term in the equation is the mean pairwise distance between the scenarios.
-III Experimentation and Results
+3 Experimentation and Results
 Compared to theoretical analysis, practical setups are more complex.
@@
 To confirm the initial theoretical findings, we proceed with evaluating the performance of the FHC algorithm with different combinations of forecast and optimization commitment in a real-world scenario.
-According to the decision matrix in Figure I, at every time step the forecast and optimization plans can be updated or retained.
-Therefore, the forecast and optimization run with commitment periods $v_{F}$ and $v_{O}$.
-respectively.
+According to the decision matrix in Figure 1, at every time step the forecast and optimization plans can be updated or retained.
+Therefore, the forecast and optimization run with commitment periods $v_{F}$ and $v_{O}$, respectively.
 We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment.
-The result is shown in Tables II(a) and II(b).
+The result is shown in Tables 2(a) and 2(b).
 The scores in the tables indicate that updating the optimization plan more frequently than the forecast offers no added benefit.
@@
 Therefore, in the following analysis, we set $v=v_{F}=v_{O}$.
-(a) FHC algorithm with deterministic forecast
+(a) deterministic FHC algorithm with point forecast
 $v_{O}$
@@
 12 0.917 0.911 0.911 0.909 0.908 0.910 0.904 0.907 0.907 0.907 0.904 0.908
-(b) FHC algorithm with stochastic forecast
+(b) stochastic FHC algorithm with probabilistic forecast
 $v_{O}$
@@
 12 0.905 0.895 0.887 0.884 0.886 0.880 0.884 0.880 0.883 0.882 0.882 0.879
-TABLE II: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12
+Table 2: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12
 Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability.
@@
 The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings.
-Similar evaluation is done for the in-sample buildings and shown in Online Appendix D  [^4] Appendix D.
-[Online] Available: https://arxiv.org/abs/2407.03368.
-It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods.
+Similar evaluation is done for the in-sample buildings and shown in A.4 It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods.
 The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
 For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours.
-The hypothesis is that this is due to the nature of the implemented method of noise simulation.
+The hypothesis is that this is due to the nature of the method of noise simulation, as outlined in [25].
 The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
@@
 The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship.
-The optimal commitment period remains at 1 for the deterministic forecast, while the stochastic forecast enables an optimal performance at a longer commitment period.
-These observations are supported by the correlation coefficients in the tables III(a) and III(b) with the correlation coefficients between the forecast properties and the optimization scores.
+The optimal commitment period remains at 1 for the point forecast, while the stochastic forecast enables an optimal performance at a longer commitment period.
+These observations are supported by the correlation coefficients in the tables 3(a) and 3(b) with the correlation coefficients between the forecast properties and the optimization scores.
 The accuracy metrics are in positive correlation with the optimization score.
@@
 Optimization Score: with switching costs
-Point Forecast + Deterministic MPC
-Probabilistic forecast + Stochastic MPC
+Point Forecast
++ Deterministic FHC
+Probabilistic forecast
++ Stochastic FHC
 Figure 5: Accuracy (Top), Horizontal (Second) and Vertical Stability (Third row) of stochastic forecast over different revision periods with data from out-of-sample buildings.
-The optimization score is shown with and without switching costs.
-The forecast error is measured with MAE and Energy Score, while the stability is measured with MAC and EMD.
-The optimization score is measured with the cost function of the FHC algorithm.
-(a) FHC algorithm with deterministic forecast
+(a) deterministic FHC algorithm
+with point forecast
 Correlation Coef. $MAE$ $MAC_{V}$ $MAC_{H}$
@@
 $cost(FHC(v))$ w/ switching costs 0.70 -0.62 0.17
-(b) FHC algorithm with stochastic forecast
-Correlation Coef. $ES$ $EMD_{V}$ $EMD_{H}$
-$cost(FHC_{stochastic}(v))$ 0.63 -0.86 -0.29
-$cost(FHC_{stochastic}(v))$ w/ switching costs 0.71 -0.06 0.47
-TABLE III: Correlation Coefficients for Forecast Metrics and Optimization Score
-IV Discussion
+(b) stochastic FHC algorithm
+with probabilistic forecast
+Correlation Coef. $ES$ $MASC_{V}$ $MASC_{H}$
+$cost(FHC_{stoch.}(v))$ 0.63 -0.86 -0.29
+$cost(FHC_{stoch.}(v))$ w/ switching costs 0.71 -0.06 0.47
+Table 3: Correlation Coefficients for Forecast Metrics and Optimization Score
+4 Discussion
 The observations from the results are consistent with the theoretical analysis of the upper bound on the competitive difference of the FHC algorithm.
@@
 The forecast error is broadly lower with shorter commitment periods, however, the horizontal and vertical stability become also relevant.
-In the theoretical analysis in Online Appendix C, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs.
+In the theoretical analysis in A.3, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs.
 Through testing with real data, we observe that switching costs can be partially mitigated by committing to more stable forecasts and optimization plans.
@@
 (b) With switching costs
-Figure 5: The difference in net load between MPC with and without switching costs.
+Figure 6: The difference in net load between FHC with and without switching costs.
 Actual net load is shown in yellow.
@@
 This is typically the case for a multi-step horizon forecast when the nearest steps are predicted at a higher accuracy.
-The situation changes for a stochastic MPC, generally hedged better against prediction errors.
+The situation changes for a stochastic FHC, generally hedged better against prediction errors.
 We can see from the results that FHC with longer commitment period becomes viable when the optimization is computed for a set of scenarios, as opposed to a deterministic point forecast.
-V Conclusion
+5 Conclusion
 In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems.
@@
 Our findings have significant implications for the design of energy management systems.
-They suggest that enhancing the stability of forecasts can improve system performance by reducing the need for frequent adjustments and managing the trade-offs between immediate and future resource allocation more effectively.
-Furthermore, we propose a novel approach to evaluating the stability of probabilistic forecasts using the Earth Mover’s Distance, which provides a robust measure of the dissimilarity between scenario sets.
+In the power system context, the switching costs are driven by imbalances, frequency deviations and additional stress on the power electronics.
+The current work suggests that enhancing the stability of forecasts can improve system performance by reducing the need for frequent adjustments and managing the trade-offs between immediate and future resource allocation more effectively.
+Furthermore, we propose $SDC$, a novel metric to evaluating the stability of probabilistic forecasts, which provides a robust measure of the dissimilarity between scenario sets.
 Looking ahead, there are questions remaining to be addressed.
@@
 These efforts would contribute to a deeper understanding of forecast-dependent optimization and improve the operational strategies of energy management systems facing an evolving energy landscape.
+Acknowledgments
+This project has received funding from the European Union’s research and innovation programme Horizon Europe under the grant agreement No.101136211
 Appendix A Cost Function and Optimization
-A-A Optimization score
+A.1 Optimization score
 Optimized battery schedules are evaluated using a set of KPIs, each of which is targeted for minimization.
@@
 The normalized electricity cost, $C$, as delineated in Eq.
-13, is the ratio of electricity spending in a given policy, $c_{policy}$, to the spending in a reference scenario without battery intervention, $c_{{nobattery}}$ The cost metric $c$ is further defined in Eq.
+12, is the ratio of electricity spending in a given policy, $c_{policy}$, to the spending in a reference scenario without battery intervention, $c_{{nobattery}}$ The cost metric $c$ is further defined in Eq.
 13 as the aggregate of the non-negative product of district-level net electricity price, $E_{h}\times T_{h}(\$)$, where $E_{h}$ denotes the electricity consumption of the district at the hour $h$ and $T_{h}$ specifies the electricity rate corresponding to that hour.
@@
 The formulation, scaled by the grid cost in the baseline no-battery scenario, is given in Eq.
-17.
+16.
 $D=\operatorname{avg}\left(\frac{R_{\text{entry}}}{R_{\text{no\ battery}}},\frac{1-L_{\text{entry}}}{1-L_{\text{no\ battery}}}\right)$ (16)
@@
 $L=\left(\frac{1}{12}\sum_{m=0}^{11}\left(\frac{\sum_{h=0}^{729}E_{730m+h}}{730\max(E_{730m},\ldots,E_{730m+729})}\right)\right)$ (18)
-A-B MPC Formulation
+A.2 Optimization Formulation
 $\mathbf{C}(S_{t},x_{t},\theta^{LA}_{t})=\sum_{i\in{price,carbon,grid}}\mathbf{C}_{i}(\tilde{S}{tt^{\prime}},\tilde{x}_{tt^{\prime}})$ (19)
@@
 In this formulation:
-  • $X_{t}^{MPC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the MPC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
+  • 1.
+    $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
+  • 2.
     $W_{s}$ Weight factor for scenario $s$, indicating the importance of different scenarios in the decision-making process.
@@
 In the deterministic case, the number of scenarios is equal to 1.
+  • 3.
     $C_{price}$, $C_{carbon}$, and $C_{grid}$ are cost functions representing the cost of electricity, the cost associated with carbon emissions, and the cost related to grid reliance, respectively.
 Each of these costs depends on the action taken ($x_{t}^{pos}$ and $x_{t}^{neg}$) and the current system state $S_{t}$.
+  • 4.
     $SOC_{t}$ denotes the state of charge of the battery at time $t$, with $SOC_{min}$ and $SOC_{max}$ being the minimum and maximum allowable states of charge, respectively.
+  • 5.
     $P_{max}$ is the maximum power with which the battery can be charged or discharged.
+  • 6.
     $\eta_{charging}$ and $\eta_{discharging}$ are the charging and discharging efficiencies of the battery, respectively.
-A-C Theoretical Analysis of Performance Bounds
+A.3 Theoretical Analysis of Performance Bounds
 In this section, we analyze the upper bound on the performance of the FHC algorithm with a limited commitment $v$.
@@
 Hence, the competitive difference of the FHC algorithm in the stochastic case will likely be lower, favoring a higher commitment level, as the first term (switching costs) decreases with $v$ at the same rate as in the deterministic case.
-A-D In-sample Performance
-(a) Performance of the FHC algorithm with deterministic forecast. Study over combinations of forecast and optimization commitment between 1 and 12
+A.4 In-sample Performance
+(a) Performance of the FHC algorithm with point forecast.
+Study over combinations of forecast and optimization commitment between 1 and 12
 $v_{O}$
@@
 12 0.495 0.495 0.495 0.495 0.495 0.495 0.496 0.496 0.495 0.496 0.496 0.495
-(b) Performance of the FHC algorithm with stochastic forecast. Study over combinations of forecast and optimization commitment between 1 and 12
+(b) Performance of the FHC algorithm with stochastic forecast.
+Study over combinations of forecast and optimization commitment between 1 and 12
 $v_{O}$
@@
 12 0.813 0.809 0.807 0.805 0.806 0.808 0.807 0.806 0.807 0.808 0.808 0.807
-TABLE IV: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12 (with data from in-sample buildings)
+Table 4: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12 (with data from in-sample buildings)
 Accuracy
@@
 Optimization Score: with switching costs
-Point Forecast + Deterministic MPC
-Probabilistic forecast + Stochastic MPC
+Point Forecast + Deterministic FHC
+Probabilistic forecast + Stochastic FHC
 Figure 7: Accuracy (Top), Horizontal (Middle) and Vertical Stability (Bottom row) of stochastic forecast over different revision periods with data from in-sample buildings

### v4

@@
 In many of these applications, decision-makers face switching costs — expenses incurred from updating operational plans.
-For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down [7], managing server load balancing [8], or covering network fees associated with energy trading [9].
+For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down [7], managing server load balancing [8], or covering network fees associated with energy trading [9, 10].
 Despite extensive research on online optimization with predictions, there remains a notable gap concerning how switching costs affect the synergy between forecasting models and optimization algorithms.
@@
 Forecast inaccuracies often lead to policy revisions, requiring adjustments to better align with the current system state and updated information.
-Striking a balance between forecast accuracy and the impact of switching costs is therefore crucial.
+Striking a balance between forecast accuracy and the impact of switching costs is crucial.
 This paper explores how this balance shapes system design and informs operational strategies.
-Building on this interplay between forecast accuracy and switching costs, we propose that the stability of forecasts plays an equally critical role.
-Since forecasts in optimization are typically updated with each new data acquisition—covering multiple future steps and undergoing regular revisions—stability becomes essential.
-A forecast that remains consistent across updates reduces the need for frequent strategic adjustments, thereby mitigating the impact of switching costs.
 Consequently, this research aims to answer the following questions:
@@
 Other example is [3], where the authors propose iterating attention over training data statically and dynamically to improve the forecast model performance in the downstream task.
-There is a noticeable gap in the literature when it comes to studying operational aspects of forecasting model and optimization deployment.
-For forecasting, the frequency of the forecast revision can be interpreted as both the frequency of the forecast model deployment and the frequency of model training.
-The latter aspect is studied by Spiliotis and Petropoulos [14], where those authors investigate different scenarios of updating the model fit for univariate exponential smoothing (ES) and univariate gradient boosting models (LightGBM).
-We are particularly interested in the applications of optimization with predictions in environments where switching costs are present.
+Among the semi-direct methods, there is a noticeable gap in the literature regarding the operational aspects of integrating forecasting models and optimization deployment.
+A relevant study by Prat et al.
+[15] addresses the issue of determining the minimum forecast horizon for storage scheduling problems in a rolling-horizon approach.
+The study introduces a verifiable condition to check if a chosen planning horizon is sufficiently long.
+Other researchers study the effect of the frequency of forecast revision, which can be interpreted as both the frequency of the forecast model deployment and the frequency of model training.
+The frequency of retraining is studied by Spiliotis and Petropoulos [16], where the authors investigate different scenarios of updating the model fit for univariate exponential smoothing (ES) and univariate gradient boosting models (LightGBM).
+We are particularly interested in studying the model deployment frequency.
+More specifically, we look at the applications of optimization with predictions in environments where switching costs are present.
 This topic has featured in research on online convex optimization (OCO) problems.
@@
 Unlike supply chain management, where freezing intervals [21, 22] and smoothing techniques [23] have been widely used to address system nervousness and the bullwhip effect [24], energy systems require more dynamic and frequent forecast updates.
-Nevertheless, switching costs are observed in energy management systems, such as ramping costs for generating units [7, 23] or toggling server load balancing [8, 24].
+Switching costs are also observed in energy management systems, such as ramping costs for generating units [7, 25] or toggling server load balancing [8, 26].
 However, the literature lacks a systematic exploration of how commitment levels influence the interplay between forecast stability and switching costs.
@@
 The setting of the problem is first outlined in the CityLearn Challenge 2022 [27] and based on a real-world dataset from a grid-interactive community.
-The setup of the problem is shown in the Figure 2.
-More generally, we consider an online optimization with switching costs.
-The problem is solved with the fixed horizon control (FHC), the variation of model predictive control (MPC) with a fixed commitment horizon.
+The CityLearn Challenge is an annual competition that focuses on the development of reinforcement learning agents for energy management in buildings.
+Participants are provided with a simulated environment that mimics real-world energy consumption and generation patterns.
+The goal is to optimize the operation of energy systems, such as batteries and HVAC systems, to minimize costs and emissions while maintaining comfort levels.
+The series of CityLearn competitions, starting from 2020, was originally set up only for Reinforcement Learning (RL) solutions, where an agent learns to make decisions by performing actions and receiving feedback from the environment.
+However, from 2022 onwards, alternative methods such as Model Predictive Control (MPC), which uses a model to predict the future behavior of a system and optimize the control actions accordingly, and Rule-Based Control (RBC), which involves using predefined rules to make decisions based on the current state of the system, have also been employed.
+The competition provides a platform and the open-source simulation environment [^1] https://github.com/intelligent-environments-lab/CityLearn for implementing the operation of energy management systems with setups that include various energy assets, such as solar panels, batteries, heat pumps, connections to the grid, and electric vehicles.
+The setup of the investigated problem is illustrated in Figure Figure 2.
+In general terms, we consider an online optimization problem with switching costs.
+The data of the competition has the hourly resolution.
+The core problem is multi-agent scheduling of energy storage over a period of a year.
+While the problem can be formulated similarly to RL, or a dynamic programming approach, in this paper, we particularly investigate the mechanics with a rolling-horizon approach.
+This means that the problem is solved iteratively over a planning horizon $H$.
+However, it is typical for only first $v$ steps to be implemented with the rest of the prediction horizon $H-v$ to be advisory, since the optimization is revised with a period of $v$ steps.
 The full details of the objective score and relevant constraints are provided in A.1.
@@
 […]
-Optimized battery schedules are evaluated using a set of KPIs.
+Optimized battery schedules are evaluated using a set of Key Performance Indicators (KPIs).
 The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$.
-The optimization score is defined as:
-Average $\text{Average Score}=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (2)
-When grid-related Key Performance Indicators (KPIs) are included, the average includes the grid score $D$.
-The grid score is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$.
+When grid-related metrics are included, the average includes the grid score $D$, which is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$.
 The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting.
-Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, in this study we consider both optimizations with and without the grid score.
-The average score with the grid cost is defined as:
-[…]
-$\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$
+Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, the combined optimization scores are defined with and without the grid score:
+Average Score $=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (2)
+Average Score (with switching costs) $=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$ (3)
 All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario.
@@
 2.2 Optimization
+The problem is solved using the Fixed Horizon Control (FHC) algorithm, which is a configuration of model predictive control (MPC) with a fixed commitment horizon.
 The optimal policy is computed for the current state of the system and the forecast for the next $H$ time steps.
@@
 2.3 Forecast
+The target variable for forecasting is the aggregate uncontrollable load of the energy community in Figure 2.
 Following the solution implemented in the winning submission of the competition, the forecasts are generated with gradient boosting regression trees and the linear least squares regression model.
 The gradient boosting implementation is LightGBM [29].
-Predictions are made for the next 24 hours with 24 independent models.
+Predictions are made directly for the next 24 hours with 24 independent models.
 The scenario sets are generated with added gaussian noise.
@@
 For example, if we have a window size of 3 and a step size of 1, we would first generate the predictions and optimize for the steps 1-3, then at a second iteration, the forecasts are issued for the steps 2-4.
-Figure 3: Vertical stability between rolling origin windows
-in time-series forecast
+Figure 3: Updating Rolling Origin Window. The forecast
+is updated every $v_{f}$ time steps and possible overlaps between prediction windows are shown.
 The rolling horizon setup enables an update of the prediction using the most recent information.
 When the commitment period is shorter than the forecast horizon, the regular updates create overlaps between prediction windows, when the forecasts associated with a particular target time in the future are released in multiple revisions.
-The overlaps between windows with a rolling origin are shown in the Figure 3.
-The overlapping interval is highlighted in light grey.
+The resulting overlapping interval is highlighted in light grey in Figure 3.
+These overlaps are considered when evaluating stability of the forecast.
 2.3.2 Stability
-As a potential factor influencing the performance of the policy, we are interested in the stability of the forecast.
-The stability of the forecast is defined as the consistency of the predictions over time.
+Stability of the forecast is defined as the consistency of the predictions over time.
 Following the categorization, first proposed in [20], we distinguish vertical and horizontal stability.
@@
 Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability.
-In Figure 5, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods.
+In Figure 5, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods between 1 and 12.
 The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings.
-Similar evaluation is done for the in-sample buildings and shown in A.4 It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods.
+Similar evaluation is done for the in-sample buildings and shown in A.4.
+It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods.
 The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
@@
 The optimal commitment period remains at 1 for the point forecast, while the stochastic forecast enables an optimal performance at a longer commitment period.
+This observation is notable as the most expected behavior is that best performance is achieved with shortest revision periods for both forecast types, as more up-to-date forecasts are provided.
 These observations are supported by the correlation coefficients in the tables 3(a) and 3(b) with the correlation coefficients between the forecast properties and the optimization scores.
@@
 The addition of switching costs makes the algorithmic problem harder as it forces current actions to depend on beliefs about future cost functions.
-In Figure 6, we can see that the shape of the final load curve is significantly different when switching costs are included into the optimization.
-Without switching costs, the battery is charged and discharged while following the price and emissions cost signals.
-The resulting net load curve (in green) is compared to the perfect scenario (in yellow), where the battery is charged and discharged following the perfect forecast.
-In the case with switching costs, the optimal policy is regulating the net load curve to a smoother shape with less drops and peaks.
-The charging and discharging actions are the means of such regulation.
-However, due to the error in the forecast, the actual net load curve (in green) is not as smooth and flat as the optimal policy (in yellow).
-It generally follows the optimal policy, but with oscillation around it.
+In Figure 6, we can see that the shape of the optimal EMS net load curve is significantly different when switching costs are included into the optimization.
+Without switching costs, the battery is charged and discharged while following the price and emissions cost signals, indicated in the bottom subplot.
+The plot also highlights the difference in net load between the storage management with oracle information (in yellow) and forecast-based management (in blue).
+The oracle information, also referred to as the perfect forecast, is the ideal scenario where the EMS uses the perfect knowledge of the future load signals.
+Without considering switching costs, it is observed that scheduling with the oracle information has higher spikes.
+We assume that the oracle information given optimizing agent the confidence to charge at maximum capacity when the price and emissions are low.
+However, the forecast-based management is more conservative in its actions, as it is uncertain about the future load signals.
+The figure illustrates the difference in net load for two scenarios: one without switching costs and one with.
+When switching costs are considered, the optimal policy smooths the net load curve, resulting in a more stable profile with fewer fluctuations.
+The charging and discharging actions contribute to this smoothing, making the aggregate net load more stable.
+However, the presence of forecast errors exacerbates the switching costs.
+The aggregate EMS-controlled net load curve (in green) is not as smooth and flat as the optimal policy (in yellow).
+While it generally follows the optimal policy, it oscillates around it.
 (a) Without switching costs
@@
 This can aggravate the impact of switching costs, as the control actions may change frequently in response to new information, even when such changes are not cost-effective in the long run.
-While traditional forward-looking RHC formulations can include terms in the objective function to penalize switching costs, the longer commitment period provides a native mechanism that respects past actions of the FHC algorithm.
-This mechanism can reduce the need for frequent adjustments, thus avoiding some of the switching costs.
-We identify two reasons for underperformance of the algorithm with a minimal commitment period (RHC, equivalent to FHC with $v_{O}=1$).
-Firstly, when the policy is computed for a period in a problem with switching cost, the past decisions are calculated as optimal with respect to future decisions given the forecast, however, after the policy is reviewed at the next iteration with an updated forecast, the optimality of the executed decision at the recent time step does not hold, since the future policies within the same horizon are reviewed.
+In line with the theoretical analysis in Appendix A.3, the upper bound on the competitive difference of the FHC algorithm is dependent on the trade-off between switching costs and forecast errors.
+While the MPC objective functions factors in the switching costs, we identify two factors that can lead to underperformance of FHC with low commitment period ($v_{O}=1$), particularly when probabilistic forecasts are used.
+Firstly, when the policy is computed for a finite period in a problem with switching cost, the past decisions are calculated as optimal with respect to future decisions given the forecast, however, after the policy is reviewed at the next iteration with an updated forecast, the optimality of the previously executed decision may be suboptimal given the updated information.
 A simple illustrative example would be the battery is discharged first with an outlook for charging in the future when the price is more favorable.
@@
 However, since the total charging capacity of batteries is finite and limited, it is advantageous to regulate switching costs within longer action windows.
-When prediction is inaccurate in the look-ahead window, the RHC is more effective than FHC with a commitment period.
-If we make predictions accurate for a small window $\gamma$ and then inaccurate for the remaining $(\omega-\gamma)$ steps of the lookahead window, the fixed horizon policy is affected by the inaccurate predictions whereas RHC only acts on the correct ones.
-This is typically the case for a multi-step horizon forecast when the nearest steps are predicted at a higher accuracy.
-The situation changes for a stochastic FHC, generally hedged better against prediction errors.
-We can see from the results that FHC with longer commitment period becomes viable when the optimization is computed for a set of scenarios, as opposed to a deterministic point forecast.
+Other important observation is that the FHC algorithm with longer commitment periods only becomes viable when the optimization is computed for a set of scenarios, as we can see from the results in Figure 5.
+It is hypothesized, that the aggregation of policies over different scenarios provides a more robust policy that hedges against the forecast errors.
+Generally, within a forecast horizon, the nearest steps are predicted with higher accuracy than further steps.
+Longer commitment periods aggravate the impact of forecast errors due to time coupling, as the suboptimal decisions made in the past are amplified in the future.
+A stochastic FHC algorithm is better equipped to handle errors in the forecast, as it aggregates policies over different scenarios, therefore, the impact of forecast errors is mitigated and longer commitment periods are more viable.
 5 Conclusion
+To summarize, the inclusion of switching costs in the FHC algorithm introduces a trade-off between forecast errors and switching costs.
+While traditional FHC formulations can penalize switching costs in the objective function, longer commitment periods provide a natural mechanism to use a stable forecast and reduce the need for frequent adjustments.
+This approach is more viable when optimization is computed over a set of scenarios, since the aggregation of policies offers a more robust policy, hedging against increasing forecast errors in a forecast horizon.
 In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems.
-In summary, our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and severity of policy revisions, thereby mitigating the financial and operational impacts of switching costs.
+Our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and switching costs of policy revisions, thereby mitigating the financial and operational impacts.
 Our findings have significant implications for the design of energy management systems.
 In the power system context, the switching costs are driven by imbalances, frequency deviations and additional stress on the power electronics.
-The current work suggests that enhancing the stability of forecasts can improve system performance by reducing the need for frequent adjustments and managing the trade-offs between immediate and future resource allocation more effectively.
+The current work suggests that enhancing the stability of forecasts leads to stability of policy which can improve system performance by managing the trade-offs between forecast accuracy and switching costs.
 Furthermore, we propose $SDC$, a novel metric to evaluating the stability of probabilistic forecasts, which provides a robust measure of the dissimilarity between scenario sets.
@@
 For future research, we recommend further exploration into methods that enhance forecast stability without compromising accuracy.
+Further research into the link between forecast quality and optimization performance is needed.
+Parameters such as forecast horizon, commitment period, and the number of scenarios are subject to a study for deeper understanding of their impact on the optimization performance and dependencies to the forecast accuracy and stability.
 It would be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading.
-Furthermore, from literature and current findings, there are indications that methods like AFHC and stochastic FHC, show more robustness in hedging against the forecast errors by aggregating policies with different underlying scenarios.
-However, they require more computational resources and are more complex to implement.
+From literature and current findings, methods like AFHC and stochastic FHC, show more robustness in hedging against the forecast errors, however, they require more computational resources and are more complex to implement.
 An interesting avenue for future research would be defining the criteria for problems where these methods are most effective.

### v5

@@
-Predict. Optimize. Revise. On Forecast and Policy Stability in Energy Management Systems
+Balancing Forecast Accuracy and Switching Costs in Online Optimization of Energy Management Systems
 Abstract
-This research addresses the challenge of integrating forecasting and optimization in energy management systems, focusing on the impacts of switching costs, forecast accuracy, and stability.
-It proposes a novel framework for analyzing online optimization problems with switching costs and enabled by deterministic and probabilistic forecasts.
-Through empirical evaluation and theoretical analysis, the research reveals the balance between forecast accuracy, stability, and switching costs in shaping policy performance.
-Conducted in the context of battery scheduling within energy management applications, it introduces a metric for evaluating probabilistic forecast stability and examines the effects of forecast accuracy and stability on optimization outcomes using the real-world case of the Citylearn 2022 competition.
-Findings indicate that switching costs significantly influence the trade-off between forecast accuracy and stability, highlighting the importance of integrated systems that enable collaboration between forecasting and operational units for improved decision-making.
-The study shows that committing to a policy for longer periods can be advantageous over frequent updates.
-Results also show a correlation between forecast stability and policy performance, suggesting that stable forecasts can mitigate switching costs.
-The proposed framework provides valuable insights for energy sector decision-makers and forecast practitioners when designing the operation of an energy management system.
+This study investigates the integration of forecasting and optimization in energy management systems, with a focus on the role of switching costs, defined as penalties incurred from frequent operational adjustments.
+We develop a theoretical and empirical framework to examine how forecast accuracy and stability interact with switching costs in online decision-making settings.
+Our analysis spans both deterministic and stochastic optimization approaches, using point and probabilistic forecasts.
+A novel metric for measuring temporal consistency in probabilistic forecasts is introduced, and the framework is validated in a real-world battery scheduling case based on the CityLearn 2022 challenge.
+Results show that switching costs significantly alter the trade-off between forecast accuracy and stability, and that more stable forecasts can reduce the performance loss due to switching.
+Contrary to common practice, the findings suggest that, under non-negligible switching costs, longer commitment periods may lead to better overall outcomes.
+These insights have practical implications for the design of intelligent, forecast-aware energy management systems.
 keywords:
 Energy Management Systems, Forecasting, Optimization, Switching Costs, Forecast Stability, Model Predictive Control
-{highlights}
-The paper evaluates the impact of forecast stability and accuracy on the downstream performance of energy management systems under switching costs.
-A novel metric, Scenario Distribution Change (SDC), is proposed to assess the stability of probabilistic scenario sets, extending the Wasserstein distance.
-Forecast accuracy and stability are shown to influence the optimal policy commitment level. There is a theoretical trade-off between switching costs and forecast errors.
-Empirical analysis is conducted using the CityLearn 2022 competition case study, highlighting the trade-offs between forecast updates, stability, and operational outcomes.
-The longer than minimal commitment periods are found to improve system performance in case of stochastic optimization with probabilistic scenario sets.
+[…]
+[cor1]Corresponding author
 1 Introduction
@@
 To effectively navigate these challenges, it is beneficial to integrate forecasting of the problem environment with the optimization of decision-making processes.
-This integration, often referred to as the predict, then optimize approach, is prevalent in energy applications like electric vehicle charging [1, 2], battery scheduling [3, 4], and energy and flexibility dispatch problems [5, 6].
+This integration, often referred to as the predict, then optimize approach, is prevalent in energy applications like electric vehicle charging Kriekinge et al.
+[2021], Tungom et al.
+[2024], battery scheduling Ruddick et al.
+[2024], Liu et al.
+[2023], and energy and flexibility dispatch problems Qu et al.
+[2024], Liu et al.
+[2025].
 In many of these applications, decision-makers face switching costs — expenses incurred from updating operational plans.
-For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down [7], managing server load balancing [8], or covering network fees associated with energy trading [9, 10].
+For example, transitioning operations from one state to another often involves costs due to physical limitations on ramping energy production up or down Wu et al.
+[2013], managing server load balancing Gupta [2021], Lin et al.
+[2012], or covering network fees associated with energy trading Yang et al.
+[2023], Levorato et al.
+[2022].
 Despite extensive research on online optimization with predictions, there remains a notable gap concerning how switching costs affect the synergy between forecasting models and optimization algorithms.
@@
 Forecast inaccuracies often lead to policy revisions, requiring adjustments to better align with the current system state and updated information.
-Striking a balance between forecast accuracy and the impact of switching costs is crucial.
+A key consideration in these systems is determining the optimal commitment period, meaning how long one can commit to a given policy before recalculating based on updated forecasts.
+This commitment level directly influences system performance, especially when switching costs are present.
+Additionally, forecast stability, the consistency of predictions across successive updates, plays a critical role in the reliability and efficacy of the optimization process.
+Striking a balance between forecast accuracy, stability, and the impact of switching costs is crucial.
 This paper explores how this balance shapes system design and informs operational strategies.
 Consequently, this research aims to answer the following questions:
-  • 1.
-    Do switching costs influence the balance between forecast accuracy and stability?
-  • 2.
-    What are the optimal strategies for forecast revision in the presence of such costs?
-  • 3.
-    How can integrated systems be designed to effectively balance these considerations for improved decision-making in energy management?
+  • Does presence of switching costs affect the optimal commitment level of the policy?
+  • How do forecast accuracy and stability influence the downstream performance of energy management systems?
+  • How can integrated systems of forecasting and optimization be designed to effectively balance these considerations for improved decision-making in energy management?
 1.1 Related work
-Literature solutions for integrating forecasting and optimization are classified based on the linkage of these subproblems. There are three categories: Direct, Indirect, and Semi-direct methods.
+Literature solutions for integrating forecasting and optimization are classified based on the level of integration between the two components.
+This classification includes three categories: Direct, Indirect, and Semi-direct methods.
 Direct methods engage with the optimization problem during training of the forecast making it an integrated approach which is known as ‘predict and optimize’.
@@
 The computational burden is partially resolved by using a simplified loss function, however more complex problems can still pose significant computational challenges.
-There are also other works that design an alternative task-specific loss function, such as reported in Mandi et al.
+There are other works that design an alternative task-specific loss function, such as reported in Mandi et al.
 [2020], and tested on a real-world energy management problem.
-Indirect methods regard the two subproblems as separate, also referred in literature as ‘predict, then optimize’.
+Indirect methods regard the two sub-problems as separate, also referred in literature as ‘predict, then optimize’.
 This approach is regarded as standard and is widely applied in practice.
@@
 [2022], an empirical evaluation of direct and indirect approaches is conducted in the context of cost-sensitive classification.
-Although the study is not directly applicable to the scheduling problems, it is interesting to note that those authors find that the indirect approach where optimization is performed on the predictions of the classifier outperforms the integrated approach.
+Although the study is not directly applicable to the scheduling problems, it is interesting to note that those authors find that the indirect approach where optimization is performed on the predictions of a classification algorithm outperforms the integrated approach.
 Therefore, the effectiveness of cost-sensitive classification may not always align with the intuitively expected benefits of integrated training within the ’predict and optimize’.
 The third category, Semi-direct methods, considers characteristics of the optimization problem but abstains from direct interaction during training of the forecast.
-These methods can take various forms, as potential improvements can occur at any stage outside of the training process.
+These methods can take various forms, as potential improvements can occur at any stage outside the training process.
 An example is found in the work by Kazmi and Paskevich [2023], where those authors propose a Bayesian Optimization approach for integrating the downstream optimization problem into the hyperparameter tuning process of the forecast model.
-Other example is [3], where the authors propose iterating attention over training data statically and dynamically to improve the forecast model performance in the downstream task.
+Other example is Huang et al.
+[2025], where the authors propose iterating attention over training data statically and dynamically to improve the forecast model performance in the downstream task.
 Among the semi-direct methods, there is a noticeable gap in the literature regarding the operational aspects of integrating forecasting models and optimization deployment.
@@
 [2024] addresses the issue of determining the minimum forecast horizon for storage scheduling problems in a rolling-horizon approach.
-The study introduces a verifiable condition to check if a chosen planning horizon is sufficiently long.
+The study introduces a verifiable condition to check if a selected planning horizon is sufficiently long.
 Other researchers study the effect of the frequency of forecast revision, which can be interpreted as both the frequency of the forecast model deployment and the frequency of model training.
-The frequency of retraining is studied by Spiliotis and Petropoulos [16], where the authors investigate different scenarios of updating the model fit for univariate exponential smoothing (ES) and univariate gradient boosting models (LightGBM).
+The frequency of retraining is studied by Spiliotis and Petropoulos [2024], where the authors investigate different scenarios of updating the model fit for univariate exponential smoothing (ES) and univariate gradient boosting models (LightGBM Ke et al.
+[2017]).
 We are particularly interested in studying the model deployment frequency.
@@
 In control theory, a widespread strategy for tackling online multi-step optimization challenges is the employment of the Receding Horizon Control (RHC) algorithm.
-Model Predictive Control (MPC) is the most common implementation of RHC that employs a dynamic model of the process and is widely used in online optimization.
 When rerun at every time step, this method commits to the first step of the future horizon while treating later decisions as advisory.
@@
 Throughout this paper, we will use the term FHC to align with the focus on commitment-specific optimization.
+In literature, the FHC/RHC algorithm is often referred to as Model Predictive Control (MPC), which is a well-established method for solving online multi-step optimization problems in control theory.
 (a) FHC
@@
 Figure 1: Diagrams of the Online Optimization algorithms. RHC: Receding Horizon Control, FHC: Fixed Horizon Control
-The research conducted by Chen et al.
-[17] highlights the challenges posed by noisy predictions in designing Online Convex Optimization (OCO) algorithms.
-Several solutions, such as Averaging Fixed Horizon Control (AFHC) and its generalized counterpart, Committed Horizon Control (CHC) [18], propose aggregation strategies over multiple revisions of plans to achieve sublinear regret and improve robustness to prediction errors.
-Similarly, the Feasible Fixed Horizon Control (FFHC) algorithm [19] incorporates a regularization term to handle multi-interval optimization problems while ensuring feasibility and robustness against ramping costs.
-While these approaches demonstrate the benefits of aggregating policies over prediction updates, the current work shifts focus to understanding the effects of commitment on forecast stability and downstream optimization in the presence of switching costs.
-Unlike these aggregation-based methods, this study seeks to explore how the commitment level impacts system performance, with a particular emphasis on the relationship between forecast stability and switching costs.
-Forecast stability has been recognized as a critical factor in planning.
-Building on the comprehensive framework by Godahewa et al.
-[20], which categorizes stability into vertical (variations across forecast updates for the same period) and horizontal (variations within a prediction window), this work investigates these dimensions in the context of energy management systems.
-Unlike supply chain management, where freezing intervals [21, 22] and smoothing techniques [23] have been widely used to address system nervousness and the bullwhip effect [24], energy systems require more dynamic and frequent forecast updates.
-Switching costs are also observed in energy management systems, such as ramping costs for generating units [7, 25] or toggling server load balancing [8, 26].
-However, the literature lacks a systematic exploration of how commitment levels influence the interplay between forecast stability and switching costs.
-By focusing on these aspects, the current research sets the stage for advancing the design and operation of energy management systems, particularly in applications involving energy storage optimization and real-time decision-making.
+The challenge of noisy predictions in Online Convex Optimization (OCO) has been extensively studied, notably by Chen et al.
+[2015], who emphasized the difficulty of designing robust online algorithms under uncertainty.
+In response, various strategies have emerged to mitigate the impact of prediction noise.
+For instance, Averaging Fixed Horizon Control (AFHC) and its generalized form, Committed Horizon Control (CHC) Chen et al.
+[2016], employ aggregation across multiple plan revisions to achieve sublinear regret and greater robustness.
+Likewise, the Feasible Fixed Horizon Control (FFHC) algorithm Christianson et al.
+[2022] introduces regularization to ensure feasibility in multi-interval settings while addressing ramping-related switching costs.
+In this study, we adopt the classical Fixed Horizon Control (FHC) algorithm with a fixed commitment level as the foundational optimization scheme.
+FHC is not only conceptually simple but also widely adopted in energy management systems, particularly in implementations of Model Predictive Control (MPC) Ahmad et al.
+[2023].
+While previous works have primarily focused on aggregating or averaging policies to improve performance, our focus shifts toward examining how the frequency of re-optimization, i.e., the commitment level, interacts with forecast stability and switching costs.
+Forecast stability itself has recently gained attention as a key dimension in sequential decision-making.
+Building on the taxonomy proposed by Godahewa et al.
+[2023], which distinguishes between vertical stability (variation across forecast updates for the same time step) and horizontal stability (variation within the forecast horizon at a single time step), we explore how these notions influence system performance in the energy domain.
+Though similar concerns have been addressed in supply chain literature, where freezing intervals Sahin et al.
+[2013], Xie et al.
+[2003] and smoothing techniques Ho* [2005] aim to reduce system nervousness and the bullwhip effect Wang and Disney [2016], energy systems impose different requirements.
+Forecasts must be updated more frequently and switching costs (such as ramping costs Wu et al.
+[2013], Tanaka [2006] or toggling server states Lin et al.
+[2012], Lee et al.
+[2021]) are often tied to physical infrastructure.
+Despite the relevance of these factors, few studies have systematically explored the link between commitment strategies, forecast stability, and switching costs in energy systems.
+This study addresses that gap by investigating how varying commitment levels affect performance, particularly in energy storage scheduling and real-time optimization contexts.
+By doing so, we aim to inform the design of more robust and responsive decision-making frameworks in energy management.
 1.2 Contents and Contributions
 The principle contributions are as follows:
-  • 1.
-    First, we formally define the problem and the metric used for evaluating the downstream performance, both with and without switching costs.
-We also conduct a theoretical analysis of the performance bounds of the Fixed Horizon Control (FHC) algorithm, highlighting how the commitment level influences the trade-off between forecast accuracy and switching costs.
-  • 2.
-    The effect of forecast accuracy and stability on the downstream performance of the optimization policy is analyzed.
-We extend the definition of forecast stability towards probabilistic forecasts and propose a metric for measuring horizontal and vertical stability of scenario sets.
-  • 3.
-    We conduct an empirical evaluation of the stability of point forecasts and probabilistic scenario sets, using deterministic and stochastic optimization approaches, respectively.
-We use the case study of the integrated forecasting and optimization energy management problem from the Citylearn 2022 competition.
-Based on the case study, we analyze the relationship between forecast accuracy, stability and the performance of the policy.
-We show that running the policy for a longer period provides a simple mechanism to avoid switching costs while maintaining or improving the performance.
-  • 4.
-    Lastly, the paper concludes with a discussion of the general implications of the results for the design of the energy management systems followed by suggested future research directions.
-2 Methodology
-2.1 Problem statement
-Figure 2: Grid diagram illustrating the coordinated battery scheduling problem.
-This paper addresses the challenge of optimizing battery storage management within a multi-agent energy management system.
-The setting of the problem is first outlined in the CityLearn Challenge 2022 [27] and based on a real-world dataset from a grid-interactive community.
-The CityLearn Challenge is an annual competition that focuses on the development of reinforcement learning agents for energy management in buildings.
-Participants are provided with a simulated environment that mimics real-world energy consumption and generation patterns.
-The goal is to optimize the operation of energy systems, such as batteries and HVAC systems, to minimize costs and emissions while maintaining comfort levels.
-The series of CityLearn competitions, starting from 2020, was originally set up only for Reinforcement Learning (RL) solutions, where an agent learns to make decisions by performing actions and receiving feedback from the environment.
-However, from 2022 onwards, alternative methods such as Model Predictive Control (MPC), which uses a model to predict the future behavior of a system and optimize the control actions accordingly, and Rule-Based Control (RBC), which involves using predefined rules to make decisions based on the current state of the system, have also been employed.
-The competition provides a platform and the open-source simulation environment [^1] https://github.com/intelligent-environments-lab/CityLearn for implementing the operation of energy management systems with setups that include various energy assets, such as solar panels, batteries, heat pumps, connections to the grid, and electric vehicles.
-The setup of the investigated problem is illustrated in Figure Figure 2.
-In general terms, we consider an online optimization problem with switching costs.
-The data of the competition has the hourly resolution.
-The core problem is multi-agent scheduling of energy storage over a period of a year.
-While the problem can be formulated similarly to RL, or a dynamic programming approach, in this paper, we particularly investigate the mechanics with a rolling-horizon approach.
-This means that the problem is solved iteratively over a planning horizon $H$.
-However, it is typical for only first $v$ steps to be implemented with the rest of the prediction horizon $H-v$ to be advisory, since the optimization is revised with a period of $v$ steps.
-The full details of the objective score and relevant constraints are provided in A.1.
-The problem includes finding an optimal state-dependent policy $X_{t}^{FHC}(S_{t}|\theta^{LA})$.
-The decision-maker is faced with a sequence of decisions, where the decision at each step is based on the current state $S_{t}$ of the system and the direct lookahead approximation $\theta^{LA}$.
-Using the framework proposed by Powell [28], we use the following formulation:
-$X_{t}^{FHC}(S_{t}|\theta^{LA})=\arg\min_{x_{t}}\Biggl{(}\sum_{i\in n}W_{i}\cdot\mathbb{C}(\textbf{x};\textbf{S})\Biggr{)}=$
+  • First, we formally define the problem and the framework for online optimization with switching costs, introducing the Fixed Horizon Control (FHC) algorithm and the concept of forecast stability.
+We develop the theoretical analysis that highlights how the commitment level influences the trade-off between forecast accuracy and switching costs.
+  • We propose a novel metric called Scenario Distribution Change (SDC) for evaluating the stability of probabilistic scenario sets, extending the concept of forecast stability from point forecasts to probabilistic forecasts.
+This metric allows for measuring both horizontal stability (within a forecast horizon) and vertical stability (across forecast updates).
+  • We conduct an empirical evaluation using the integrated forecasting and optimization energy management problem from the Citylearn 2022 competition.
+Through extensive testing of different commitment periods, we analyze the relationship between forecast accuracy, stability, and policy performance, demonstrating that in the presence of switching costs, longer periods of commitment can lead to improved decision-making and system performance.
+  • Lastly, we discuss the implications of our findings for the design of energy management systems, highlighting how the balance between forecast accuracy, stability, and switching costs shapes optimal operational strategies in real-world applications.
+2 Problem Formulation and Framework
+This section begins by introducing a general problem framework for online optimization with switching costs, defining the mathematical formulation and related key concepts.
+Next, we present the Fixed Horizon Control algorithm as the primary solution method, discussing how commitment periods affect decision-making.
+The concept of forecast stability is also introduced to analyze its impact on system performance.
+2.1 Online Optimization with Switching Costs
+Online optimization with switching costs is a class of problems characterized by sequential decision-making under uncertainty.
+In these problems, decisions must be made in real-time as new information becomes available, with only partial knowledge of future conditions.
+What makes this class of problems particularly challenging is the presence of time-coupling effects, where decisions at one time step directly affect the state of the system and the options available at subsequent time steps.
+The general problem is defined as finding an optimal state-dependent policy $X_{t}(S_{t}|\theta^{LA})$.
+The decision-maker faces a sequence of decisions, where each decision is based on the current system state $S_{t}$ and the direct look-ahead approximation $\theta^{LA}$.
+Using the framework proposed by Powell [2019], the problem can be formulated as follows:
+$X_{t}(S_{t}|\theta^{LA})=\arg\min_{x_{t}}\Biggl{(}\sum_{i\in n}W_{i}\cdot\mathbb{C}(\textbf{x};\textbf{S})\Biggr{)}=$
 $\arg\min_{x_{t+1},\ldots,x_{t+H}}\Biggl{(}\sum_{t^{\prime}=t+1}^{t+H}\sum_{i\in n}W_{i}\cdot\Bigl{[}\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ (1)
-$+\beta\left\lVert\tilde{y}_{tt^{\prime}}-\tilde{y}_{t(t^{\prime}-1)}\right\rVert\Bigr{]}\Biggr{)}$
+$+\beta\left\lVert\tilde{S}_{tt^{\prime}}-\tilde{S}_{t(t^{\prime}-1)}\right\rVert\Bigr{]}\Biggr{)}$
 In this formulation:
-  • 1.
-    $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharging $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$.
-This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
-  • 2.
-    $\mathbb{C}(\textbf{x};\textbf{S})$ is the cost function, which is a function of actions $\textbf{x}=x_{t},...x_{t+H}$ and states $\textbf{S}=S_{t},...S_{t+H}$.
-  • 3.
-    $W_{s}$ Weight factor for scenario $s$, indicating the importance of different scenarios in the decision-making process.
-We assume that the weight factors are equal for all scenarios.
-In the deterministic case, the number of scenarios is equal to 1.
-  • 4.
-    $\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ Represents the cost of the set of decisions $x_{t+1},...,x_{t+H}$ given the future forecast scenarios generated by the model $\theta^{LA}_{i,tt^{\prime}}$
-  • 5.
+  • $\mathbb{C}(\textbf{x};\textbf{S})$ denotes the total cost function, which is a function of the decision variables x and the system state S
+  • $W_{i}$ represents the weight associated with different cost components or scenarios.
+For deterministic optimization, only one scenario exists, while stochastic approaches consider multiple weighted scenarios.
+  • $\mathit{h}(x_{tt^{\prime}};\theta^{LA}_{i,tt^{\prime}})$ captures the direct operational cost at time $t^{\prime}$ when following decision $x_{tt^{\prime}}$ under forecast scenario $\theta^{LA}_{i,tt^{\prime}}$.
 […]
-Optimized battery schedules are evaluated using a set of Key Performance Indicators (KPIs).
-The optimization targets the minimization of the equally weighted sum of the normalized electricity cost $C$ and carbon emissions $G$.
-When grid-related metrics are included, the average includes the grid score $D$, which is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$.
-The grid score, $D$, introduces switching costs in the optimization by accounting for the ramping costs associated with frequent changes in forecasting.
-Therefore, in order to provide a comprehensive analysis of the impact of switching costs on the optimization, the combined optimization scores are defined with and without the grid score:
-Average Score $=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (2)
-Average Score (with switching costs) $=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$ (3)
-All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario.
-The full formulation of the cost function, along with the definition of each of the KPIs and the formulation of the MPC problem is given in A.2.
-Furthermore, In A.3, we conduct a theoretical analysis of the performance bounds on the solving algorithm in an online optimization problem with switching costs.
-Following the line of proof in Chen et al.
-[18], we derive the upper bound on the competitive difference between the FHC algorithm and the optimal policy.
-The analysis is conducted for the deterministic case, where the forecast is a point forecast, then extended to the probabilistic case, where the forecast is represented as a set of scenarios.
-We discuss the role of policy commitment and draw a parallel to stability.
-A trade off is demonstrated between forecast accuracy and switching costs, with the commitment level regulating the balance between the two.
-2.2 Optimization
-The problem is solved using the Fixed Horizon Control (FHC) algorithm, which is a configuration of model predictive control (MPC) with a fixed commitment horizon.
-The optimal policy is computed for the current state of the system and the forecast for the next $H$ time steps.
-An online optimization algorithm typically determines the optimal actions $x_{t+1},...,x_{t+H}$ given a prediction window of length $H$.
+2.2 Fixed Horizon Control Algorithm
+To address online optimization problems with switching costs, the Fixed Horizon Control (FHC) algorithm is employed.
+FHC is a variant of model predictive control with a fixed commitment horizon.
+At each decision point, the algorithm computes an optimal policy for the current state of the system and forecasts for the next $H$ time steps, determining the optimal actions $x_{t+1},...,x_{t+H}$ given a prediction window of length $H$.
+A key parameter in FHC is the commitment period $v$, which defines how many steps of the computed policy are implemented before recomputing.
 Every $v$ time steps, the optimization is rerun to generate a new plan.
-This algorithm is referred to as a Fixed Horizon Control (FHC) algorithm.
-In the context of the online optimization algorithms, the term “commitment” refers to the duration or number of time steps for which the optimized policy is committed to before the optimization process is rerun to generate a new plan.
-This concept plays a crucial role in determining how frequently the optimization is updated based on new information, accumulation of forecast errors and changes in the system’s state.
-An FHC algorithm with the shortest commitment is the Receding Horizon Control (RHC) algorithm.
-In RHC, the commitment is $v=1$, indicating that the optimization is rerun at every time step.
-Similar to the optimization, the forecast is also run in a rolling origin fashion.
-The optimal policy is then applied to the system for the next $v_{F}$ time steps and the process is repeated every $v_{F}$ steps.
-In summary, both the forecast and the optimization have the same rolling origin setup and need to be updated at certain intervals while committing to $v$ steps.
-Ultimately, at each time step, the decision-maker is faced with a decision to either reuse the current plan or to revise it.
+When $v=1$, the algorithm becomes Receding Horizon Control (RHC), where optimization is performed at every time step, implementing only the first step of each plan.
+Both the forecast and optimization follow a rolling origin setup and require updates at certain intervals while committing to $v$ steps.
+At each time step, the decision-maker faces a choice to either reuse the current plan or to revise it.
 Similarly, the forecast can be reused or revised.
-The resulting decision matrix is shown in Table 1.
+This decision framework is illustrated in Table 1.
+It is important to note that not all quadrants in this decision matrix are equally practical or beneficial.
+Specifically, updating the forecast without updating the optimization plan (bottom-left quadrant) offers limited value, as new information is obtained but not acted upon.
+Conversely, updating the optimization without updating the forecast (top-right quadrant) can be valuable in situations where the system state changes in ways unrelated to the forecast variables.
+In most practical implementations, the most logical configurations are either reusing both forecast and plan (top-left) or revising both simultaneously (bottom-right).
 Table 1: Decision Matrix illustrating the possible actions based on the decisions made in the forecasting and optimization stages.
@@
 forecasting & plan
-2.3 Forecast
-The target variable for forecasting is the aggregate uncontrollable load of the energy community in Figure 2.
-Following the solution implemented in the winning submission of the competition, the forecasts are generated with gradient boosting regression trees and the linear least squares regression model.
-The gradient boosting implementation is LightGBM [29].
-Predictions are made directly for the next 24 hours with 24 independent models.
-The scenario sets are generated with added gaussian noise.
-The details of the forecast models and the generation of the scenario sets are outlined in the competition paper by Nweye et al.
-[27] and the code repository of the winner solution is available on GitLab [30].
-The evaluation conducted in this study is available in the code repository [31].
-While typical forecasting models use a building’s own historical data for training, in this case, we apply forecasts to out-of-sample buildings using a model trained on an isolated set of buildings.
-This setup follows the problem setting given in the Citylearn Challenge 2022 [27].
-Although this train-test splitting is not typical in applied tasks, the setup is still relevant to real-world scenarios, as the lack of historical data within the test subset is compensated by the availability of perfect weather forecast and similarity between the train and test sets.
 2.2.1 Rolling Origin Window
-In practical applications, the forecast is issued periodically for a finite number of future time steps.
-This process is run iteratively in a rolling origin fashion, where the forecast is updated every $v_{f}$ time steps.
-A schematic of this process is shown in Figure 3.
-This approach involves continuously moving the time window forward by a certain step size after each forecasting iteration.
-For example, if we have a window size of 3 and a step size of 1, we would first generate the predictions and optimize for the steps 1-3, then at a second iteration, the forecasts are issued for the steps 2-4.
-Figure 3: Updating Rolling Origin Window. The forecast
-is updated every $v_{f}$ time steps and possible overlaps between prediction windows are shown.
-The rolling horizon setup enables an update of the prediction using the most recent information.
-When the commitment period is shorter than the forecast horizon, the regular updates create overlaps between prediction windows, when the forecasts associated with a particular target time in the future are released in multiple revisions.
-The resulting overlapping interval is highlighted in light grey in Figure 3.
-These overlaps are considered when evaluating stability of the forecast.
+In practical applications, forecasts are issued periodically for a finite number of future time steps.
+This process runs iteratively in a rolling origin fashion, where the forecast is updated every $v_{f}$ time steps.
+Figure 2 illustrates this approach, which involves continuously moving the time window forward by a certain step size after each forecasting iteration.
+  Input Data
+ $\rightarrow$ Forecast Horizon
+$t_{0}$$t_{1}$$t_{2}$$t_{3}$$t_{4}$$t_{5}$$t_{6}$$t_{7}$$t_{8}$orecast1Forecast2Forecast3Forecast4Forecast Origins
+Figure 2: Rolling Origin Forecasting with Fixed Input Window and Forecast Horizon.
+The diagram illustrates how forecasts with horizon $H=2$ are updated every $v_{f}=1$ time steps, creating overlaps between prediction windows.
+The rolling horizon setup enables prediction updates using the most recent information.
+When the commitment period is shorter than the forecast horizon, regular updates create overlaps between prediction windows, as forecasts for a particular future time point are released in multiple revisions.
+These overlaps are considered when evaluating forecast stability.
 2.2.2 Stability
-Stability of the forecast is defined as the consistency of the predictions over time.
-Following the categorization, first proposed in [20], we distinguish vertical and horizontal stability.
-The difference between the two concepts is illustrated in Figure 4.
+Forecast stability is defined as the variability of predictions over time.
+Following the categorization proposed by Godahewa et al.
+[2023], we distinguish between vertical and horizontal stability, as illustrated in Figure 3.
 (a) Vertical Stability
@@
 Figure 3: Diagrams demonstrating the concepts of vertical and horizontal stability between predictions.
-The vertical stability is the difference between the predictions produced at different origins for the same time period.
-If the forecast revisions are vertically stable, the corresponding adjustment costs in policy are set to be minimal.
-Horizontal stability characterizes the difference between predictions generated at the same origin.
-The horizontally stable forecasts are meant to reduce fluctuations between time steps within a prediction window.
-Similarly to the evaluation in [32], a stability metric is measured for point forecasts using the mean absolute change (MAC).
-The metric measures the change between the predictions issued at different origins — ${MAC}_{V}$, or the change of predictions within the horizon window generated at the same origin — ${MAC}_{H}$
-$\operatorname{MAC}_{V}=\frac{\sum_{i=1}^{H-1}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i\mid t-1}\right|}{H-1}$ (4)
-$\operatorname{MAC}_{H}=\frac{\sum_{i=2}^{H}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i-1\mid t}\right|}{H-1}$ (5)
+Vertical stability refers to the consistency between predictions made at different origins for the same target time period.
+Horizontal stability, on the other hand, characterizes the variance of predictions within the same forecast window.
+3 Theoretical analysis of performance bounds
+This section presents an analysis of the performance bounds for Fixed Horizon Control (FHC) algorithms with commitment level $v$.
+We derive upper bounds on the competitive difference, the gap between FHC and the optimal clairvoyant policy.
+Our analysis reveals a fundamental trade-off between forecast accuracy and switching costs that depends critically on how frequently policies are updated.
+We first develop these bounds in a deterministic setting, then extend them to stochastic optimization where forecasts are represented as scenario ensembles.
+Finally, we establish the theoretical connection between forecast stability and switching costs, demonstrating why stable forecasts can significantly improve performance in environments with high switching costs.
+We adopt the framework introduced by Chen et al.
+[2016] for analyzing online convex optimization (OCO) problems with switching costs and noisy predictions.
+In this setting, a decision-maker sequentially selects actions $x_{t}\in\mathcal{F}\subset\mathbb{R}^{n}$ over a time horizon $T$, facing unknown convex cost functions $h_{t}(x_{t})$ and incurring a switching cost $\beta\|S_{t}-S_{t-1}\|$.
+$\operatorname{cost}(ALG)=\sum_{t=1}^{T}h_{t}(x_{t})+\sum_{t=1}^{T}\beta\|S_{t}-S_{t-1}\|$ (2)
+The true state $y_{t}$ at each time $t$ is unknown at the time of decision-making and must be predicted.
+Forecasts $\hat{y}_{t|\tau}$ are generated at time $\tau<t$ and are corrupted by noise, modeled as:
+$y_{t}-\hat{y}_{t|\tau}=\sum_{s=\tau+1}^{t}f(t-s)e(s),$ (3)
+where $e(s)$ are i.i.d. zero-mean noise terms with covariance matrix $R_{e}$, and $f(\cdot)$ is a deterministic impulse response function capturing the correlation structure of the forecast errors.
+The performance of an online algorithm is assessed using the competitive difference, defined as the expected cost gap between the online algorithm and the offline optimal (oracle) policy.
+Formally, an online algorithm ALG, is said to have a competitive difference of $\rho(T)$ if:
+$\sup_{\hat{y}}\mathbb{E}_{e}\left[\operatorname{cost}(\text{ALG})-\operatorname{cost}(\text{OPT})\right]\leq\rho(T)$ (4)
+where the expectation is taken with respect to the prediction noise sequence $\{e(t)\}_{t=1}^{T}$.
+The offline optimal policy OPT has full access to future outcomes and is thus not affected by forecast errors.
+In contrast, the performance of online algorithms like Fixed Horizon Control (FHC) or Receding Horizon Control (RHC) is directly impacted by the structure and magnitude of the forecast noise.
+Following this framework, we derive bounds on the expected performance degradation of FHC-type algorithms as a function of the commitment level $v$, the forecast error correlation $\|f_{v}\|$, and the switching cost penalty $\beta$.
+As a starting point, we refer to Theorem 1 in Chen et al.
+[2016], and adapt it to our context.
+$\operatorname{cost}(FHC(v))\leq\operatorname{cost}(OPT)+2M_{k}\beta D+2G\sum_{t=1}^{T}\|S_{t}-S_{t\mid\phi_{k}(t)}\|$ (5)
+where $M$ is the number of updates, $k$ is the commitment level, $D$ is the maximum switching cost, and $G$ is the Lipschitz constant of the cost function
+Here, in contrast to the original paper, we derive the expectation over the forecast horizon for the FHC algorithm with commitment level $v$:
+$\mathbb{E}[\operatorname{cost}(FHC(v))]$ $\leq\operatorname{cost}(OPT)+2M_{k}\beta D+2G\sum_{t=1}^{T}\mathbb{E}\left[\|S_{t}-S_{t\mid t-k}\|\right]$ (6)
+$\leq\operatorname{cost}(OPT)+2(T/v)\beta D+2GT\cdot\|f_{v}\|^{\alpha}$ (7)
+as $M_{k}$ is the number of optimization updates, which is equal to $T/v$ for the FHC(v) algorithm.
+$\mathbb{E}\operatorname{cost}(FHC)\leq$ (8)
+$\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2G\mathbb{E}\sum_{t=1}^{T}\left\|S_{t}-S_{t\mid t-\phi^{1}(t)}\right\|_{2}^{\alpha}$
+$\mathbb{E}\operatorname{cost}(FHC)\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2GT\left\|f_{v}\right\|^{\alpha}.$ (9)
+The term $\left\|f_{v}\right\|^{\alpha}$ is the $\alpha$-norm of the prediction error covariance, where $\alpha$ is the exponent of the Hölder condition.
+We assume $\alpha\geq 1$ to ensure that the cost function is convex.
+3.1 Deterministic Optimization
+In practical applications, forecast errors often exhibit exponentially decaying temporal correlations.
+We model this behavior using a decay parameter $a\in(0,1)$ and define the Frobenius norm of the correlation impulse response as:
+$\|f(s)\|_{F}=\begin{cases}ca^{s},&s\geq 0\\ 0,&s<0\end{cases}$ (10)
+where $c$ is a constant scaling factor and $s$ is the time lag between the forecast target time and forecast issuance.
+We assume zero-mean i.i.d.
+noise vectors $e(s)$ with covariance $\mathbb{E}[e(s)e(s)^{T}]=R_{e}$ and trace $\operatorname{trace}(R_{e})=\sigma^{2}$.
+The cumulative forecast error norm up to commitment level $v$ becomes:
+$\left\|f_{v}\right\|^{2}$ $=\sum_{s=0}^{v}\operatorname{trace}(R_{e}f(s)^{T}f(s))=\sum_{s=0}^{v}\|R_{e}^{1/2}\|_{F}^{2}\cdot\|f(s)\|_{F}^{2}$
+$=\sum_{s=0}^{v}c^{2}\sigma^{2}a^{2s}=c^{2}\sigma^{2}\cdot\frac{1-a^{2(v+1)}}{1-a^{2}}$
+Taking the square root and using the Lipschitz continuity of $h$ with constant $G$, we have:
+$\|f_{v}\|=c\sigma\sqrt{\frac{1-a^{2(v+1)}}{1-a^{2}}}$ (11)
+Substituting into the expected cost bound for the FHC algorithm, we obtain:
+$\mathbb{E}[\operatorname{cost}(FHC(v))]\leq\mathbb{E}[\operatorname{cost}(OPT)]+\frac{2T\beta D}{v}+2GT\|f_{v}\|$ (12)
+Define:
+$A$ $=2T\beta D$
+$B$ $=\frac{2GTc\sigma}{\sqrt{1-a^{2}}}$
+Then the bound becomes:
+$\mathbb{E}[\operatorname{cost}(FHC(v))]\leq\mathbb{E}[\operatorname{cost}(OPT)]+\frac{A}{v}+B\cdot\sqrt{1-a^{2(v+1)}}$ (13)
+This formulation clearly reflects the trade-off.
+The term $\frac{A}{v}$ is the cost of policy switching and decreases with larger commitment $v$.
+The term $B\cdot\sqrt{1-a^{2(v+1)}}$ captures forecast error accumulation and increases with $v$ since $a^{2(v+1)}\to 0$ as $v$ increases.
+The role of each parameter is summarized in Table 2.
+Therefore, the expected competitive difference exhibits a U-shaped curve with respect to $v$, where a minimal cost is achieved by balancing the cost of frequent switching and the degradation due to forecast staleness.
+This quantifies the intuition that updating too frequently incurs high switching costs, while updating too infrequently leads to performance loss due to outdated forecasts.
+This function run with different sets of parameters is demonstrated in Figure 7.
+It is observed that the low switching costs lead to a linear increase in the expected competitive difference, while the high switching costs lead to a formation of a U-shaped curve.
+Table 2: Key parameters in performance bound
+Symbol Represents Interpretation
+$A$ Switching Cost Weight Scales with horizon $T$, penalty $\beta$, and max switching distance $D$. Larger $A$ increases the switching cost.
+$B$ Forecast Error Sensitivity Incorporates cost weight $G$, horizon $T$, and noise parameters. Amplifies penalty for forecast inaccuracy.
+$a$ Correlation Decay Rate Controls how quickly forecast errors become uncorrelated over time. Lower values of $a$ mean errors at future steps are largely independent.
+Figure 4: Illustration of the FHC expected competitive difference as a function of the commitment level for (a) Low and High switching costs with predictions of exponentially decaying correlations, and (b) Stochastic and Deterministic optimizations.
+Stochastic scenarios exhibit a lower level of forecast error sensitivity ($B_{s}<B$), shifting the optimal commitment level to higher values.
+3.2 Stochastic Optimization over Scenarios
+We now analyze the performance of Fixed Horizon Control (FHC) under a stochastic optimization setting where forecasts are provided as a set of $n$ equiprobable scenarios, denoted $\{\xi_{i}\}_{i=1}^{n}$, each representing a possible trajectory of future values over the horizon $v$.
+For each scenario $\xi_{i}$, the FHC algorithm yields a policy $x^{(i)}=(x^{(i)}_{t})_{t=1}^{T}$.
+The overall cost is computed by averaging over the scenario outcomes:
+$\mathbb{E}\left[\operatorname{cost}(FHC_{\text{stochastic}}(v))\right]=\frac{1}{n}\sum_{i=1}^{n}\mathbb{E}\left[\operatorname{cost}(FHC(v,\xi_{i}))\right].$ (14)
+Let $f_{v}^{(i)}\in\mathbb{R}^{d}$ denote the cumulative forecast error over horizon $v$ under scenario $\xi_{i}$.
+The expected cost includes a term of the form $\|f_{v}^{(i)}\|^{\alpha}$, which contributes to the second term in the performance bound.
+Assuming $\alpha\geq 1$, the function $x\mapsto\|x\|^{\alpha}$ is convex, and Jensen’s inequality yields:
+$\left\|\frac{1}{n}\sum_{i=1}^{n}f_{v}^{(i)}\right\|^{\alpha}\leq\frac{1}{n}\sum_{i=1}^{n}\left\|f_{v}^{(i)}\right\|^{\alpha}.$ (15)
+This inequality implies that the effective contribution of forecast uncertainty to the cost is lower in the stochastic case than in the deterministic case, since:
+$2GT\left\|\mathbb{E}[f_{v}^{(i)}]\right\|^{\alpha}\leq 2GT\cdot\frac{1}{n}\sum_{i=1}^{n}\left\|f_{v}^{(i)}\right\|^{\alpha}.$ (16)
+Thus, scenario averaging introduces a variance reduction effect, which improves robustness to forecast noise. This insight leads to the following bound:
+$\mathbb{E}\left[\operatorname{cost}(FHC_{\text{stochastic}}(v))\right]\leq\mathbb{E}\left[\operatorname{cost}(OPT)\right]+\frac{A}{v}+B_{s}(1-a^{2(v+1)})^{1/2},$ (17)
+where $B_{s}<B$ captures the effect of scenario averaging.
+This result highlights a key theoretical benefit of stochastic FHC: the aggregation over multiple forecast paths leads to a lower effective forecast variance and therefore lower forecast error sensitivity.
+It is shown in Figure 4(b), that the vertex of the U-shaped curve shifts to higher values of $v$ in the stochastic case.
+The fundamental trade-off between switching costs (decreasing with $v$) and forecast accuracy (typically decreasing with more frequent updates) persists, but the stochastic case enables a longer commitment horizon before the forecast error term dominates.
+3.3 Role of Forecast Stability in Optimization Performance
+The theoretical performance bounds in Equation (9) consist of two main components: (i) a term proportional to the switching cost, $\frac{2T\beta D}{v}$, and (ii) a forecast error penalty, $2GT\|f_{v}\|^{\alpha}$.
+While the second term captures the accuracy of the forecast over the horizon $v$, the first term reflects the cost incurred by changing policies across planning updates.
+In this section, we formalize the connection between the stability of forecasts and the incurred switching costs.
+Let $x_{t}$ denote the decision policy applied at time $t$, and let this policy be a deterministic function of the forecast $\hat{y}_{t}$, such that:
+$x_{t}=\mathcal{M}(\hat{y}_{t}),$ (18)
+where $\mathcal{M}:\mathbb{R}^{H}\to\mathbb{R}^{d}$ is a policy generation function (e.g., the solution to an optimization problem over horizon $H$), and is assumed to be Lipschitz continuous with constant $L_{\mathcal{M}}$:
+$\|\mathcal{M}(\hat{y}_{t})-\mathcal{M}(\hat{y}_{t-v})\|\leq L_{\mathcal{M}}\cdot\|\hat{y}_{t}-\hat{y}_{t-v}\|.$ (19)
+Using this, the switching cost term can be bounded as:
+$\sum_{t=1}^{T/v}\beta\|S_{t}-S_{t-v}\|\leq\beta L_{\mathcal{M}}\sum_{t=1}^{T/v}\|\hat{y}_{t}-\hat{y}_{t-v}\|.$ (20)
+The term $\|\hat{y}_{t}-\hat{y}_{t-v}\|$ represents the change in forecast between updates, which corresponds to the concept of vertical forecast stability—the consistency of predictions made at different time origins for the same target time.
+A common empirical proxy for this is the Mean Absolute Change (MAC) metric:
+$\operatorname{MAC}_{V}=\frac{1}{H-1}\sum_{i=1}^{H-1}|\hat{y}_{t+i|t}-\hat{y}_{t+i|t-v}|.$ (21)
+Thus, improving the vertical stability of forecasts directly reduces the magnitude of policy changes, and therefore the incurred switching costs.
+This establishes a theoretical link between empirical measures of forecast stability and the first term in the performance bound (9).
+The MAC metric and appropriate stability metrics for the stochastic case are discussed in detail in the next section.
+4 Experimentation and Results
+In this section, we transition from theoretical analysis to practical implementation and evaluation.
+While the theoretical model provides important insights, real-world energy management systems face additional complexities including non-trivial forecast error correlations, potential non-convexity in optimization problems, and the challenges of coordinating multiple distributed assets.
+We present empirical results from applying the proposed framework to a battery scheduling problem, evaluating how different commitment periods affect system performance under varying forecast qualities.
+First, the current section outlines the experimental setup and case study, followed by a detailed analysis of the results.
+The analysis focuses on the relationship between forecast accuracy, stability, and the impact of switching costs on the optimal decisions made in the battery scheduling process.
+4.1 Application to Battery Scheduling in Energy Management Systems
+The decision-maker must balance the benefits of plan revisions against the costs of switching.
+This balance is particularly critical in energy management systems, where rapid changes in operation can strain infrastructure, reduce equipment lifespan, and lead to additional expenses or inefficiencies.
+To investigate these trade-offs in a realistic setting, we apply the concepts of online optimization with switching costs to battery storage management within a multi-agent energy management system.
+Our case study is based on the CityLearn Challenge 2022 Nweye et al.
+[2022], which provides a comprehensive dataset from an actual grid-interactive community.
+The central challenge involves multi-agent scheduling of energy storage across a one-year period with hourly granularity.
+Figure 5 illustrates the problem setup.
+The data of 17 buildings used in this study is derived from a real-world zero net energy community in Fontana, California, USA, previously studied for grid integration of zero net energy developments as part of the California Solar Initiative program Narayanamurthy et al.
+[2016].
+Each building has a battery with specific charging/discharging characteristics, and the system must coordinate these resources efficiently.
+The objective is to develop control policies that minimize grid electricity costs, carbon emissions and ramping while increasing the load factor.
+The associated ramping costs constitute the switching costs in this context.
+The combined objective score with the grid cost is defined as:
+[…]
+$\operatorname{avg}\left(\frac{C_{\text{ALG}}}{C_{\text{no\ battery}}},\frac{G_{\text{ALG}}}{G_{\text{no\ battery}}},D\right)$
+where $C$ are electricity costs, $G$ are carbon emissions, and $D$ is the grid score.
+The grid score is computed as the mean of the normalized ramping KPI $R$ and the Load Factor KPI $(1-L)$.
+Full details of the objective score and relevant constraints are provided in A.1.
+The proposed framework was implemented using the CityLearn simulation environment, an open-source platform designed for evaluating energy management systems in grid-interactive communities [^1] https://github.com/intelligent-environments-lab/CityLearn.
+Originally created for reinforcement learning (RL), the platform now also supports alternative methods like Model Predictive Control (MPC) and rule-based control.
+It simulates a range of energy assets, including batteries, solar panels, heat pumps, and electric vehicles, all of which play a role in optimizing energy consumption across a grid-connected community.
+Figure 5: System setup (a) and data flow (b) for the battery scheduling problem.
+The data flow diagram illustrates how input features are processed through forecasting models with varying update frequencies to generate load predictions, which then inform the optimization stage to produce battery control actions.
+4.2 Implementations of Forecasting and Optimization
+The system architecture consists of two primary components: forecasting and optimization, as shown in Figure 5(b).
+The forecasting module generates load predictions, which serve as inputs to the optimization model.
+Both forecasting and optimization follow a rolling horizon approach.
+While the commitment period $v$ is commonly the same for both stages, updates are not inherently required to be synchronized.
+Therefore, we distinguish between $v_{F}$ (forecast update frequency) and $v_{O}$ (optimization update frequency) when necessary.
+For the forecasting task, we predict the aggregate uncontrollable load of the energy community.
+The training set comprises data of 5 buildings, with the remaining 12 buildings used for validation and testing.
+The forecasts are generated using a combination of gradient boosting regression trees (LightGBM Ke et al.
+[2017]) and linear least squares regression.
+The method follows the procedure of the winning solution, outlined in Nweye et al.
+[2022].
+We train the forecasting model across all buildings, then apply it to 7 out-of-sample buildings for testing, a practice commonly used in scenarios with limited historical data but similar building characteristics, as discussed in Grabner et al.
+[2023].
+The forecasts are generated for a 24-hour horizon, with 24 separate models predicting each hour.
+To account for forecast uncertainty, we generate multiple forecast scenarios by adding Gaussian noise to the base predictions.
+The optimization model, implemented with Pyomo Hart et al.
+[2017] and solved using Gurobi Gurobi Optimization, LLC [2024], aims to minimize a composite cost function that incorporates electricity costs, ramping penalties, switching costs, and carbon emissions.
+The optimization problem is solved for the same 24-hour horizon as the forecasting model, subject to revision every $v_{O}$ steps.
+At the optimization stage, both deterministic and stochastic forecast settings are considered.
+In the deterministic case, a single forecast trajectory is used to determine the optimal control policy.
+In the stochastic case, multiple forecast scenarios are considered, and the optimization is performed over 75 scenarios, minimizing the expected cost.
+This allows the system to better handle forecast uncertainty and improves robustness by optimizing expected performance across scenarios.
+The optimization and forecasting steps are coordinated through the CityLearn platform, which integrates data from multiple sources, including weather forecasts, building characteristics, and historical consumption patterns.
+The resulting control policies, which dictate how the batteries should charge and discharge, are evaluated based on their ability to reduce grid electricity costs, minimize ramping costs, and lower carbon emissions.
+The implementation is available on GitHub [^2] https://github.com/ujohn33/Predict-Optimize-Revise and is fully reproducible, with optimization logs available upon request.
+The detailed mathematical formulation of the optimization problem and the constraints imposed on the system are provided in A.2.
+4.3 Evaluation metrics
+Similarly to the evaluation in [Van Belle et al., 2023], a stability metric is measured for point forecasts using the mean absolute change (MAC).
+The metric measures the change between the predictions issued at different origins — ${MAC}_{V}$, or the variation of predictions within the horizon window — ${MAC}_{H}$
+$\operatorname{MAC}_{V}=\frac{1}{H-1}\sum_{i=1}^{H-1}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i\mid t-1}\right|$ (23)
+$\operatorname{MAC}_{H}=\frac{1}{H-1}\sum_{i=2}^{H}\left|\hat{y}_{t+i\mid t}-\hat{y}_{t+i-1\mid t}\right|$ (24)
 where $\hat{y}_{t+i\mid t}$ is the forecast for time $t+i$ generated at time $t$, $H$ is the forecast horizon, and $t$ is the current time step.
@@
 By incorporating the dimension of the metric space, SDC is conceptually analogous to the Mean Absolute Change (MAC) used for point forecasts.
-To assess stability across time, SDC evaluates both horizontal stability (within a forecast horizon) and vertical stability (across forecast updates).
-For $N$ number of scenarios, these are defined as:
+To assess stability across time, SDC evaluates both vertical stability (across forecast updates) and horizontal stability (within a forecast horizon).
+For $N$ number of scenarios and a forecasting horizon $H$, these are defined as:
+Vertical Stability (SDCV):
+$\operatorname{SDC}_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (25)
 Horizontal Stability (SDCH):
-$\operatorname{SDC}_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i\mid t-1,j}\right|$ (6)
-Vertical Stability (SDCV):
-$\operatorname{SDC}_{V}=\frac{1}{H-1}\sum_{i=2}^{H}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (7)
-2.3.4 Accuracy of Determinstic Forecasts
+$\operatorname{SDC}_{H}=\frac{1}{H-1}\sum_{i=1}^{H-1}\frac{1}{N}\sum_{j=1}^{N}\left|\hat{y}_{t+i\mid t,j}-\hat{y}_{t+i-1\mid t,j}\right|$ (26)
+4.3.2 Accuracy of Point Forecasts
 For point forecasts, the MAE is utilized to quantify the forecast accuracy.
@@
 $\operatorname{MAE}=\Biggl{(}\frac{1}{H}\sum_{i=1}^{H}\left|y_{i}-\hat{y}_{i}\right|\Biggr{)}$ (27)
-where $y_{i}$ is the actual value and $\hat{y}_{i}$ is the forecasted value at time $i$.
 4.3.3 Accuracy of Probabilistic Forecasts
@@
 $ES=\left(\frac{1}{N}\sum_{j=1}^{N}\left\|y_{i}-\hat{y}_{ij}\right\|^{p}-\frac{1}{2N^{2}}\sum_{j=1}^{N}\sum_{k=1}^{N}\left\|\hat{y}_{ij}-\hat{y}_{ik}\right\|^{p}\right)^{\frac{1}{p}}$ (28)
-where $y_{i}$ is the actual value, $\hat{y}_{ij}$ is the $j$-th scenario of the forecasted value at time $i$, $N$ is the number of scenarios, and $p$ is the order of the norm.
-The second term in the equation is the mean pairwise distance between the scenarios.
-3 Experimentation and Results
-Compared to theoretical analysis, practical setups are more complex.
-They often deal with non-trivial correlation structures in forecast errors and the potential for non-convex behaviors in optimization.
-This complexity is due to the necessity for cooperative behavior across multiple buildings.
-Additionally, various temporal dependencies emerge from switching costs and system feedback.
-To confirm the initial theoretical findings, we proceed with evaluating the performance of the FHC algorithm with different combinations of forecast and optimization commitment in a real-world scenario.
-According to the decision matrix in Figure 1, at every time step the forecast and optimization plans can be updated or retained.
-Therefore, the forecast and optimization run with commitment periods $v_{F}$ and $v_{O}$, respectively.
-We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment.
-The result is shown in Tables 2(a) and 2(b).
+4.4 Results
+According to the decision matrix in Table 1, at every time step the forecast and optimization plans can be updated or retained.
+We evaluate the performance of the FHC algorithm with different combinations of forecast and optimization commitment, $v_{F}$ and $v_{O}$, respectively.
+The result is shown in Table 3.
 The scores in the tables indicate that updating the optimization plan more frequently than the forecast offers no added benefit.
@@
 Therefore, in the following analysis, we set $v=v_{F}=v_{O}$.
-(a) deterministic FHC algorithm with point forecast
-$v_{O}$
-$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
+Table 3: Optimization KPI scores for combinations of forecast ($v_{F}$) and optimization ($v_{O}$) commitment periods between 1 and 12 hours
+Optimization Commitment Period ($v_{O}$)
+Forecast ($v_{F}$) 1 2 3 4 5 6 7 8 9 10 11 12
+Panel A: Deterministic FHC with point forecast
 1 0.899
@@
 12 0.917 0.911 0.911 0.909 0.908 0.910 0.904 0.907 0.907 0.907 0.904 0.908
-(b) stochastic FHC algorithm with probabilistic forecast
-$v_{O}$
-$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
+Panel B: Stochastic FHC with probabilistic forecast
 1 0.875
@@
 12 0.905 0.895 0.887 0.884 0.886 0.880 0.884 0.880 0.883 0.882 0.882 0.879
-Table 2: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12
-Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability.
-In Figure 5, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods between 1 and 12.
-The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings.
-Similar evaluation is done for the in-sample buildings and shown in A.4.
-It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods.
-The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
-For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours.
-The hypothesis is that this is due to the nature of the method of noise simulation, as outlined in [27].
-The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
-The plots of the optimization scores show different behavior depending on the presence of the switching costs.
-The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship.
-The optimal commitment period remains at 1 for the point forecast, while the stochastic forecast enables an optimal performance at a longer commitment period.
-This observation is notable as the most expected behavior is that best performance is achieved with shortest revision periods for both forecast types, as more up-to-date forecasts are provided.
-These observations are supported by the correlation coefficients in the tables 3(a) and 3(b) with the correlation coefficients between the forecast properties and the optimization scores.
-The accuracy metrics are in positive correlation with the optimization score.
-When switching costs are factored in, the inverse correlation between the vertical stability and the cost becomes less pronounced.
-Yet, the correlation between the horizontal stability and the cost changes from negative to positive correlation when switching costs are included.
+Note: Diagonal elements represent equal commitment periods for both forecast and optimization.
+Lower values indicate better performance.
+Empty cells (—) represent invalid combinations where $v_{O}>v_{F}$.
 Accuracy
@@
 Figure 6: Accuracy (Top), Horizontal (Second) and Vertical Stability (Third row) of stochastic forecast over different revision periods with data from out-of-sample buildings.
-(a) deterministic FHC algorithm
-with point forecast
-Correlation Coef. $MAE$ $MAC_{V}$ $MAC_{H}$
-$cost(FHC(v))$ 0.95 -0.94 -0.27
-$cost(FHC(v))$ w/ switching costs 0.70 -0.62 0.17
-(b) stochastic FHC algorithm
-with probabilistic forecast
-Correlation Coef. $ES$ $MASC_{V}$ $MASC_{H}$
-$cost(FHC_{stoch.}(v))$ 0.63 -0.86 -0.29
-$cost(FHC_{stoch.}(v))$ w/ switching costs 0.71 -0.06 0.47
-Table 3: Correlation Coefficients for Forecast Metrics and Optimization Score
-4 Discussion
-The observations from the results are consistent with the theoretical analysis of the upper bound on the competitive difference of the FHC algorithm.
-We observe that the optimal commitment period is dependent on the forecast properties.
-The inclusion of switching costs in the optimization problem changes the relationship between desired forecast properties and the optimization score.
-The forecast error is broadly lower with shorter commitment periods, however, the horizontal and vertical stability become also relevant.
-In the theoretical analysis in A.3, we see that for a more realistic model of the prediction noise with exponentially decaying correlation, a trade-off emerges between the forecast error and the switching costs.
-Through testing with real data, we observe that switching costs can be partially mitigated by committing to more stable forecasts and optimization plans.
-The addition of switching costs makes the algorithmic problem harder as it forces current actions to depend on beliefs about future cost functions.
-In Figure 6, we can see that the shape of the optimal EMS net load curve is significantly different when switching costs are included into the optimization.
+Further, we investigate the relationship between the performance of the FHC algorithm and properties of the forecast, namely accuracy and stability.
+In Figure 6, the accuracy and stability of the deterministic and stochastic forecasts are evaluated for different commitment periods between 1 and 12.
+The comparison is made between the optimization with and without switching costs for the forecasts generated for out-of-sample buildings.
+Similar evaluation is done for the in-sample buildings and shown in A.3.
+It is observed from the plots that the forecast error generally decreases with shorter commitment periods, while the vertical stability is better with longer commitment periods.
+The horizontal stability demonstrates a trade-off where the optimal level is reached at a certain commitment period between 1 and 12 hours.
+For stochastic forecast error, measured with the Energy Score, we observe an anomaly when the lowest error is observed for scenarios updated every 8 hours.
+The hypothesis is that this is due to the nature of the method of noise simulation, as outlined in Nweye et al.
+[2022].
+The magnitude of added gaussian noise is proportional to the value of the respective point prediction, which produces low variation in the regions around zero, where this specific configuration captures most accurately, therefore the resulting scenario sets capture the real distribution with a low variation.
+Nevertheless, as seen in the optimization score plots, the global minimum is reached at a shorter commitment period, indicating that updating the schedule more frequently is not only beneficial for using the most accurate forecast but also for correcting the previous decisions.
+Furthermore, it is observed that the presence of switching costs has a significant impact on the relation of performance to the commitment period.
+The optimization score without switching costs is optimal with shorter commitment periods, while the addition of switching costs adds more fluctuation to this relationship.
+Similar findings are devised from the average-case analysis in Section 3.1, where higher switching costs can potentially lead to a formation of a minimum beyond lowest commitment period.
+The optimal commitment period remains at 1 for the point forecast, while the stochastic forecast enables an optimal performance at a longer commitment period.
+This observation is notable as, conventionally, the best performance is achieved with the shortest revision periods.
+This observation is also consistent with the theoretical analysis in Section 3.2, as the stochastic optimization reduces sensitivity to forecast errors, leading to a more stable performance.
+These observations are supported by the correlation coefficients in the Table 4 with the correlation coefficients between the forecast properties and the optimization scores.
+The accuracy metrics are in positive correlation with the optimization score.
+When switching costs are factored in, the inverse correlation between the vertical stability and the cost becomes less pronounced.
+Yet, the correlation between the horizontal stability and the cost changes from negative to positive correlation when switching costs are included.
+Figure 7 demonstrates an example from the Citylearn case when the shape of the optimal EMS net load curve is significantly different when switching costs are included into the optimization.
 Without switching costs, the battery is charged and discharged while following the price and emissions cost signals, indicated in the bottom subplot.
@@
 Without considering switching costs, it is observed that scheduling with the oracle information has higher spikes.
-We assume that the oracle information given optimizing agent the confidence to charge at maximum capacity when the price and emissions are low.
+We assume that the oracle information gives the optimizing agent the confidence to charge at maximum capacity when the price and emissions are low.
 However, the forecast-based management is more conservative in its actions, as it is uncertain about the future load signals.
@@
 While it generally follows the optimal policy, it oscillates around it.
+Table 4: Correlation Coefficients between Forecast Metrics and Optimization Performance
+Point Forecast (Deterministic FHC) Probabilistic Forecast (Stochastic FHC)
+Correlation Coefficient $MAE$ $MAC_{V}$ $MAC_{H}$ $ES$ $SDC_{V}$ $SDC_{H}$
+Without switching costs 0.95 -0.94 -0.27 0.63 -0.86 -0.29
+With switching costs 0.70 -0.62 0.17 0.71 -0.06 0.47
 (a) Without switching costs
 (b) With switching costs
-Figure 6: The difference in net load between FHC with and without switching costs.
-Actual net load is shown in yellow.
-The base load is shown in green.
-The net load following the perfect forecast is shown in red.
-The electricity price is shown in blue.
-Minimal commitment periods $v$ of the control algorithm, combined with imperfect predictions, can lead to a scenario where the control strategy is constantly “chasing” the optimal state based on imperfect prediction.
-This can aggravate the impact of switching costs, as the control actions may change frequently in response to new information, even when such changes are not cost-effective in the long run.
-In line with the theoretical analysis in Appendix A.3, the upper bound on the competitive difference of the FHC algorithm is dependent on the trade-off between switching costs and forecast errors.
-While the MPC objective functions factors in the switching costs, we identify two factors that can lead to underperformance of FHC with low commitment period ($v_{O}=1$), particularly when probabilistic forecasts are used.
-Firstly, when the policy is computed for a finite period in a problem with switching cost, the past decisions are calculated as optimal with respect to future decisions given the forecast, however, after the policy is reviewed at the next iteration with an updated forecast, the optimality of the previously executed decision may be suboptimal given the updated information.
-A simple illustrative example would be the battery is discharged first with an outlook for charging in the future when the price is more favorable.
-Once the forecast and policy updates, the updated policy is discharging again, however the battery has been discharged earlier.
-The second reason is that while the objective cost function does include the term for the switching costs $\beta||{x_{t}}-x_{t-1}||$, the switching costs are computed pairwise between loads at close time instances.
-However, since the total charging capacity of batteries is finite and limited, it is advantageous to regulate switching costs within longer action windows.
-Other important observation is that the FHC algorithm with longer commitment periods only becomes viable when the optimization is computed for a set of scenarios, as we can see from the results in Figure 5.
-It is hypothesized, that the aggregation of policies over different scenarios provides a more robust policy that hedges against the forecast errors.
-Generally, within a forecast horizon, the nearest steps are predicted with higher accuracy than further steps.
-Longer commitment periods aggravate the impact of forecast errors due to time coupling, as the suboptimal decisions made in the past are amplified in the future.
-A stochastic FHC algorithm is better equipped to handle errors in the forecast, as it aggregates policies over different scenarios, therefore, the impact of forecast errors is mitigated and longer commitment periods are more viable.
+Figure 7: Net load comparison: Without switching costs, energy management with perfect forecast (yellow) generates high peaks versus conservative forecast-based control (blue).
+With switching costs, the policy smoothes fluctuations.
+The base load, being the prediction target variable, is shown in blue.
+Price and emissions signals are shown in the bottom subplot.
+5 Discussion
+This study provides theoretical and empirical evidence for the nuanced role that commitment periods play in forecast-based optimization under switching costs.
+Our results validate the theoretical performance bounds developed in Section 3, showing that the trade-off between forecast inaccuracy and switching costs emerges most prominently when switching penalties are non-negligible.
+Notably, this trade-off is not universal: in problems without meaningful switching costs, shorter commitment periods are almost always preferred.
+However, when switching costs are significant, such as in systems with physical constraints, ramping penalties, or contract-driven operational inertia, the choice of commitment level becomes a key decision variable in its own right.
+The presence of switching costs changes the optimization landscape in fundamental ways.
+It introduces temporal coupling between successive decisions, requiring current actions to anticipate the cost of future revisions.
+This often leads to smoother and more conservative policies, as illustrated in Figure 7, the addition of switching costs results in a visibly flattened net load curve.
+Such behavior is desirable in many real-world settings: battery degradation, generator ramping, thermal comfort zones, and load shifting incentives all impose implicit or explicit switching penalties that must be carefully managed.
+Although such costs are context-specific and often application-dependent, they are far from rare in modern energy systems.
+Importantly, we show that this trade-off is more tractable under stochastic optimization.
+Section 3.2 demonstrates that scenario averaging introduces a form of regularization, lowering the effective sensitivity to forecast errors.
+This effect is reflected in both the theoretical bounds and the experimental results.
+In practice, this means that higher commitment levels, normally problematic due to the accumulation of outdated forecast errors, can be made viable when robust scenario-based policies are used.
+The performance gains from longer commitment periods become visible only when switching costs are high and stochastic optimization is employed, as shown in Figure 6.
+This aligns with the theoretical prediction that the second term in the cost bound (forecast error contribution) becomes less dominant when multiple forecast scenarios are aggregated.
+From an operational perspective, this has several implications.
+First, minimizing switching costs is not solely about reducing the number of policy changes.
+It also requires designing forecasts that are stable across time.
+As formalized in Section 3.3, vertical forecast stability directly mitigates the switching cost penalty.
+Stability metrics such as Mean Absolute Change (MAC) or Scenario Distribution Change (SDC) can thus serve as practical proxies to monitor and optimize for reduced control variability in real systems.
+Second, while the importance of forecast accuracy has long been emphasized in predictive control, this study shows that stability is an equally critical, yet underappreciated property, especially when switching costs are present.
+Systems that update policies too frequently may suffer from over-correction, chasing short-term forecast noise at the expense of long-term efficiency.
 6 Conclusion
@@
 In conclusion, this study highlights the critical role that switching costs, forecast accuracy, and stability play in the design and operation of energy management systems.
-Our analysis shows that in systems where forecasts guide decision making, stability of predictions can reduce the frequency and switching costs of policy revisions, thereby mitigating the financial and operational impacts.
+Our analysis shows that in systems where forecasts guide decision-making, stability of predictions can reduce the frequency and switching costs of policy revisions, thereby mitigating the financial and operational impacts.
 Our findings have significant implications for the design of energy management systems.
@@
 The current work suggests that enhancing the stability of forecasts leads to stability of policy which can improve system performance by managing the trade-offs between forecast accuracy and switching costs.
-Furthermore, we propose $SDC$, a novel metric to evaluating the stability of probabilistic forecasts, which provides a robust measure of the dissimilarity between scenario sets.
-Looking ahead, there are questions remaining to be addressed.
-For future research, we recommend further exploration into methods that enhance forecast stability without compromising accuracy.
-Further research into the link between forecast quality and optimization performance is needed.
-Parameters such as forecast horizon, commitment period, and the number of scenarios are subject to a study for deeper understanding of their impact on the optimization performance and dependencies to the forecast accuracy and stability.
-It would be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading.
-From literature and current findings, methods like AFHC and stochastic FHC, show more robustness in hedging against the forecast errors, however, they require more computational resources and are more complex to implement.
-An interesting avenue for future research would be defining the criteria for problems where these methods are most effective.
+Furthermore, we propose $SDC$, a novel metric to evaluate the stability of scenario sets.
+Our study has several important limitations that should be acknowledged.
+First, our theoretical analysis assumes convex cost functions and exponentially decaying forecast errors, which may not fully capture the complexity of real-world energy systems.
+Second, while we demonstrate the relationship between forecast stability and optimization performance on a single case study, further validation across different energy management contexts and geographical locations would strengthen our findings.
+Additionally, our definition of switching costs primarily focuses on grid ramping, which serves as a proxy for the more complex switching costs in real energy systems, including wear-and-tear, efficiency losses, and network fees.
+Future work should address these limitations by testing across diverse datasets, and exploring more sophisticated methods for enhancing forecast stability without compromising accuracy.
 These efforts would contribute to a deeper understanding of forecast-dependent optimization and improve the operational strategies of energy management systems facing an evolving energy landscape.
+It would also be beneficial to extend this work to other applications with switching costs, such as supply chain management and financial trading.
 Acknowledgments
@@
 The optimization score is defined as:
-Average $\text{Average Score}=\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}}\right)$ (10)
+Average $\text{Average Score}=\operatorname{avg}\left(\frac{C_{\text{ALG}}}{C_{\text{no\ battery}}},\frac{G_{\text{ALG}}}{G_{\text{no\ battery}}}\right)$ (29)
 When grid-related KPIs are included, the average includes the grid score $D$.
@@
 […]
-$\operatorname{avg}\left(\frac{C_{\text{entry}}}{C_{\text{no\ battery}}},\frac{G_{\text{entry}}}{G_{\text{no\ battery}}},D\right)$
+$\operatorname{avg}\left(\frac{C_{\text{ALG}}}{C_{\text{no\ battery}}},\frac{G_{\text{ALG}}}{G_{\text{no\ battery}}},D\right)$
 All the scores range from 0 to 1 because they are normalized by the no-battery scenario, representing the improvement as a fraction of the metric compared to the no-battery scenario.
 The normalized electricity cost, $C$, as delineated in Eq.
-12, is the ratio of electricity spending in a given policy, $c_{policy}$, to the spending in a reference scenario without battery intervention, $c_{{nobattery}}$ The cost metric $c$ is further defined in Eq.
+31, is the ratio of electricity spending in a given policy, $c_{policy}$, to the spending in a reference scenario without battery intervention, $c_{{nobattery}}$ The cost metric $c$ is further defined in Eq.
 32 as the aggregate of the non-negative product of district-level net electricity price, $E_{h}\times T_{h}(\$)$, where $E_{h}$ denotes the electricity consumption of the district at the hour $h$ and $T_{h}$ specifies the electricity rate corresponding to that hour.
@@
 The formulation, scaled by the grid cost in the baseline no-battery scenario, is given in Eq.
-16.
-$D=\operatorname{avg}\left(\frac{R_{\text{entry}}}{R_{\text{no\ battery}}},\frac{1-L_{\text{entry}}}{1-L_{\text{no\ battery}}}\right)$ (16)
+35.
+$D=\operatorname{avg}\left(\frac{R_{\text{ALG}}}{R_{\text{no\ battery}}},\frac{1-L_{\text{ALG}}}{1-L_{\text{no\ battery}}}\right)$ (35)
 The ramping KPI, $R$, reflects the smoothness of the district’s load profile.
@@
 A.2 Optimization Formulation
-$\mathbf{C}(S_{t},x_{t},\theta^{LA}_{t})=\sum_{i\in{price,carbon,grid}}\mathbf{C}_{i}(\tilde{S}{tt^{\prime}},\tilde{x}_{tt^{\prime}})$ (19)
+$\mathbf{C}(S_{t},x_{t},\theta^{LA}_{t})=\sum_{i\in{price,carbon,grid}}\mathbf{C}_{i}(\tilde{S}{tt^{\prime}},\tilde{S}_{tt^{\prime}})$ (38)
 $=$ (39)
@@
 In this formulation:
-  • 1.
-    $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the lookahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
-  • 2.
+  • $X_{t}^{FHC}(S_{t}|\theta^{LA})$ Represents the optimal set of decisions (battery charging $x_{t}^{pos}$ and discharing $x_{t}^{neg}$ actions) made by the FHC at time $t$ based on the current state $S_{t}$ and the look-ahead approximation model $\theta^{LA}$ This model enables predicting future costs within a horizon $H$, taking into account the expected electric load and PV generation.
   • $W_{s}$ Weight factor for scenario $s$, indicating the importance of different scenarios in the decision-making process.
@@
 In the deterministic case, the number of scenarios is equal to 1.
-  • 3.
   • $C_{price}$, $C_{carbon}$, and $C_{grid}$ are cost functions representing the cost of electricity, the cost associated with carbon emissions, and the cost related to grid reliance, respectively.
 Each of these costs depends on the action taken ($x_{t}^{pos}$ and $x_{t}^{neg}$) and the current system state $S_{t}$.
-  • 4.
   • $SOC_{t}$ denotes the state of charge of the battery at time $t$, with $SOC_{min}$ and $SOC_{max}$ being the minimum and maximum allowable states of charge, respectively.
-  • 5.
   • $P_{max}$ is the maximum power with which the battery can be charged or discharged.
-  • 6.
   • $\eta_{charging}$ and $\eta_{discharging}$ are the charging and discharging efficiencies of the battery, respectively.
-A.3 Theoretical Analysis of Performance Bounds
-In this section, we analyze the upper bound on the performance of the FHC algorithm with a limited commitment $v$.
-We focus on the cost difference between the optimal policy and the policy produced by the FHC algorithm.
-Specifically, we examine the impact of the commitment $v$, which defines how often the policy and forecast are updated.
-Our goal is to show how the performance bound is influenced by a trade-off between forecast accuracy and switching costs.
-To illustrate this, we consider two models of forecast error: independent and identically distributed (i.i.d.) noise, noise with exponentially decaying correlation and a probabilistic forecast.
-We start with the deterministic case (point forecast) and then extend to the probabilistic case (set of scenarios).
-Following the derivation in Chen et al.
-[18], we derive the average-case performance of the FHC algorithm by taking the expectation of the cost difference with respect to the prediction noise $(e(t))^{T}_{t=1}$.
-Here, OPT represents the optimal policy that would achieve the lowest possible cost if there were perfect information and no prediction errors.
-$\mathbb{E}\operatorname{cost}(FHC)\leq$ (26)
-$\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2G\mathbb{E}\sum_{t=1}^{T}\left\|y_{t}-y_{t\mid t-\phi^{1}(t)}\right\|_{2}^{\alpha}$
-$\mathbb{E}\operatorname{cost}(FHC)\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2GT\left\|f_{v}\right\|^{\alpha}.$ (27)
-The term $\left\|f_{v}\right\|^{\alpha}$ is the $\alpha$-norm of the prediction error covariance, where $\alpha$ is the exponent of the Hölder condition.
-The two noise models (i.i.d.
-and exponentially decaying correlation) are explained below.
-Independent and Identically Distributed Noise
-The i.i.d. noise is a simple and understandable baseline for forecast error, without time dependency. For i.i.d. noise, we have:
-$f(s)=\begin{cases}I,&s=0\\ 0,&\text{otherwise.}\end{cases}$
-The statistical properties of prediction errors in this context are:
-$\mathbb{E}\left[e(s)e(s)^{T}\right]=R_{e}$
-$\operatorname{trace}(R_{e})=\sigma^{2}$
-This leads to the upper bound on the FHC cost:
-$\mathbb{E}\operatorname{cost}(FHC(v))\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2GT\sigma^{\alpha}$ (28)
-The upper bound on the competitive difference depends on $v$ only in the first term (in the denominator).
-Therefore, the optimal cost is minimized when $v=H$, i.e., the commitment level reaches the maximum allowable duration and corresponds to the forecast horizon length.
-The first term represents switching costs, so a shorter commitment period leads to more frequent updates and higher switching costs.
-The second term, representing i.i.d.
-noise, is independent of $v$.
-Exponentially Decaying Correlation
-A more realistic model of prediction noise uses correlations that decay exponentially over time. For this, given $a<1$, the Frobenius norm of the correlation structure is:
-$\|f(s)\|_{F}=\begin{cases}ca^{s},&s\geq 0\\ 0,&s<0\end{cases}$ (29)
-We then express the prediction error covariance as:
-$\left\|f_{v}\right\|^{2}$ $=\sum_{s=0}^{v}\operatorname{trace}(R_{e}f(s)^{T}f(s))=\sum_{s=0}^{v}\left\langle R_{e}^{1/2},f(s)\right\rangle^{2}$
-$\leq\sum_{s=0}^{v}\left(\left\|R_{e}^{1/2}\right\|_{F}\|f(s)\|_{F}\right)^{2}$
-$=\sum_{s=0}^{v}c^{2}\sigma^{2}a^{2s}=c^{2}\sigma^{2}\frac{1-a^{2v}}{1-a^{2}}$
-Given that $h$ exhibits G-Lipschitz continuity ($\alpha=1$), the norm of the correlation function $\|f_{v}\|$ is bounded by:
-$\left\|f_{v}\right\|\leq\frac{c\sigma(1-a^{2(v+1)}/2)}{1-a^{2}}$ (30)
-Plugging this into the expression for the expected competitive difference of the FHC algorithm, we get:
-$\mathbb{E}\operatorname{cost}(FHC(v))\leq$ (31)
-$\mathbb{E}\operatorname{cost}(OPT)+2\left[\frac{T}{v}\right]\beta D+2GT\sigma^{\alpha}\frac{c(1-a^{2(v+1)})/2}{1-a^{2}}$
-$\leq\mathbb{E}\operatorname{cost}(OPT)+2\left[\frac{T}{v}\right]\beta D+2GTc\sigma\left(\frac{a^{2v+2}-2}{2(a^{2}-1)}\right)$
-To simplify, we group the parameters independent of $v$:
-$A=2T\beta D\quad B=\frac{1}{a^{2}-1}\quad C=GTc\sigma a^{2}$
-Thus, the expected competitive difference of the FHC algorithm is:
-$\mathbb{E}\operatorname{cost}(FHC(v))\leq\mathbb{E}\operatorname{cost}(OPT)+\frac{A}{v}+B(Ca^{2v}-1)$ (32)
-The competitive difference behavior depends on two main components.
-The term $\frac{A}{v}$ decreases with $v$, while the term $B(Ca^{2v}-1)$ increases with $v$.
-Since $Ca^{2v}$ approaches 0 as $v$ increases, the second term is bound by $B$.
-Therefore, the minimal competitive difference is a trade-off between switching costs (first term) and forecast errors (second term).
-Stochastic Optimization
-For stochastic optimization with Fixed Horizon Control (FHC) applied across $n$ scenarios, we examine how the stochastic nature influences the competitive difference compared to an optimal solution.
-The expected cost for the stochastic algorithm across all scenarios is:
-$\mathbb{E}\operatorname{cost}(FHC_{stochastic}(v))=\sum_{i=1}^{n}\mathbb{E}\operatorname{cost}(FHC(v,i))$ (33)
-Taking the expectation of the competitive difference across all scenarios:
-$\mathbb{E}\operatorname{cost}(FHC_{stochastic}(v))\leq$ (34)
-$\mathbb{E}\operatorname{cost}(OPT)+\frac{2T\beta D}{v}+2G\mathbb{E}\left[\sum_{t=1}^{T}\left\|y_{t}-y_{t\mid t-\phi^{1}(t)}\right\|_{2}^{\alpha}\right]$
-In the stochastic case, the forecast errors across the scenario ensemble would be lower in expectation than in any singular, deterministic scenario.
-This is because the stochastic approach averages out the errors across multiple scenarios, reducing the overall forecast error.
-Hence, the competitive difference of the FHC algorithm in the stochastic case will likely be lower, favoring a higher commitment level, as the first term (switching costs) decreases with $v$ at the same rate as in the deterministic case.
-A.4 In-sample Performance
-(a) Performance of the FHC algorithm with point forecast.
-Study over combinations of forecast and optimization commitment between 1 and 12
-$v_{O}$
-$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
+A.3 In-sample Performance
+Table 5: Optimization KPI scores for combinations of forecast ($v_{F}$) and optimization ($v_{O}$) commitment periods between 1 and 12 hours (with data from in-sample buildings)
+Optimization Commitment Period ($v_{O}$)
+Forecast ($v_{F}$) 1 2 3 4 5 6 7 8 9 10 11 12
+Panel A: Deterministic FHC with point forecast
 1 0.486
@@
 12 0.495 0.495 0.495 0.495 0.495 0.495 0.496 0.496 0.495 0.496 0.496 0.495
-(b) Performance of the FHC algorithm with stochastic forecast.
-Study over combinations of forecast and optimization commitment between 1 and 12
-$v_{O}$
-$v_{F}$ 1 2 3 4 5 6 7 8 9 10 11 12
+Panel B: Stochastic FHC with probabilistic forecast
 1 0.798
@@
 12 0.813 0.809 0.807 0.805 0.806 0.808 0.807 0.806 0.807 0.808 0.808 0.807
-Table 4: Optimization KPIs score for combinations of forecast and optimization commitment between 1 and 12 (with data from in-sample buildings)
+Note: Diagonal elements represent equal commitment periods for both forecast and optimization.
+Lower values indicate better performance.
+Empty cells (—) represent invalid combinations where $v_{O}>v_{F}$.
 Accuracy
+Stability: vertical (top) and horizontal (bottom)
 Optimization Score: without switching costs
-Stability: vertical (top) and horizontal (bottom)
 Optimization Score: with switching costs
-Point Forecast + Deterministic FHC
-Probabilistic forecast + Stochastic FHC
-Figure 7: Accuracy (Top), Horizontal (Middle) and Vertical Stability (Bottom row) of stochastic forecast over different revision periods with data from in-sample buildings
+Point Forecast
++ Deterministic FHC
+Probabilistic forecast
++ Stochastic FHC
+Figure 8: Accuracy (Top), Horizontal (Second) and Vertical Stability (Third row) of stochastic forecast over different revision periods with data from in-sample buildings.